### v1

# Khintchine dichotomy for self-similar measures

Abstract.
We establish the analogue of Khintchine’s theorem for all self-similar probability measures on the real line. When specified to the case of the Hausdorff measure on the middle-thirds Cantor set, the result is already new and provides an answer to an old question of Mahler. The proof consists in showing effective equidistribution in law of expanding upper-triangular random walks on $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$, a result of independent interest.
2010 Mathematics Subject Classification:
Primary 37A99, 11J83; Secondary 22F30.
W.H. is supported by the National Key R&D Program of China (No. 2022YFA1007500) and the National Natural Science Foundation of China (No. 12288201).
H.Z. acknowledges the support of startup grant of Soochow University.
1. Introduction
A Borel probability measure $\sigma$ on the real line $\mathbb{R}$ is called self-similar if it satisfies
(1) $\sigma=\sum_{i=1}^{\mathtt{m}}\lambda_{i}\,\phi_{i\star}\sigma$
for some integer $m\geq 1$, some probability vector $(\lambda_{1},\cdots,\lambda_{\mathtt{m}})\in\mathbb{R}_{>0}^{\mathtt{m}}$, and some affine maps $\phi_{1},\dotsc,\phi_{\mathtt{m}}:\mathbb{R}\to\mathbb{R}$ without common fixed point. This includes Hausdorff measures on missing digit Cantor sets. For example, the one on the middle-thirds Cantor set satisfies (1) with $\lambda_{1}=\lambda_{2}=1/2$ and $\phi_{1}:x\mapsto x/3$ and $\phi_{2}:x\mapsto x/3+2/3$. Another standard definition of self-similar measures adds the restriction that the maps $\phi_{i}$ are all contracting, we do not make such an assumption.
It is particularly intriguing to explore the Diophantine properties of points within the support of a self-similar measure. This research topic was proposed by Mahler in [31, Section 2], asking how close irrational numbers in the middle-thirds Cantor set can be approximated by rational numbers. One approach to framing Mahler’s question is by investigating whether Khintchine’s theorem extends to the middle-thirds Cantor measure (as asked by Kleinbock-Lindenstrauss-Weiss in [23, Section 10.1]).
Let us recall the classical Khintchine theorem. Here and hereafter, $\psi:\mathbb{N}\to\mathbb{R}_{>0}$ is a function that will be referred to as an approximation function. A point $x\in\mathbb{R}$ is called $\psi$-approximable if there exist infinitely many $(p,q)\in\mathbb{Z}\times\mathbb{N}$ such that
$|qx-p|<\psi(q).$
Denote by $W(\psi)$ the set of $\psi$-approximable points in $\mathbb{R}$. The classical Khintchine theorem for the Lebesgue measure [20, 21] states that given a non-increasing approximation function $\psi$, the set $W(\psi)$ has null Lebesgue measure if $\sum\psi(q)<\infty$, and full Lebesgue measure if $\sum\psi(q)=\infty$.
In this paper, we establish the analogue of Khintchine’s theorem for self-similar measures on $\mathbb{R}$.
Theorem A (Khintchine’s theorem for self-similar measures).
Let $\sigma$ be a self-similar probability measure on $\mathbb{R}$, let $\psi:\mathbb{N}\to\mathbb{R}_{>0}$ be a non-increasing function. Then
(2) $\sigma(W(\psi))=\left\{\begin{array}[]{ll}0&\text{if }\sum_{q\in\mathbb{N}}\psi(q)<\infty,\\ &\\ 1&\text{if }\sum_{q\in\mathbb{N}}\psi(q)=\infty.\end{array}\right.$
Let us briefly present the state of the art surrounding Khintchine’s theorem on fractals.
For the convergence part, the case $\psi(q)=1/q^{1+\varepsilon}$ was treated by Weiss [40] for measures satisfying certain decay conditions, comprising the case of the middle-thirds Cantor measure. Weiss’ result was later generalized to friendly measures on $\mathbb{R}^{d}$ for arbitrary positive integer $d$ by Kleinbock-Lindenstrauss-Weiss [23]. See also the related work of Pollington-Velani [32] on absolutely friendly measures.
For the divergence part, the case $\psi(q)=\varepsilon/q$ was treated by Einsiedler-Fishman-Shapira [16] for missing digit Cantor measures. Simmons-Weiss [36] then significantly generalized their result, promoting it to self-similar measures on $\mathbb{R}^{d}$ arising from a contractive IFS $(\phi_{i})_{1\leq i\leq m}$ satisfying the open set condition.
All the above works focus on specific approximation functions $\psi$. Under the sole condition that $\psi$ is non-increasing, Khalil and Luethi [19] were able to extend Khintchine’s theorem to self-similar measures $\sigma$ on $\mathbb{R}^{d}$, provided $\sigma$ has large dimension and the underlying IFS $(\phi_{i})_{1\leq i\leq m}$ is contractive, rational, and satisfies the open set condition. In particular, they derived Khintchine’s theorem for one-missing digit Cantor sets in base $5$. With a different approach, based on Fourier analysis, Yu [42] also achieved the convergence part of Khintchine’s theorem for general approximation functions, provided $\sigma$ is a missing digit Cantor measure with sufficiently fast Fourier decay. The divergence part was very recently settled by Datta-Jana [14] under similar conditions, reaching some cases that are not covered by [19] such as a $4$-missing digit Cantor set in base $450$.
All the aforementioned works impose various constraints for the Khintchine dichotomy (2) to hold for a fractal measure. Specifically, none of them establishes (2) in the case of the middle-thirds Cantor measure advertised by Mahler. A is already innovative in this regard, and goes way beyond this example.
Other related research topics. As Mahler pointed out in [31], it is also interesting to investigate intrinsic Diophantine approximation on a Cantor set. This means asking how well points on a fractal set can be approximated by rational points sitting inside the fractal set itself. We refer to recent works [39, 11] and references therein for related research in this direction.
In addition to fractals, Khintchine’s theorem has been extensively studied on submanifolds of $\mathbb{R}^{d}$. Major works in this area include [26, 8, 9].
A is derived from an effective equidistribution result in homogeneous dynamics which we now present. Consider the real algebraic group $G=\operatorname{SL}_{2}(\mathbb{R})$, a lattice $\Lambda\subseteq G$, and the quotient space $X=G/\Lambda$, endowed with the standard hyperbolic metric (§2.1) and the Haar probability measure $m_{X}$. Write $\operatorname{inj}(x)$ the injectivity radius at $x\in X$2.1). Denote by $B^{\infty}_{\infty,1}(X)$ the set of smooth functions on $X$ which are bounded and have bounded order-$1$ derivatives, write ${\mathcal{S}}_{\infty,1}(\cdot)$ the corresponding Sobolev norm (§2.1). For $t>0$, $s\in\mathbb{R}$, write $a(t),u(s)\in G$ the elements given by
$a(t)=\begin{pmatrix}t^{1/2}&0\\ 0&t^{-1/2}\end{pmatrix}\qquad u(s)=\begin{pmatrix}1&s\\ 0&1\end{pmatrix}.$
Theorem B (Effective equidistribution of expanding fractals).
Let $\sigma$ be a self-similar probability measure on $\mathbb{R}$. There exists a constant $c>0$ such that for all $t>1$, $x\in X$, $f\in B^{\infty}_{\infty,1}(X)$, we have
(3) $\int_{\mathbb{R}}f(a(t)u(s)x)\,\mathrm{d}\sigma(s)=\int_{X}f\,\mathrm{d}m_{X}\,+\,O\bigl{(}\operatorname{inj}(x)^{-1}t^{-c}{\mathcal{S}}_{\infty,1}(f)\bigr{)}$
where the implicit constant in $O(\cdot)$ only depends on $\Lambda$ and $\sigma$.
B states the exponential equidistribution of the measure $\sigma$ seen on a piece of horocycle based at $x$ and expanded by the action of the geodesic flow. The exponent $c$ in the rate of equidistribution is uniform in $x$, however, equidistribution may take more time to start when $x$ is high in the cusp. This is reflected by the term $\operatorname{inj}(x)^{-1}$ in the rate.
The link between homogeneous dynamics and Diophantine approximation is known as Dani’s correspondence [13]. In [27], Kleinbock-Margulis explicitely demonstrated how to use dynamics to obtain a new proof of the classical Khintchine theorem for the Lebesgue measure, see also the variant [38] by Sullivan. This dynamical perspective laid the foundation for many subsequent works generalizing Khintchine’s theorem in various aspects, see e.g. [24, 12, 19]. In particular, the implication from B to A is given in the work of Khalil-Luethi [19] (assuming contractive IFS with the open set condition for the divergence part). Note that for this implication, it is only necessary to consider the case $\Lambda=\operatorname{SL}_{2}(\mathbb{Z})$ and prove the above equidistribution statement for specific points $x$ related to $\sigma$. Our equidistribution statement is valid for all lattices $\Lambda\subseteq G$ and all $x\in X$.
Besides its applications to Diophantine approximation, B is interesting in its own right. It can be seen as a fractal and effective version of Ratner’s equidistribution theorem for unipotent flows on $X$. Recall that Ratner’s theorem states that any unipotent orbit on a finite-volume homogeneous space equidistributes within the smallest finite-volume homogeneous subspace that contains it. Unfortunately, the proof gives no information on the rate of equidistribution. Over the past years, substantial efforts were made to obtain an effective version of Ratner’s theorem, in other terms quantify the equidistribution of large but bounded pieces of unipotent orbits. In the case where the unipotent orbit arises from the action of a horospherical subgroup, Kleinbock and Margulis established in [25, 28] the effective equidistribution of expanding translates under the corresponding diagonal flow. More recently, significant progress on effective Ratner was made by Einsiedler-Margulis-Venkatesh [17], Strömbergsson [37], Kim [22], Lindenstrauss-Mohammadi [29], Lindenstrauss-Mohammadi-Wang [30], and Yang [41]. We note that these works focus on the expanding translates of the Haar measure on a piece of a unipotent orbit. In [19], Khalil and Luethi obtain the first effective equidistribution of expanding fractal measures on a unipotent orbit in $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$. They argue under the assumption that the underlying IFS is contractive, rational, satisfies the open set condition, and the measure $\sigma$ is thick enough. They also require that the starting point $x$ belongs to a specific countable set related to the IFS. B generalizes Khalil-Luethi’s equidistribution result in $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$ by allowing an arbitrary lattice and eliminating the constraints on the IFS, the measure $\sigma$, and the starting point.
We prove B from the point of view of random walks. The connection between the asymptotic behaviour of an expanding fractal and that of a random walk is rooted in the work of Simmons-Weiss [36] and then further exploited in [33, 34, 19, 15, 1]. In our paper, this connection takes the form of Lemma 5.4.
We establish the following effective equidistribution in law for random walks driven by expanding upper triangular matrices on $X$. In the statement below, $\mathbb{R}^{2}$ is endowed with the usual Euclidean structure and we write $e_{1}:=(1,0)\in\mathbb{R}^{2}$.
Theorem C (Effective equidistribution for random walks).
Let $\mu$ be a finitely supported probability measure on the group
$\{\,a(t)u(s):t>0,\,s\in\mathbb{R}\,\}\subseteq G.$
Assume that the support of $\mu$ is not simultaneously diagonalisable, and $\mu$ satisfies $\int_{G}\log\|ge_{1}\|\,\mathrm{d}\mu(g)>0$. Then there exists a constant $c>0$ such that for all $x\in X$, $n\geq 1$ and $f\in B^{\infty}_{\infty,1}(X)$, we have
$\mu^{*n}*\delta_{x}(f)=m_{X}(f)+O\bigl{(}\operatorname{inj}(x)^{-1}e^{-cn}{\mathcal{S}}_{\infty,1}(f)\bigr{)}$
where the implicit constant in $O(\cdot)$ only depends on $\Lambda$ and $\mu$.
The proof of C is inspired by [4], where the first-named and second-named authors establish effective equidistribution for random walks on $X$ which are driven by a Zariski-dense probability measure on $G$. In our context however, the acting group is solvable. The proof consists of three phases. Each step concerns the dimension of the distribution of the random walk at some scale.
First, we show that the random walks gains some initial positive dimension: there exist constants $\kappa>0$, $A>0$ determined by $\mu$ such that for every small $\rho>0$, $x,y\in X$, every $n\geq|\log\rho|+A|\log\operatorname{inj}(x)|$,
${\mu^{*n}}*\delta_{x}(B_{\rho}y)\leq\rho^{\kappa}$
where the notation $B_{\rho}y$ refers to the open ball of radius $\rho$ centered at $y$ in $X$.
Second, we bootstrap the value of the exponent $\kappa$ arbitrarily close to $3$, say up to $3-\varepsilon$ provided $\rho\leq\rho_{0}(\varepsilon,\mu)$ and $n\geq C_{0}(\varepsilon,\mu)|\log\rho|+A|\log\operatorname{inj}(x)|$. The method is based on the multislicing argument from [4].
Finally, once the dimension is close to full, we conclude using the spectral gap of the convolution operator $f\mapsto\mu*f$ acting on $L^{2}(X)$.
B follows from C, using Lemma 5.4 and a probabilistic argument. A is obtained from B, case $\Lambda=\operatorname{SL}_{2}(\mathbb{Z})$, thanks to [19, Th.9.1, Th.12.1] (up to a few adaptations to remove the assumptions of contractiveness and open set condition from [19, Th.12.1 ]).
Allowing $\lambda$ to have infinite support. Our method allows for slightly more general statements, extending the aforementioned Khintchine dichotomy and equidistribution results to measures arising from randomized IFS with potentially infinite support, provided a finite exponential moment.
We let $\operatorname{Aff}(\mathbb{R})$ denote the affine group of $\mathbb{R}$. For any $\phi\in\operatorname{Aff}({\mathbb{R}})$, we let $\mathtt{r}_{\phi}\in\mathbb{R}^{*}$, $\mathtt{b}_{\phi}\in\mathbb{R}$ denote the unique numbers such that
(4) $\phi(t)=\mathtt{r}_{\phi}t+\mathtt{b}_{\phi},\quad\forall t\in\mathbb{R}.$
We say a probability measure $\lambda$ on $\operatorname{Aff}(\mathbb{R})$ has a finite exponential moment if there exists $\varepsilon>0$ such that
(5) $\int_{\operatorname{Aff}(\mathbb{R})}|\mathtt{r}_{\phi}|^{\varepsilon}+|\mathtt{r}_{\phi}^{-1}|^{\varepsilon}+|\mathtt{b}_{\phi}|^{\varepsilon}\,\mathrm{d}\lambda(\phi)<\infty.$
Theorem A’.
Let $\lambda$ be a probability measure on $\operatorname{Aff}(\mathbb{R})$ with finite exponential moment and such that $\operatorname{supp}\lambda$ does not have a global fixed point. Let $\sigma$ be a probability measure on $\mathbb{R}$ satisfying $\lambda*\sigma=\sigma$. Then $\sigma$ satisfies the Khintchine dichotomy (2).
Theorem B’.
Every measure $\sigma$ as in A’ satisfies the effective equidistribution for expanding translates from Equation 3.
Recall a probability measure $\mu$ on $G$ has finite exponential moment if there exists $\varepsilon>0$ such that
(6) $\int_{G}\|g\|^{\varepsilon}\,\mathrm{d}\mu(g)<\infty.$
Theorem C’.
C is valid when the finite support assumption on $\mu$ is relaxed into a finite exponential moment condition.
Further comments. We have been informed that B intersects an ongoing work of Khalil-Luethi-Weiss. In a paper in preparation, they show that a self-similar measure in $\mathbb{R}^{d}$ seen on a piece of unipotent orbit and expanded by the action of any diagonal flow equidistributes (without rate) toward the Haar measure in $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$, provided the self-similar measure arises from a contractive rational IFS with equal contraction ratios and no rotation in the linear part. For $d=1$, this result is also covered by B. Their method of proof is different from ours, as it relies on measure rigidity results for $S$-arithmetic random walks and builds on the work of Benoist-Quint [5, 6, 7].
Structure of the paper. In Section 2, we fix notations for the rest of the paper, we highlight basic facts concerning self-similar measures, and we recall some recurrence properties of the $\mu$-walk on $X$. In Section 3, we deduce positive dimension of ${\mu^{*n}}*\delta_{x}$ at exponentially small scales. In Section 4, we bootstrap the dimension until it reaches a number arbitrarily close to $3=\dim X$. In Section 5, we deduce the equidistribution statements, namely B’ and C’. In Section 6, we obtain the analogue of Khintchine’s theorem for self similar measures, A’.
Acknowledgements. The authors thank Nicolas de Saxcé for sharing his insight on random walks and Diophantine approximation. W.H. and H.Z. thank Barak Weiss for enlightening discussions. H.Z. thanks Ronggang Shi for his encouragement.
2. Preliminaries
In this section, we set up notations and collect basic facts that will be useful for the rest of the paper.
2.1. Notation and Conventions
Throughout this paper, $G=\operatorname{SL}_{2}(\mathbb{R})$, $\Lambda\subseteq G$ is a lattice, and $X=G/\Lambda$.
Metric. We fix a basis $(e_{-},e_{0},e_{+})$ of the Lie algebra $\mathfrak{g}=\operatorname{Lie}(G)$ given by
$e_{-}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}\,\,\,\,\,\,e_{0}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\,\,\,\,\,\,e_{+}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$
We assume throughout that $G$ is endowed with the unique right-invariant Riemannian metric for which $(e_{-},e_{0},e_{+})$ is orthonormal. This induces a distance on $G$ and the quotient $X$ that we denote by $\operatorname{dist}$ in both cases. Given $\rho>0$, we write $B_{\rho}$ to denote the open ball of radius $\rho>0$ centered at the neutral element $\operatorname{Id}$ in $G$. Then the open ball of radius $\rho$ centered a point $x\in X$ can be written as $B_{\rho}x$.
The injectivity radius of $X$ at a point $x$ is
$\operatorname{inj}(x)=\inf\{\,\rho>0\,:\,\text{the map $B_{\rho}\rightarrow X,g\mapsto gx$ is injective}\,\}.$
Sobolev norms. Set $\Xi_{l}$ the words on the alphabet $\{e_{-},e_{0},e_{+}\}$ of length at most $l$. Each $D\in\Xi_{l}$ acts as a differential operator on the space of smooth functions $C^{\infty}(X)$. Given $f\in C^{\infty}(X)$, $k,l\in\mathbb{N}\cup\{\infty\}$, we set
${\mathcal{S}}_{k,l}(f)=\sum_{D\in\Xi_{l}}\|Df\|_{L^{k}}$
where $\|\cdot\|_{L^{k}}$ refers to the $L^{k}$-norm for the Haar probability measure on $X$. We let $B^{\infty}_{k,l}(X)$ denote the space of bounded smooth functions $f$ on $X$ such that ${\mathcal{S}}_{k,l}(f)<\infty$.
Haar measure. Let $m_{G}$ denote the Haar measure on $G$ normalised so that the $G$-invariant Borel measure $m_{X}$ it induces on $X$ is a probability measure.
Driving measures $\lambda$ and $\mu$. Let $\operatorname{Aff}(\mathbb{R})^{+}$ denote the group of orientation preserving affine transformations of the real line. Denote by
$P=\{\,a(t)u(s):t>0,\,s\in\mathbb{R}\,\}\subseteq G$
the subgroup of upper triangular matrices with positive diagonal entries. For every $g\in P$, we let $\mathtt{r}_{g}\in\mathbb{R}_{>0}$ and $\mathtt{b}_{g}\in\mathbb{R}$ be the unique numbers such that
$g=a(\mathtt{r}_{g})^{-1}u(\mathtt{b}_{g})=\begin{pmatrix}\mathtt{r}_{g}^{-1/2}&\mathtt{r}_{g}^{-1/2}\mathtt{b}_{g}\\ 0&\mathtt{r}_{g}^{1/2}\end{pmatrix}$
We identify $P$ with $\operatorname{Aff}(\mathbb{R})^{+}$ by mapping $g\in P$ with the similarity $x\mapsto\mathtt{r}_{g}x+\mathtt{b}_{g}$. This is an anti-isomorphism between the two groups.
Fix a probability measure $\lambda$ on $\operatorname{Aff}(\mathbb{R})^{+}$ with support $\operatorname{supp}(\lambda)$, denote by $\mu$ the corresponding probability measure on $P$ via the above anti-isomorphism. Throughout this paper, $\lambda$ and $\mu$ determine each other in this way. For $n\in\mathbb{N}$, we write $\lambda^{*n}=\lambda*\dotsm*\lambda$ to denote the $n$-fold convolution of $\lambda$ with itself, we define $\mu^{*n}$ similarly.
We assume that $\lambda$, and equivalently $\mu$, has a finite exponential moment. With our notations, this means there exists $\varepsilon>0$ such that
$\int_{P}|\mathtt{r}_{g}|^{\varepsilon}+|\mathtt{r}_{g}^{-1}|^{\varepsilon}+|\mathtt{b}_{g}|^{\varepsilon}\,\mathrm{d}\mu(g)<\infty.$
We assume that $\operatorname{supp}\lambda$ does not have a global fixed point in $\mathbb{R}$. This amounts to saying that $\operatorname{supp}(\mu)$ does not have two common fixed points on the projective line, or alternatively, that the matrices in $\operatorname{supp}(\mu)$ are not simultaneously diagonalisable.
Self-similar measure $\sigma$. Throughout this paper, we let $\sigma$ denote a probability measure on $\mathbb{R}$ that is $\lambda$-stationary, which means
$\sigma=\int_{\operatorname{Aff}(\mathbb{R})}\phi_{\star}\sigma\,\mathrm{d}\lambda(\phi).$
By a theorem of Bougerol-Picard [10, Theorem 2.5], the existence of such $\sigma$ is equivalent to the condition:
(7) $\int_{P}\log\mathtt{r}_{g}\,\mathrm{d}\mu(g)<0,$
i.e. the random walk on $\mathbb{R}$ driven by $\lambda$ is contractive in average. Moreover, provided existence, the measure $\sigma$ is uniquely determined by $\lambda$, see [10, Corollary 2.7].
Lyapunov exponent. Let $\operatorname{Ad}:G\rightarrow\operatorname{Aut}(\mathfrak{g})$ be the adjoint representation. We denote by $\ell$ the top Lyapunov exponent associated to $\operatorname{Ad}_{\star}\mu$. It is determined only by the diagonal terms and is equal to
$\ell=-\int_{P}\log\mathtt{r}_{g}\,\mathrm{d}\mu(g)>0.$
[…]
2.2. Regularity of the self-similar measure
We first recall that the measure $\sigma$ has finite moment of positive order and is Hölder regular. We often refer to the second property as having positive dimension (at all scales).
Lemma 2.1 (Moment and Hölder-regularity of $\sigma$).
There exists $\gamma>0$ such that
$(i)\,\,\int_{\mathbb{R}}|s|^{\gamma}\,\mathrm{d}\sigma(s)<\infty,\qquad\quad(ii)\,\,\forall r>0,\,\sup_{s\in\mathbb{R}}\sigma(s+[-r,r])\ll r^{\gamma}.$
Proof.
(i) follows from [18, Proposition 5.1] and (ii) is a consequence of [2, Theorem 2.12]. ∎
Given an integer $n\in\mathbb{N}$, denote by $\sigma^{(n)}$ the image measure of $\mu^{*n}$ under the map $g\in P\mapsto\mathtt{b}_{g}\in\mathbb{R}$. Equivalently, $\sigma^{(n)}={\lambda^{*n}}*\delta_{0}$, where $\delta_{0}$ denotes the Dirac measure at $0\in\mathbb{R}$. We show that the measures $\sigma^{(n)}$ have a uniformly finite positive moment, and uniform positive dimension above an exponentially small scale. For this, we first observe that $\sigma^{(n)}$ converges toward $\sigma$ at exponential rate. We denote by $\operatorname{Lip}(\mathbb{R})$ the space of bounded Lipschitz functions on $\mathbb{R}$ with the norm $\|f\|_{\operatorname{Lip}}=\|f\|_{\infty}+\sup_{s\neq t}\frac{|f(s)-f(t)|}{|s-t|}$.
Lemma 2.2.
There exists $\varepsilon>0$ such that for all $n\geq 0$, all $f\in\operatorname{Lip}(\mathbb{R})$, we have
$|\sigma^{(n)}(f)-\sigma(f)|\ll e^{-\varepsilon n}\|f\|_{\operatorname{Lip}}.$
Proof.
We may assume $\|f\|_{\operatorname{Lip}}=1$. Then we have,
$|\sigma^{(n)}(f)-\sigma(f)|$ $=|\lambda^{*n}*\delta_{0}(f)-\lambda^{*n}*\sigma(f)|$
$\leq\int_{\operatorname{Aff}(\mathbb{R})^{+}\times\mathbb{R}}|f(\phi(0))-f(\phi(s))|\,\mathrm{d}(\lambda^{*n}\otimes\sigma)(\phi,s)$
(8) $\leq\int_{\operatorname{Aff}(\mathbb{R})^{+}\times\mathbb{R}}\min(2,\,\mathtt{r}_{\phi}|s|)\,\mathrm{d}(\lambda^{*n}\otimes\sigma)(\phi,s),$
where $\mathtt{r}_{\phi}>0$ denotes the dilation factor in the affine map $\phi$, see (4). Using the principle of large deviations and that $\lambda$ is contracting in average (7), we have for $\varepsilon\lll 1$,
(9) $\lambda^{*n}\{\,\phi\,:\,\mathtt{r}_{\phi}>e^{-\ell n/2}\,\}\ll e^{-\varepsilon n}.$
On the other hand, up to taking smaller $\varepsilon$, Lemma 2.1(i) guarantees that
(10) $\sigma\{\,s\,:\,|s|>e^{\ell n/4}\,\}\ll e^{-\varepsilon n}.$
The claim follows from the combination of (2.2), (9), (10). ∎
We now deduce our claim on the measures $\sigma^{(n)}$.
Lemma 2.3 (Moment and Hölder-regularity of $\sigma^{(n)}$).
There exists $\gamma>0$ such that
$(i)\quad\sup_{n\geq 1}\int_{\mathbb{R}}|s|^{\gamma}\,\mathrm{d}\sigma^{(n)}(s)<\infty$
and
$(ii)\quad\forall n\geq 1,\,\forall r>e^{-n},\quad\sup_{s\in\mathbb{R}}\sigma^{(n)}(s+[-r,r])\ll r^{\gamma}.$
Proof.
Fix $\gamma,\varepsilon\in(0,1)$ as in Lemma 2.2.
For (i), we need to check that the map $t\mapsto\sup_{n}\sigma^{(n)}\{\,s:|s|\geq t\,\}$ has polynomial decay as $t\to+\infty$. Given $t>2$, Lemma 2.2 and Lemma 2.1(i) imply that for every $n\geq 0$, one has
$\sigma^{(n)}\{\,s\,:\,|s|\geq t\,\}\leq\sigma\{\,s\,:\,|s|\geq t-1\,\}+O(e^{-\varepsilon n})\ll t^{-\gamma}+e^{-\frac{\varepsilon}{2}n}.$
Let $R>0$ be a parameter. The above justifies that uniformly in $n$, we have polynomial decays of tail probabilities of $\sigma^{(n)}$ for $t\leq e^{Rn}$. Taking $R\ggg 1$, the exponential moment assumption on $\lambda$ takes care of the case $t>e^{Rn}$, using the observation
$\sigma^{(n)}\{\,s\,:\,|s|\geq t\,\}\leq\lambda^{\otimes n}\Bigl{\{}(\phi_{1},\dots,\phi_{n})\,:\,n\prod_{k=1}^{n}\max(1,\mathtt{r}_{\phi_{k}},|\mathtt{b}_{\phi_{k}}|)\geq t\Bigr{\}}$
and the Markov inequality. This justifies (i), with a potentially smaller value of $\gamma$.
Let us check (ii). For $n\geq 0$, $s\in\mathbb{R}$ and $r\geq e^{-\frac{\varepsilon}{2}n}$, Lemma 2.2 guarantees
$\sigma^{(n)}([s-r,s+r])\leq\sigma([s-2r,s+2r])+O(e^{-\frac{\varepsilon}{2}n})\ll r^{\gamma},$
whence the claim (with $\frac{\varepsilon}{2}\gamma$ in place of $\gamma$ to treat all scales above $e^{-n}$). ∎
2.3. Recurrence of the random walk
We recall the following result of non-escape of mass for the $\mu$-walk on $X$.
Proposition 2.4 (Effective recurrence on $X$).
There exist constants $c,c^{\prime}>0$ depending on $\mu$ only such that for every $x\in X$, $n\in\mathbb{N}$, and $\rho>0$, we have
$\mu^{*n}*\delta_{x}\{\operatorname{inj}<\rho\}\ll({\operatorname{inj}^{-c}(x)}e^{-c^{\prime}n}+1)\rho^{c}.$
This result is proven in [34] in a more general setting. In our context, we present a brief proof for completeness.
Proof of Proposition 2.4.
[…]
(11) $\mu^{*n}*\delta_{x}(\operatorname{sys}^{-c})\ll e^{-c^{\prime}n}\operatorname{sys}(x)^{-c}+1.$
Let $k,R\geq 2$, $c\in(0,1)$ be parameters to be specified later depending on $\mu$. Recall also the notation $\ell=-\int_{G}\log\mathtt{r}_{g}\,\mathrm{d}\mu(g)>0$. Given $x\in X$, we write $w_{x}=v_{x}/\|v_{x}\|$ and set
$E:=\{g\in G\,:\,\|g\|>R^{k}\}\quad\text{and}\quad F_{x}:=\{g\in G\,:\,\|gw_{x}\|<e^{k\ell/4}\}.$
Assume $\operatorname{sys}(x)<R^{-k}$. Note that for any $g\in G$, we have $\operatorname{sys}(gx)\geq\|g\|^{-1}\operatorname{sys}(x)$, and also $\operatorname{sys}(gx)=\|gv_{x}\|\geq e^{k\ell/4}\operatorname{sys}(x)$ if $g\notin(E\cup F_{x})$. It follows that
(12) $\mu^{*k}*\delta_{x}(\operatorname{sys}^{-c})$ $\leq\int_{E\cup F_{x}}\|g\|^{c}\,\mathrm{d}\mu^{*k}(g)\operatorname{sys}(x)^{-c}+e^{-k\ell c/4}\operatorname{sys}(x)^{-c}.$
Moreover, using the Cauchy-Schwarz inequality and the submultiplicativity of the norm, we have
(13) $\int_{E\cup F_{x}}\|g\|^{c}\,\mathrm{d}\mu^{*k}(g)\leq\mu^{*k}(E\cup F_{x})^{1/2}M_{2c}(\mu)^{k/2}$
where, given $s>0$, we write $M_{s}(\mu):=\int_{G}\|g\|^{s}\,\mathrm{d}\mu(g)$ the $s$-exponential moment of $\mu$.
We claim that for some $\alpha=\alpha(\mu)>0$, and up to taking parameters $k,R\ggg 1$, we may further bound
(14) $\mu^{*k}(E\cup F_{x})\leq e^{-k\alpha}.$
Indeed, on the one hand, by the Markov inequality, for any $\varepsilon>0$,
$\mu^{*k}(E)\leq\left(M_{\varepsilon}(\mu)R^{-\varepsilon}\right)^{k}$
whence the claim on $\mu^{*k}(E)$ provided $R\ggg 1$. On the other hand, writing $w_{x}=(w_{x,1},w_{x,2})\in\mathbb{R}^{2}$, we have by looking at the first coordinate,
$\|gw_{x}\|\geq\mathtt{r}^{-1/2}_{g}|w_{x,1}+\mathtt{b}_{g}w_{x,2}|$
whence
$\mu^{*k}(F_{x})$ $\leq\mu^{*k}\{g\in G\,:\,|w_{x,1}+\mathtt{b}_{g}w_{x,2}|<\mathtt{r}^{1/2}_{g}e^{k\ell/4}\}$
$\leq\mu^{*k}\{g\in G\,:\,|w_{x,1}+\mathtt{b}_{g}w_{x,2}|<e^{-k\ell/5}\}+Ce^{-k\varepsilon},$
where $C,\varepsilon>0$ are constants depending on $\mu$ arising from the large devation principle for $\log\mathtt{r}_{g}$. Now the desired bound on $\mu^{*k}(F_{x})$ follows provided $k\ggg 1$: when $|w_{x,2}|\leq e^{-k\ell/10}$ it is a consequence of the moment estimate on $\mathtt{b}_{g}$, see Lemma 2.3(i), and otherwise it stems from its non-concentration estimate, see Lemma 2.3(ii). This finishes the proof of (14).
Combining (13) and (14), we get
(15) $\int_{E\cup F_{x}}\|g\|^{c}\,\mathrm{d}\mu^{*k}(g)\leq e^{-k\alpha/2}M_{2c}(\mu)^{k/2}\leq e^{-k\alpha/3},$
where the second inequality assumes additionally $c\lll 1$. Plugging (15) into (12), we have shown in the end that for $k,R\ggg 1$, and $c\lll 1$, for any $x\in X$ with $\operatorname{sys}(x)<R^{-k}$, we have
(16) $\mu^{*k}*\delta_{x}(\operatorname{sys}^{-c})\leq r\operatorname{sys}(x)^{-c}$
where $r=r(\mu,k,c)\in(0,1)$. On the other hand, if $\operatorname{sys}(x)\geq R^{-k}$ we have by finite exponential moment of $\mu$ and the inequality $\operatorname{sys}(gx)\geq\|g\|^{-1}\operatorname{sys}(x)$ that
(17) $\mu^{*k}*\delta_{x}(\operatorname{sys}^{-c})\leq C$
where $C=C(\mu,R,k,c)>0$.
Together (16) and (17) yield for every $x\in X$,
(18) $\mu^{*k}*\delta_{x}(\operatorname{sys}^{-c})\leq r\operatorname{sys}(x)^{-c}+C.$
The claim (11) follows from (18) inductively, setting $c^{\prime}=|\log r|$ and using that the sum $\sum_{k\geq 0}r^{k}C$ is finite. This finishes the proof of the proposition in the case where $\Lambda=\operatorname{SL}_{2}(\mathbb{Z})$.
For general $\Lambda$, note we may assume that $\Lambda$ is not cocompact, note also that each cusp of $X$ has a neighborhood that is isometric to a neighborhood of the cusp in $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$. Hence the behavior of the walk outside of a compact set is captured by the behavior in $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$, more precisely the proof of (11) (with $\operatorname{inj}$ in place of $\operatorname{sys}$ here) still goes through. ∎
3. Positive dimension
We show that the $n$-step distribution of the $\mu$-walk starting from a point $x$ acquires positive dimension at an exponential rate, tempered by the possibility that $x$ may be high in the cusp.
Proposition 3.1 (Positive dimension).
There exists $A,\kappa>0$ such that for every $x\in X$, $\rho>0$, $n\geq|\log\rho|+A|\log\operatorname{inj}(x)|$, we have
(19) $\forall y\in X,\quad\mu^{*n}*\delta_{x}(B_{\rho}y)\ll\rho^{\kappa}.$
Proof.
Let $\kappa>0$ be a parameter to specify below. Let $\rho\in(0,1/10)$, $n\geq|\log\rho|$, $x,y\in X$, and assume
(20) $\mu^{*n}*\delta_{x}(B_{\rho}y)\geq\rho^{\kappa}.$
Fix $\alpha=\alpha(\mu)>0$ so small that, setting $m=\lfloor\alpha|\log\rho|\rfloor$, we have
$\max_{h\in\operatorname{supp}\mu}\lVert\operatorname{Ad}(h)\rVert^{m}\leq\rho^{-1/2}.$
Writing $\mu^{*n}*\delta_{x}=\mu^{*m}*\mu^{*(n-m)}*\delta_{x}$, Equation (20) implies that
$\mu^{*(n-m)}*\delta_{x}(Z)\geq\rho^{2\kappa}\,\,\text{ where }\,\,Z:=\{z:\mu^{*m}*\delta_{z}(B_{\rho}y)\geq\rho^{2\kappa}\},$
up to assuming $\rho$ small enough in terms of $\kappa$. Indeed,
$\rho^{\kappa}\leq\mu^{*m}*\mu^{*(n-m)}*\delta_{x}(B_{\rho}y)$ $=\int_{Z\cup(X\smallsetminus Z)}\mu^{*m}*\delta_{z}(B_{\rho}y)\,\mathrm{d}\mu^{*(n-m)}*\delta_{x}(z)$
$\leq\rho^{2\kappa}+\mu^{*(n-m)}*\delta_{x}(Z),$
so we obtain $\mu^{*(n-m)}*\delta_{x}(Z)\geq\rho^{\kappa}-\rho^{2\kappa}\geq\rho^{2\kappa}$, provided $\rho\leq 2^{-1/\kappa}$.
We now show that $Z$ must be included in a small neighborhood of the cusp. Fix $z\in Z$. By the choice of $m$ and the lower bound $\mu^{*m}*\delta_{z}(B_{\rho}y)\geq\rho^{2\kappa}$, we have
(21) ${\mu^{*m}}\{\,g\,:\,g^{-1}y\in B_{\rho^{1/2}}z\,\}\geq\rho^{2\kappa}.$
On the other hand, by Lemma 2.3(i), we have for some $\gamma=\gamma(\mu)>0$, and $\rho\lll 1$,
(22) ${\mu^{*m}}\{\,g\,:\,|\mathtt{b}_{g}|\leq\rho^{-4\gamma^{-1}\kappa}\,\}\geq 1-\rho^{3\kappa}.$
By finite exponential moment of $\mu$, there exists also $\varepsilon>0$ depending only on $\mu$ such that
(23) $\mu^{*m}\{g\,:\,\log\mathtt{r}_{g}\in[(\ell-1)m,(\ell+1)m]\}\geq 1-\rho^{\alpha\varepsilon}.$
Let $C>1$ be a parameter to be specified below depending on $\mu$ only. Cutting the intervals $[-\rho^{-4\gamma^{-1}\kappa},\rho^{-4\gamma^{-1}\kappa}]$ and $[(\ell-1)m,(\ell+1)m]$ into subintervals of length $\rho^{C\kappa}$, then using the pigeonhole principle, we deduce from (21) (22), (23) that there exists $(r_{0},b_{0})\in\mathbb{R}^{2}$ with $r_{0}>0$, $|b_{0}|\leq\rho^{-4\gamma^{-1}\kappa}$ such that the set
$E:=\{g\,:\,g^{-1}y\in B_{\rho^{1/2}}z\text{ and }|1-\mathtt{r}_{g}r^{-1}_{0}|\leq\rho^{C\kappa}\text{ and }|\mathtt{b}_{g}-b_{0}|\leq\rho^{C\kappa}\}$
has $\mu^{*m}$-measure
(24) $\mu^{*m}(E)\geq\frac{\rho^{2\kappa}-\rho^{3\kappa}-\rho^{\alpha\varepsilon}}{\lceil 2\rho^{-4\gamma^{-1}\kappa}\rho^{-C\kappa}\rceil\lceil 2m\rho^{-C\kappa}\rceil}\geq\rho^{4C\kappa}$
where the last lower bound assumes $C\geq 4\gamma^{-1}$, $3\kappa\leq\alpha\varepsilon$, and $\rho\lll_{\kappa}1$.
Consider $g_{1},g_{2}\in E$, in particular $\operatorname{dist}(g_{1}^{-1}y,g_{2}^{-1}y)\ll\rho^{1/2}$. Recalling $g^{-1}_{i}=u(-\mathtt{b}_{g_{i}})a(\mathtt{r}_{g_{i}})$ and setting $z_{2}=a(\mathtt{r}_{g_{2}})y$, the inequality can be rewritten
$\operatorname{dist}(u(-\mathtt{b}_{g_{1}})a(\mathtt{r}_{g_{1}}\mathtt{r}^{-1}_{g_{2}})z_{2},u(-\mathtt{b}_{g_{2}})z_{2})\ll\rho^{1/2}.$
In view of the condition $|\mathtt{b}_{g_{2}}|\leq\rho^{-4\gamma^{-1}\kappa}$, this implies
(25) $\operatorname{dist}(u(\mathtt{b}_{g_{2}}-\mathtt{b}_{g_{1}})a(\mathtt{r}_{g_{1}}\mathtt{r}^{-1}_{g_{2}})z_{2},z_{2})\ll\rho^{1/2-O(\kappa)}\ll\rho^{1/4}$
provided $\kappa\lll 1$. Moreover, the combination of (24) and the non-concentration estimate from Lemma 2.3(ii) allows to choose the elements $g_{1},g_{2}\in E$ such that
(26) $|\mathtt{b}_{g_{2}}-\mathtt{b}_{g_{1}}|\geq\rho^{\gamma^{-1}5C\kappa},$
provided $\kappa\lll_{C}1$ to guarantee $\rho^{\gamma^{-1}5C\kappa}\geq e^{-m}$ as required by Lemma 2.3(ii). Observing that
$\operatorname{dist}(u(\mathtt{b}_{g_{2}}-\mathtt{b}_{g_{1}})a(\mathtt{r}_{g_{1}}\mathtt{r}^{-1}_{g_{2}}),\operatorname{Id})$ $\simeq|\mathtt{b}_{g_{2}}-\mathtt{b}_{g_{1}}|+|1-\mathtt{r}_{g_{1}}\mathtt{r}^{-1}_{g_{2}}|$ $\in[\rho^{\gamma^{-1}5C\kappa},4\rho^{C\kappa}]$
we deduce from (25) that for $\kappa\lll_{C}1$ so that $\rho^{\gamma^{-1}5C\kappa}\geq\rho^{1/5}$, we have
$\operatorname{inj}(z_{2})\ll\rho^{C\kappa}.$
Recalling $\operatorname{dist}(z,z_{2})\leq\operatorname{dist}(z,g_{2}^{-1}y)+\operatorname{dist}(g_{2}^{-1}y,z_{2})\ll\rho^{-4\gamma^{-1}\kappa}$, we get
$\operatorname{inj}(z)\ll\rho^{\frac{C}{2}\kappa}$
provided $C\ggg 1$.
In conclusion, we have shown that for $C\ggg 1$, for $\kappa\lll_{C}1$, for $\rho\lll_{\kappa}1$, and $n\geq m\simeq\alpha|\log\rho|$, we have
${\mu^{*(n-m)}}*\delta_{x}\{\operatorname{inj}\leq\rho^{C\kappa}\}\geq\rho^{2\kappa}$
By the effective recurrence statement from Proposition 2.4, this is absurd if $n-m\ggg{|\log\operatorname{inj}(x)|}$. This concludes the proof of the proposition. ∎
4. Dimensional bootstrap
In this section, we explain how the positive dimension estimate for $\mu^{*n}*\delta_{x}$ established in the previous section can be upgraded to a high-dimension estimate, up to applying more convolutions by $\mu$ and throwing away some small part of the measure. We will need the following definition.
Definition 4.1 (Robustness).
Let $\alpha,\tau>0$, $I\subseteq(0,1]$. A Borel measure $\nu$ on $X$ is $(\alpha,{\mathcal{B}}_{I},\tau)$-robust if $\nu$ can be decomposed as the sum of two Borel measures $\nu=\nu^{\prime}+\nu^{\prime\prime}$ such that $\nu^{\prime\prime}(X)\leq\tau$, and $\nu^{\prime}$ satisfies
(27) $\nu^{\prime}\{\operatorname{inj}<\sup I\}=0,$
as well as for all $\rho\in I$, $y\in X$,
(28) $\nu^{\prime}(B_{\rho}y)\leq\rho^{3\alpha}$
Condition (28) means that $\nu^{\prime}$ has normalised dimension at least $\alpha$ with respect to balls of radius $\rho$. Note that in this definition, $\nu$ may not be a probability measure, this flexibility will be convenient for us.
The goal of the section is to establish the following high dimension estimate.
Proposition 4.1 (High dimension).
Let $\kappa\in(0,1/10)$. For $\eta,\rho\lll_{\kappa}1$ and for all $n\ggg_{\kappa}|\log\rho|+|\log\operatorname{inj}(x)|$, the measure $\mu^{*n}*\delta_{x}$ is $(1-\kappa,B_{\rho},\rho^{\eta})$-robust.
4.1. Multislicing
The proof of Proposition 4.1 relies on a multislicing estimate established in [4]. We recall the case of interest in our context.
We consider $\Theta$ a measurable space. We let $(\varphi_{\theta})_{\theta\in\Theta}$ denote a measurable family of $C^{2}$-embeddings $\varphi_{\theta}:B^{\mathbb{R}^{3}}_{1}\rightarrow\mathbb{R}^{3}$, and $(L_{\theta})_{\theta\in\Theta}$ a measurable family of constants $L_{\theta}\geq 1$ such that each map $\varphi_{\theta}$ is $L_{\theta}$-bi-Lipschitz:
$\forall x,y\in B^{\mathbb{R}^{3}}_{1},\quad\frac{1}{L_{\theta}}\|x-y\|\leq\|\varphi_{\theta}(x)-\varphi_{\theta}(y)\|\leq L_{\theta}\|x-y\|$
and has second order derivatives bounded by $L_{\theta}$:
$\forall x,h\in B^{\mathbb{R}^{3}}_{1},\quad\lVert\varphi_{\theta}(x+h)-\varphi_{\theta}(x)-(D_{x}\varphi_{\theta})(h)\rVert\leq L_{\theta}\lVert h\rVert^{2}.$
Given $\rho>0$, we denote by ${\mathcal{D}}_{\rho}$ (resp. ${\mathcal{R}}_{\rho}$) the collection of subsets of $\mathbb{R}^{3}$ that are translates of the $\rho$-cube $[0,\rho]^{3}$ (resp. the rectangle $R_{\rho}:=[0,1]e_{1}+[0,\rho^{1/2}]e_{2}+[0,\rho]e_{3}$).
We will also need to measure the angle between subspaces in $\mathbb{R}^{3}$. For each $k=1,2,3$, endow $\wedge^{k}\mathbb{R}^{3}$ with the unique Euclidean with respect to which the standard basis is orthonormal. Given subspaces $V,W\subseteq\mathbb{R}^{3}$, we set
$\operatorname{d_{\measuredangle}}(V,W)=\|v\wedge w\|$
where $v,w$ are unit vectors in $\wedge^{*}\mathbb{R}^{3}$ spanning respectively the lines $\wedge^{\dim V}V$, $\wedge^{\dim W}W$.
The multislicing estimate presented in 4.2 below takes as input a Borel measure $\nu$ on the unit ball $B^{\mathbb{R}^{3}}_{1}$ that has normalized dimension at least $\alpha$ with respect to balls of radius above $\rho$. The output is a dimensional gain when the balls are replaced by (non-linear) rectangles of the form $(\varphi^{-1}_{\theta}(x+R_{\rho}))_{x\in\mathbb{R}^{3}}$ provided $\theta$ is chosen almost typically via a probability measure for which $\varphi_{\theta}$ satisfies suitable bounds on the derivatives as well as non-concentration estimates.
Theorem 4.2 (Multislicing [4]).
Given $\kappa\in(0,1/2)$, there exist $\varepsilon=\varepsilon(\kappa)>0$ and $\rho_{0}=\rho_{0}(\kappa)>0$ such that the following holds for all $\rho\in(0,\rho_{0}]$.
Let $\nu$ be a Borel measure on $B^{\mathbb{R}^{3}}_{1}$ satisfying: $\exists\alpha\in(\kappa,1-\kappa)$, $\forall r\in[\rho,\rho^{\varepsilon}]$,
$\sup_{Q\in{\mathcal{D}}_{r}}\nu(Q)\leq r^{3\alpha}.$
Let $\Xi$ be a probability measure on $\Theta$ satisfying:
  • (i)
    $\Xi\{\,\theta\in\Theta:L_{\theta}\leq\rho^{-\varepsilon}\,\}=1.$
  • (ii)
    $\forall k\in\{1,2\}$, $\forall x\in B^{\mathbb{R}^{3}}_{1}$, $\forall r\in[\rho,\rho^{\varepsilon}]$, $\forall W\in\operatorname{Gr}(\mathbb{R}^{3},3-k)$,
    $\Xi\{\,\theta\in\Theta:\operatorname{d_{\measuredangle}}((D_{x}\varphi_{\theta})^{-1}V_{k},W)\leq r\,\}\leq r^{\kappa},$
    where $V_{k}=\operatorname{span}_{\mathbb{R}}(e_{1},\dotsc,e_{k})$.
Then there exists ${\mathcal{F}}\subseteq\Theta$ such that $\sigma({\mathcal{F}})\geq 1-\rho^{\varepsilon}$ and for every $\theta\in{\mathcal{F}}$, there exists $A_{\theta}\subseteq B^{\mathbb{R}^{3}}_{1}$ with $\nu(A_{\theta})\geq 1-\rho^{\varepsilon}$ and satisfying
$\sup_{Q\in{\mathcal{R}}_{\rho}}\nu_{|A_{\theta}}(\varphi_{\theta}^{-1}Q)\leq\rho^{\frac{3}{2}\alpha+\varepsilon}.$
Proof.
This is a special case of [4, Theorem 2.2]. ∎
4.2. Straightening charts
In order to apply the multislicing estimates from 4.2, we need special macroscopic charts in which the preimage by $g\in\operatorname{supp}\mu^{*n}$ of a ball looks like a rectangle. The goal of the present subsection is to define those charts.
Recall that $\mathfrak{g}$ admits the rootspace decomposition
$\mathfrak{g}=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{+},$
where
$\mathfrak{g}_{-}=\mathbb{R}e_{-},\quad\mathfrak{g}_{0}=\mathbb{R}e_{0}\quad\text{and}\quad\mathfrak{g}_{+}=\mathbb{R}e_{+}.$
We then define $\Psi:\mathfrak{g}\rightarrow G$ by the formula: $\forall(v_{-},v_{0},v_{+})\in\mathfrak{g}_{-}\times{\mathfrak{g}_{0}}\times\mathfrak{g}_{+}$,
$\Psi(v_{-}+v_{0}+v_{+})=\exp(v_{-})\exp(v_{0})\exp(v_{+}).$
Recall also the notation
$a(t)=\begin{pmatrix}t^{1/2}&0\\ 0&t^{-1/2}\end{pmatrix}.$
The next lemma tells us that, in the chart $\Psi$, the image of a ball $B_{\rho}$ in $G$ by some diagonal element $a(t)$ with small $t>0$ is included in a rectangle whose volume is comparable.
Lemma 4.3.
There is an absolute constant $r_{0}>0$ such that for any $t,\rho\in(0,1)$ with $|t^{-1}\rho|\leq r_{0}$ and any $h\in G$, there is $w\in\mathfrak{g}$ such that
(29) $\{\,v\in B^{\mathfrak{g}}_{r_{0}}:\Psi(v)\in a(t)B_{\rho}h\,\}\subseteq\operatorname{Ad}(a(t))B^{\mathfrak{g}}_{10\rho}+w.$
This result is a particular case of [4, Lemma 4.10]. We give a shorter proof in our context for completeness.
Proof.
Fix a vector $w$ in the left hand side of (29). If $v$ belongs to the left hand side of (29) as well, then by the triangle inequality, we have
$\Psi(v)\in a(t)B_{2\rho}a(t^{-1})\Psi(w).$
We can choose $r_{0}>0$ small so that the image of $\Psi$ contains $B_{2r_{0}}$. In particular, using that conjugation commutes with the exponential map, there is $u=u_{-}+u_{0}+u_{+}\in\mathfrak{g}$ such that
$u_{-}\in B^{\mathfrak{g}_{-}}_{t^{-1}\rho},\quad u_{0}\in B^{\mathfrak{g}_{0}}_{\rho},\quad u_{+}\in B^{\mathfrak{g}_{+}}_{t\rho}$
and
$\Psi(v)=\Psi(u)\Psi(w).$
Consider $x,y,s\in\mathbb{R}$ with $s\neq 0$ and $1+xy\neq 0$. Note that we have in $G$ the following equality
(30) $\begin{pmatrix}1&x\\ 0&1\end{pmatrix}\begin{pmatrix}1&0\\ y&1\end{pmatrix}\begin{pmatrix}s&0\\ 0&s^{-1}\end{pmatrix}=\begin{pmatrix}1&0\\ y^{\prime}&1\end{pmatrix}\begin{pmatrix}s^{\prime}&0\\ 0&s^{\prime-1}\end{pmatrix}\begin{pmatrix}1&x^{\prime}\\ 0&1\end{pmatrix}$
where
$x^{\prime}=\frac{x}{(1+xy)s^{2}},\quad s^{\prime}=(1+xy)s,\quad y^{\prime}=\frac{y}{1+xy}.$
Similarly,
(31) $\begin{pmatrix}s&0\\ 0&s^{-1}\end{pmatrix}\begin{pmatrix}1&0\\ y&1\end{pmatrix}=\begin{pmatrix}1&0\\ s^{-2}y&1\end{pmatrix}\begin{pmatrix}s&0\\ 0&s^{-1}\end{pmatrix}.$
Observe that
$\Psi(v)=\Psi(u)\Psi(w)=\exp(u_{-})\exp(u_{0})\exp(u_{+})\exp(w_{-})\exp(w_{0})\exp(w_{+}).$
Assuming $r_{0}\lll 1$, applying (30) to the factor $\exp(u_{+})\exp(w_{-})\exp(w_{0})$ then (31) to the factor $\exp(u_{0})\exp(w^{\prime}_{-})$, we obtain
$\Psi(v)\in\exp(B^{\mathfrak{g}_{-}}_{10t^{-1}\rho}+w_{-})\exp(B^{\mathfrak{g}_{0}}_{10\rho}+w_{0})\exp(B^{\mathfrak{g}_{+}}_{10t\rho}+w_{+}).$
Noting that $\Psi$ is injective (by direct computation again), this finishes the proof. ∎
In view of Lemma 4.3 and the formula
$g=a(\mathtt{r}^{-1}_{g})u(\mathtt{b}_{g}),$
we define a family of straightening charts $(\varphi_{\theta})$ as follows. Let
$\Theta=u(\mathbb{R}).$
Fix $r_{1}>0$ such that $\Psi$ is a smooth diffeomorphism between $B^{\mathfrak{g}}_{r_{1}}$ and a neighborhood $\mathcal{O}$ of $\operatorname{Id}\in G$. Given $\theta\in\Theta$, define $\varphi_{\theta}:\mathcal{O}\rightarrow\mathfrak{g}$ by
$\varphi_{\theta}:=\operatorname{Ad}(\theta^{-1})\circ(\Psi_{|B^{\mathfrak{g}}_{r_{1}}})^{-1}.$
Using that $\psi$ commutes with conjugation, we have the alternative formula $\varphi_{\theta}=(\Psi_{|\operatorname{Ad}(\theta^{-1})B^{\mathfrak{g}}_{r_{1}}})^{-1}\circ{\mathscr{C}}_{\theta^{-1}}$ where ${\mathscr{C}}_{\theta^{-1}}:h\mapsto\theta^{-1}h\theta$. Note that $\varphi_{\theta}$ is $L_{\theta}$-bi-Lipschitz and satisfies $\|\varphi_{\theta}\|_{C^{2}}\leq L_{\theta}$ for some quantity
(32) $L_{\theta}:=L\|\theta\|^{4}.$
where $L>1$ is a constant depending only on $r_{1}$.
Given an element $g\in P$, write
$g^{-1}=\theta_{g}a(\mathtt{r}_{g})$
with
(33) $\theta_{g}:=u(-\mathtt{b}_{g})\in\Theta.$
Lemma 4.3 tells us that for any $h\in G$, $\varphi_{\theta_{g}}(g^{-1}B_{\rho}h)$ is essentially an additive translate of the rectangle $\operatorname{Ad}(a(\mathtt{r}_{g}))B^{\mathfrak{g}}_{\rho}$ provided that $\mathtt{r}_{g}\in(0,1)$, and that both $g^{-1}B_{\rho}h$ and $a(\mathtt{r}_{g})B_{\rho}h\theta_{g}$ sit inside a prescribed (macroscopic) neighborhood of the identity.
4.3. Control of the charts
We now check that the charts ($\varphi_{\theta}$) from the previous section satisfy distortion control and non-concentration estimates. These will be required in order to apply 4.2 in the next section. The constants $r_{0}$ from Lemma 4.3 and $r_{1}>0$ in the definition of $\varphi_{\theta}$ are assumed fixed in a canonical way (so that dependence on them does not appear in subscript asymptotic notations).
Recall $L_{\theta}$ and $\theta_{g}$ are respectively defined in (32) and (33).
Lemma 4.4 (Distortion control).
Given $\varepsilon>0$, there exists $\gamma=\gamma(\mu,\varepsilon)>0$ such that for $n\ggg_{\varepsilon}1$, we have
$\mu^{*n}\{\,g\,:\,L_{\theta_{g}}>e^{\varepsilon n}\,\}\leq e^{-\gamma n}.$
Proof.
This is a direct consequence of $L_{\theta_{g}}\ll\|\theta_{g}\|^{4}\ll(1+|\mathtt{b}_{g}|)^{4}$ and Lemma 2.3(i), stating that the variable $\mathtt{b}_{g}$, where $g\overset{law}{\sim}\mu^{*n}$, has a moment of positive order that is bounded independently of $n$. ∎
Set $\mathfrak{g}_{-,0}=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}$.
Lemma 4.5 (Non-concentration).
There exists a constant $\kappa>0$ such that for $n\ggg 1$, $h\in B_{r_{0}}$ and $\rho\geq e^{-n}$, we have
$\forall W\in\operatorname{Gr}(T_{h}G,2),\quad\mu^{*n}\{\,g\,:\,\operatorname{d_{\measuredangle}}((D_{h}\varphi_{\theta_{g}})^{-1}\mathfrak{g}_{-},W)\leq\rho\,\}\ll\rho^{\kappa},$
and
$\forall W\in\operatorname{Gr}(T_{h}G,1),\quad\mu^{*n}\{\,g\,:\,\operatorname{d_{\measuredangle}}((D_{h}\varphi_{\theta_{g}})^{-1}\mathfrak{g}_{-,0},W)\leq\rho\,\}\ll\rho^{\kappa}.$
Proof.
Unwrapping definitions, we observe that the distribution of subspaces $h\mapsto(D_{h}\varphi_{\theta})^{-1}\mathfrak{g}_{-}$ is right-invariant and coincides with $\operatorname{Ad}(\theta)\mathfrak{g}_{-}$ at the identity. The same holds for $(D_{h}\varphi_{\theta})^{-1}\mathfrak{g}_{-,0}$. Recalling that $\theta_{g}=u(-\mathtt{b}_{g})$ and $\sigma^{(n)}$ is the law of $\mathtt{b}_{g}$ as $g\overset{law}{\sim}\mu^{*n}$, we are then led to proving the following non-concentration estimates:
$\sup_{W\in\operatorname{Gr}(\mathfrak{g},2)}\sigma^{(n)}\{s\,:\,\operatorname{d_{\measuredangle}}(\operatorname{Ad}(u(-s))\mathfrak{g}_{-},W)\leq\rho\}\ll\rho^{\kappa}.$
$\sup_{W\in\operatorname{Gr}(\mathfrak{g},1)}\sigma^{(n)}\{s\,:\,\operatorname{d_{\measuredangle}}(\operatorname{Ad}(u(-s))\mathfrak{g}_{-,0},W)\leq\rho\}\ll\rho^{\kappa}.$
Let us check the first estimate, where $W\in\operatorname{Gr}(\mathfrak{g},2)$. Set $e_{-,0}=e_{-}\wedge e_{0}$, $e_{-,+}=e_{-}\wedge e_{+}$, $e_{0,+}=e_{0}\wedge e_{+}$. Write $\wedge^{2}W=\mathbb{R}(ae_{-,0}+be_{-,+}+ce_{0,+})$ where $a,b,c\in\mathbb{R}$ with $\max(|a|,|b|,|c|)=1$. Then direct computation yields for any $s\in\mathbb{R}$,
$\operatorname{d_{\measuredangle}}(\operatorname{Ad}(u(-s))\mathfrak{g}_{-},W)\simeq|as^{2}-bs-c|.$
In the case $\max(|a|,|b|)\leq\rho^{1/10}$, we must have $|c|=1$ while $|as^{2}-bs|\leq 1/2$ with $\sigma^{(n)}$-probability at least $1-O(\rho^{\gamma^{-1}/20})$ for some $\gamma=\gamma(\mu)$, see (Lemma 2.3(i)), whence the claim. In the case $\max(|a|,|b|)\geq\rho^{1/10}$, the desired non-concentration follows from combining Lemma 4.6 below and Lemma 2.3 (items (i) and (ii)), the latter claiming that $\sigma^{(n)}$ has finite moment and positive dimension at scales above $e^{-n}$, uniformly in $n$.
The second estimate is similar: writing $W=\mathbb{R}(ae_{+}+be_{0}+ce_{-})$ where $\max(\lvert a\rvert,\lvert b\rvert,\lvert c\rvert)=1$, we find $\operatorname{d_{\measuredangle}}(\operatorname{Ad}(u(-s))\mathfrak{g}_{-},W)\simeq\lvert a-2bs-cs^{2}\rvert$. ∎
The next elementary lemma was used in the proof Lemma 4.5 of above.
Lemma 4.6.
Let $(a,b,c)\in\mathbb{R}^{3}$ and $\varepsilon\in(0,1)$ with $\max(|a|,|b|)\geq\varepsilon$, let $\rho\in(0,1)$. The set $E:=\{s\in[-\rho^{-1/4},\rho^{-1/4}]\,:\,|as^{2}+bs+c|\leq\rho\}$ is included in at most two balls of radius $8\varepsilon^{-1}\rho^{1/4}$.
Proof.
We may assume $8\varepsilon^{-1}\rho^{1/4}\leq 2\rho^{-1/4}$, i.e. $\rho^{1/2}\leq\varepsilon/4$. If $s_{1},s_{2}\in E$, then $|as_{1}^{2}+bs_{1}-as_{2}^{2}+bs_{2}|\leq 2\rho$, i.e.
$|(s_{1}-s_{2})(b+a(s_{1}+s_{2}))|\leq 2\rho.$
Then either $|s_{1}-s_{2}|\leq 2\rho^{1/2}$ or $|b+a(s_{1}+s_{2})|\leq\rho^{1/2}$. In the second case, the conditions $\max(|a|,|b|)\geq\varepsilon$ and $\rho^{1/2}\leq\varepsilon/4$ force $a\geq\rho^{1/4}\varepsilon/4$, then $s_{1}$ belongs to the ball of radius $8\varepsilon^{-1}\rho^{1/4}$ and center $(-a^{-1}b-s_{2})$. ∎
4.4. Dimension increment
In this subsection, we apply the multislicing estimate from 4.2 to show that convolution by a well chosen power of $\mu$ increases dimensional properties of a measure at a given scale.
Proposition 4.7 (Dimension increment).
Let $\kappa,\varepsilon,\rho\in(0,1/10)$, $\alpha\in{[\kappa,1-\kappa]}$, $\tau\geq 0$ be some parameters. Consider on $X$ a Borel measure $\nu$ which is $(\alpha,{\mathcal{B}}_{[\rho,\rho^{\varepsilon}]},\tau)$-robust. Denote by $n_{\rho}\geq 0$ the integer part of $\frac{1}{2\ell}|\log\rho|$.
Assume $\varepsilon,\rho\lll_{\kappa}1$, then
[…]
Remark.
Recall here that $\ell$ denotes the Lyapunov exponent of the $\operatorname{Ad}_{\star}\mu$-walk on $\mathfrak{g}$. Hence our choice for $n_{\rho}$ guarantees that the operator norm $\|\operatorname{Ad}g^{-1}\|$ is roughly $\rho^{-1/2}$ when $g\overset{law}{\sim}{\mu^{*n_{\rho}}}$.
Proof.
In the proof, we may allow $\rho$ to be small enough in terms of $\varepsilon$ (not just $\kappa$). We may also assume $\tau=0$. We will write $n=n_{\rho}$.
Note that the compact set $X_{\rho^{\varepsilon}}:=\{\operatorname{inj}\geq\rho^{\varepsilon}\}$ can be covered by $\rho^{-O(\varepsilon)}$ balls of radius $\rho^{2\varepsilon}$, more precisely
$X_{\rho^{\varepsilon}}\subseteq\cup_{i\in I}B_{\rho^{2\varepsilon}}x_{i}$
where $\sharp I\leq\rho^{-O(\varepsilon)}$, $x_{i}\in X_{\rho^{\varepsilon}}$ for all $i$. As $\nu$ is supported on $X_{\rho^{\varepsilon}}$, we can then write
$\nu=\sum_{i\in I}\nu_{i}*\delta_{x_{i}}$
where $\nu_{i}$ is a Borel measure on $G$ with support in $B_{\rho^{2\varepsilon}}$. Note that the assumption that $\nu$ is $(\alpha,{\mathcal{B}}_{[\rho,\rho^{\varepsilon}]},0)$-robust implies that each $\nu_{i}$ is $(\alpha,{\mathcal{B}}_{[\rho,\rho^{\varepsilon}]},0)$-robust.
We now apply 4.2 to each $\nu_{i}$. We consider the family of charts $\varphi_{\theta}:\mathcal{U}\rightarrow\mathfrak{g}$ introduced in Section 4.3. In order to guarantee the distortion control requirement for $\varphi_{\theta}$, we introduce the renormalized truncation of $\mu^{*n}$ defined by
$\mu^{\prime}_{n}=\frac{\mu^{*n}_{|L_{\theta_{g}}\leq\rho^{-\varepsilon}}}{\mu^{*n}\{L_{\theta_{g}}\leq\rho^{-\varepsilon}\}}.$
By Lemma 4.4, this probability measure satisfies $\|\mu^{\prime}_{n}-\mu^{*n}\|\leq\rho^{\gamma}$ for some $\gamma=\gamma(\mu,\varepsilon)>0$. In particular, provided $\rho\lll_{\varepsilon}1$, the measure $\mu^{\prime}_{n}$ also satisfies the non-concentration estimates from Lemma 4.5. This allows us to apply 4.2 with $\Xi$ the law of $\theta_{g}$ when $g\overset{law}{\sim}{\mu^{\prime}_{n}}$ ($n=n_{\rho}$). We obtain some constant $\varepsilon_{1}>0$ depending only on $\kappa$, $\mu$ such that up to assuming $\varepsilon\lll_{\kappa}1$, $\rho\lll_{\kappa,\varepsilon}1$, there exists $\mathcal{G}_{i}\subseteq P$ with ${\mu^{\prime}_{n}}(\mathcal{G}_{i})\geq 1-\rho^{\varepsilon_{1}}$ satisfying for every $g\in\mathcal{G}_{i}$, that there exists a component measure $\nu_{i,g}\leq\nu_{i}$ with $\nu_{i,g}(G)\geq\nu_{i}(G)-\rho^{\varepsilon_{1}}$ and such that
(34) $\sup_{Q\in{\mathcal{R}}_{\rho}}\nu_{i,g}(\varphi_{\theta_{g}}^{-1}Q)\leq\rho^{\frac{3}{2}\alpha+\varepsilon_{1}}.$
On the other hand, the large deviation principle for the walk on $\mathbb{R}$ driven by $-\log\mathtt{r}_{g}\,\mathrm{d}\mu(g)$ guarantees that
the set $\mathcal{G}_{\mathtt{r}}=\{\,g\,:\,{\mathtt{r}_{g}^{-1}}\in[\rho^{-1/2+\varepsilon},\rho^{-1/2-\varepsilon}]\,\}\,\,$ satisfies $\,\,{\mu^{\prime}_{n}(\mathcal{G}_{\mathtt{r}})}\geq 1-\rho^{\varepsilon_{2}}$
for some $\varepsilon_{2}=\varepsilon_{2}(\mu,\varepsilon)>0$.
Setting $\mathcal{G}_{i,\mathtt{r}}=\mathcal{G}_{i}\cap\mathcal{G}_{\mathtt{r}}$ and using Lemma 4.3, observe that for $i\in I$, $g\in\mathcal{G}_{i,\mathtt{r}}$, for any ball $B_{\rho^{1/2}}y$ where $y\in X$, the intersection $(g^{-1}B_{\rho^{1/2}}y)\cap B_{\rho^{2\varepsilon}}x_{i}$ lifted to $B_{\rho^{2\varepsilon}}$ is included in at most $\rho^{-O(\varepsilon)}$ blocks of the form $\varphi_{\theta_{g}}^{-1}Q$ where $Q\in{\mathcal{R}}_{\rho}$. Hence, we get from (34),
(35) $\sup_{y\in X}\delta_{g}*\nu_{i,g}(B_{\rho^{1/2}}y)\leq\rho^{\frac{3}{2}\alpha+\varepsilon_{1}-O(\varepsilon)}.$
Setting $\mathcal{G}=\cap_{i\in I}\mathcal{G}_{i,\mathtt{r}}$ and recalling $\sharp I\leq\rho^{-O(\varepsilon)}$, we have $\mu^{\prime}_{n}(\mathcal{G})\geq 1-\rho^{\varepsilon_{1}-O(\varepsilon)}-\rho^{\varepsilon_{2}}$. We deduce
$\mu^{*n}(\{L_{\theta_{g}}\leq\rho^{-\varepsilon}\}\cap\mathcal{G})\geq 1-\rho^{\varepsilon_{1}-O(\varepsilon)}-\rho^{\varepsilon_{2}}-\rho^{\gamma}.$
Letting
$m_{n}^{\prime\prime}={\mu^{*n}}*\nu-\int_{G}\sum_{i}\delta_{g}*\nu_{i,g}\operatorname{{\bf 1}}_{\{L_{\theta_{g}}\leq\rho^{-\varepsilon}\}\cap\mathcal{G}}\,\mathrm{d}\mu^{*n}(g),$
and taking $\varepsilon\lll_{\varepsilon_{1}}1$, we have $\|m_{n}^{\prime\prime}\|\leq\rho^{\varepsilon_{3}}$ where $\varepsilon_{3}=\varepsilon_{3}(\varepsilon,\varepsilon_{1},\varepsilon_{2},\gamma)>0$, while we see from (35) that $m_{n}^{\prime}:=({\mu^{*n}}*\nu)-m_{n}^{\prime\prime}$ satisfies
$\sup_{y\in X}m^{\prime}_{n}(B_{\rho^{1/2}}y)\leq\rho^{\frac{3}{2}\alpha+\varepsilon_{1}/2}.$
We now have checked the required dimensional increment. In order to conclude, we also need check that $\mu^{*n}*\nu$ does not give too much mass to the cusp. This directly follows from Proposition 2.4 implying that for some constants $c,c^{\prime}>0$ depending on $\mu$, we have
$\mu^{*n}*\delta_{x}\{\operatorname{inj}<\rho^{1/2}\,\}\ll({\operatorname{inj}^{-c}(x)}e^{-c^{\prime}n}+1)\rho^{c/2},$
while by assumption, $\nu$ is supported on $\{\operatorname{inj}\geq\rho^{\varepsilon}\}$ with $\varepsilon<c^{\prime}/(2\ell c)$, and $n=n_{\rho}$. ∎
4.5. Proof of high dimension
We are finally able to show Proposition 4.1, namely that ${\mu^{*n}}*\delta_{x}$ reaches high dimension exponentially fast. The proof starts from positive dimension given by Proposition 3.1 and then proceeds by small increments using Proposition 4.7. Note however that Proposition 4.7 assumes non-concentration on a wide range of scales but the output dimensional increment only concerns a specific scale. Hence we need combine those single-scale increments to allow iterating the bootstrap. For this, we rely on the following lemma.
Lemma 4.8.
Let $\alpha,s,\rho\in(0,1]$, $\tau\in\mathbb{R}_{>0}$ be parameters. If $\nu$ is $(\alpha,\mathcal{B}_{r},\tau)$-robust for all $r\in[\rho,\rho^{s}]$, then for any $\varepsilon\in(0,\alpha)$, the measure $\nu$ is $(\alpha-\epsilon,\mathcal{B}_{[\rho,\rho^{s}]},\lceil\frac{\log s}{\log(1-\varepsilon)}\rceil\tau)$-robust.
Proof.
This is just a combination of two observations (1) if $\nu$ is $(\alpha,\mathcal{B}_{r},\tau)$-robust, then for every $t\in(0,1)$, it is $(t\alpha,\mathcal{B}_{[r^{1/t},r]},\tau)$-robust; (2) if $\nu$ is $(\alpha,\mathcal{B}_{I_{1}},\tau_{1})$-robust and $(\alpha,\mathcal{B}_{I_{2}},\tau_{2})$-robust, then $\nu$ is $(\alpha,\mathcal{B}_{I_{1}\cup I_{2}},\tau_{1}+\tau_{2})$-robust. See [4, Lemma 4.5] for details. ∎
Proof of Proposition 4.1.
Let $A>0$ be a large enough constant depending on the initial data $\mu$. Combining Proposition 3.1 and Proposition 2.4, we may assume $\kappa>0$ small enough from the start, so that for any $M>0$, for every $\rho\lll_{M}1$ and $n\geq M|\log\rho|+A|\log\operatorname{inj}(x)|$, the measure
[…]
By Proposition 4.7, there is some small constant $\varepsilon=\varepsilon(\mu,\kappa)>0$, such that up to imposing from the start $M\ggg_{\kappa}1$, we have for all $\rho\lll_{\kappa,M}1$, all $n\geq(\frac{1}{2\ell}+1)M|\log\rho|+A|\log\operatorname{inj}(x)|$ and $r\in[\rho^{2M\varepsilon},\rho^{1/(2M\varepsilon)}]$,
[…]
These estimates for single scales can be combined using Lemma 4.8 to get under the same conditions:
[…]
The argument in the last paragraph can be applied iteratively, adding at each step $k$ the value $+\varepsilon$ to the dimension provided the latter is not yet above $1-\kappa$ and provided $M\ggg_{\kappa,k}1$. As the value of $\varepsilon$ only depends on $\kappa$, we reach dimension $1-\kappa$ in a finite number of steps. This concludes the proof.
5. From high dimension to equidistribution
We consider the one-parameter family of probability measures $(\eta_{t})_{t>0}$ on $G$ given by
$\eta_{t}=a(t)u(s)\,\mathrm{d}\sigma(s).$
We show Proposition 5.1, stating that as $t\to+\infty$, a probability measure on $X$ with dimension close to $3$ equidistributes under the $\eta_{t}$-process toward the Haar measure on $X$, and does so with exponential rate. From this we deduce B’ (whence B) and C’ (whence C).
Proposition 5.1.
There exist $\kappa,\rho_{0}>0$ such that the following holds for all $\rho\in(0,\rho_{0}]$, $\tau>0$.
Let $\nu$ be a Borel measure on $X$ that is $(1-\kappa,{\mathcal{B}}_{\rho},\tau)$-robust and has mass at most $1$. Then for any $t\in[\rho^{-1/4},\rho^{-1/2}]$, for any $f\in B^{\infty}_{\infty,1}(X)$ with $m_{X}(f)=0$, we have
$|\eta_{t}*\nu(f)|\leq(\rho^{\kappa}+\tau){\mathcal{S}}_{\infty,1}(f).$
The argument relies on the quantitative decay of correlations for $X$. Consider the unitary representation of $G$ on $L^{2}(X)$ defined by the formula $g.f=f\circ g^{-1}$. From the combination of [35, Theorem 2.10] and [17, Equation 6.9], we know there exists $\delta_{0}=\delta_{0}(\Lambda)>0$ such that for any function $f\in B^{\infty}_{2,1}(X)$, any $g\in G$, we have
(36) $\lvert\langle g.f,f\rangle\rvert\ll\lVert g\rVert^{-\delta_{0}}{\mathcal{S}}_{2,1}(f)^{2}.$
From this we deduce a spectral gap for the family of Markov operators $P_{\eta_{t}}$. Recall that $P_{\eta_{t}}$ is the operator acting on non-negative measurable functions on $X$ by the formula
$P_{\eta_{t}}f(x)=\int_{G}f(gx)\,\mathrm{d}\eta_{t}(g).$
$P_{\eta_{t}}$ extends continuously into an operator on $L^{2}(X)$ of norm $1$.
Proposition 5.2 (Spectral gap for $P_{\eta_{t}}$).
There exists $c>0$ such that for any function $f\in B^{\infty}_{2,1}(X)$ with $m_{X}(f)=0$, we have
$\forall t>1,\quad\|P_{\eta_{t}}f\|_{L^{2}}\ll t^{-c}{\mathcal{S}}_{2,1}(f).$
Proof.
Using (36), we have
$\|P_{\eta_{t}}f\|^{2}_{L^{2}}$ $=\iint_{G^{2}}\langle g^{-1}.f,h^{-1}.f\rangle\,\mathrm{d}\eta_{t}(g)\,\mathrm{d}\eta_{t}(h)$
$=\iint_{G^{2}}\langle hg^{-1}.f,f\rangle\,\mathrm{d}\eta_{t}(g)\,\mathrm{d}\eta_{t}(h)$
$\ll{\mathcal{S}}_{2,1}(f)^{2}\iint_{G^{2}}\lVert hg^{-1}\rVert^{-\delta_{0}}\,\mathrm{d}\eta_{t}(g)\,\mathrm{d}\eta_{t}(h)$
Plugging in the definition of $\eta_{t}$
$hg^{-1}\,\mathrm{d}\eta_{t}(g)\,\mathrm{d}\eta_{t}(h)=u(t(s_{1}-s_{2}))\,\mathrm{d}\sigma(s_{1})\,\mathrm{d}\sigma(s_{2}),$
we get
$\lVert P_{\eta_{t}}f\rVert^{2}_{L^{2}}\ll\iint_{\mathbb{R}^{2}}\max\{1,t|s_{1}-s_{2}|\}^{-\delta_{0}}\,\mathrm{d}\sigma(s_{1})\,\mathrm{d}\sigma(s_{2}){\mathcal{S}}_{2,1}(f)^{2}.$
Finally, the Hölder regularity of $\sigma$ from Lemma 2.1(ii) implies that
$\sigma^{\otimes 2}\underbrace{\{(s_{1},s_{2}):t|s_{1}-s_{2}|\leq t^{1/2}\}}_{E}\ll t^{-c},$
for some constant $c=c(\sigma)>0$. Hence,
$\iint_{\mathbb{R}^{2}}\max\{1,t\lvert s_{1}-s_{2}\rvert\}^{-\delta_{0}}\,\mathrm{d}\sigma(s_{1})\,\mathrm{d}\sigma(s_{2})$
$=$ $\iint_{E}1\,\mathrm{d}\sigma\otimes\sigma+\iint_{\mathbb{R}^{2}\smallsetminus E}(t\lvert s_{1}-s_{2}\rvert)^{-\delta_{0}}\,\mathrm{d}\sigma(s_{1})\,\mathrm{d}\sigma(s_{2})$
$\ll$ $t^{-c}+t^{-\delta_{0}/2},$
concluding the proof. ∎
To prove Proposition 5.1 we mollify the measure $\nu$ at some scale $\rho>0$: let $\nu_{\rho}$ be the Borel measure on $X$ defined by
$\nu_{\rho}(f)=\frac{1}{m_{G}(B_{\rho})}\int_{X}\int_{B_{\rho}}f(gx)\,\mathrm{d}m_{G}(g)\,\mathrm{d}\nu(x),$
where $f$ denotes here any non-negative measurable function on $X$.
Note that if $\nu$ is supported on the compact set $\{\operatorname{inj}\geq\rho\}$, then by a change of variable $g\in B_{\rho}\mapsto gx\in B_{\rho}x$ and the Fubini-Lebesgue theorem, we have
$\nu_{\rho}(f)$ $=\frac{1}{m_{G}(B_{\rho})}\iint_{X\times X}\operatorname{{\bf 1}}_{y\in B_{\rho}x}f(y)\,\mathrm{d}m_{X}(y)\,\mathrm{d}\nu(x)$
$=\frac{1}{m_{G}(B_{\rho})}\int_{X}f(y)\int_{X}\operatorname{{\bf 1}}_{x\in B_{\rho}y}\,\mathrm{d}\nu(x)\,\mathrm{d}m_{X}(y).$
This implies that $\nu$ is absolutely continuous with respect to $m_{X}$ and its Radon-Nikodym derivative is
(37) $\frac{\mathrm{d}\nu_{\rho}}{\mathrm{d}m_{X}}(x)=\frac{\nu(B_{\rho}x)}{m_{G}(B_{\rho})}.$
In particular, if $\nu$ is $(1-\kappa,{\mathcal{B}}_{\rho},0)$-robust, then
$\Bigl{\lVert}\frac{\mathrm{d}\nu_{\rho}}{\mathrm{d}m_{X}}\Bigr{\rVert}_{L^{\infty}}\ll\rho^{-3\kappa}.$
Proof of Proposition 5.1.
We let $\kappa,r,\rho_{0}>0$ be parameters to specify below, $\rho,\tau,\nu$ as in the proposition, and consider a test function $f\in B^{\infty}_{\infty,1}(X)$ with zero average. Clearly we may assume $\tau=0$, i.e. $\nu$ is $(1-\kappa,{\mathcal{B}}_{\rho},0)$-robust.
We can write for any $n\geq 0$,
$|\eta_{t}*\nu(f)|=\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu\right\rvert\leq\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu_{\rho}\right\rvert+\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu_{\rho}-\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu\right\rvert.$
The first term is bounded by
$\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu_{\rho}\right\rvert$ $\leq\lVert P_{\eta_{t}}f\rVert_{L^{1}}\Bigl{\lVert}\frac{\mathrm{d}\nu_{\rho}}{\mathrm{d}m_{X}}\Bigr{\rVert}_{L^{\infty}}$
$\leq\lVert P_{\eta_{t}}f\rVert_{L^{2}}\Bigl{\lVert}\frac{\mathrm{d}\nu_{\rho}}{\mathrm{d}m_{X}}\Bigr{\rVert}_{L^{\infty}}$
$\ll t^{-c}{\mathcal{S}}_{2,1}(f)\rho^{-3\kappa},$
where the last inequality uses the spectral gap estimate from Proposition 5.2 on the one hand, and the assumption that $\nu$ is $(1-\kappa,{\mathcal{B}}_{\rho},0)$-robust on the other hand.
From the definition of $\nu_{\rho}$, the second term is bounded by
$\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu_{\rho}-\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu\right\rvert\leq\rho{\mathcal{S}}_{\infty,1}(P_{\eta_{t}}f)\ll\rho t{\mathcal{S}}_{\infty,1}(f).$
Put together, we have obtained
$|\eta_{t}*\nu(f)|\ll\bigl{(}t^{-c}\rho^{-3\kappa}+\rho t){\mathcal{S}}_{\infty,1}(f)\ll\rho^{\kappa}{\mathcal{S}}_{\infty,1}(f)$
where the last upper bound holds for $t\in[\rho^{-1/4},\rho^{-1/2}]$, up to choosing $\kappa=c/16$ and $\rho_{0}$ is small enough in terms of $\kappa$. ∎
We now address the
Proof of B’.
Note first that until now, we considered a measure $\lambda$ supported on $\operatorname{Aff}(\mathbb{R})^{+}$ while B’ allows for a measure $\lambda$ on $\operatorname{Aff}(\mathbb{R})$. We reduce easily to the $\operatorname{Aff}(\mathbb{R})^{+}$-case via the following lemma.
Lemma 5.3.
We may assume the measure $\lambda$ is supported on $\operatorname{Aff}(\mathbb{R})^{+}$.
Proof.
Set $\Omega=\operatorname{Aff}(\mathbb{R})^{\mathbb{N}}$, consider the stopping time $\tau_{+}:\Omega\rightarrow\mathbb{N}$ defined for $\underline{\phi}=(\phi_{i})_{i\geq 1}\in\Omega$ by
$\tau_{+}(\underline{\phi})=\inf\{n\geq 1\,:\,\mathtt{r}_{\phi_{1}\circ\dots\circ\phi_{n}}>0\}.$
Write $\lambda^{*\tau_{+}}=\int_{\Omega}\delta_{\phi_{1}\circ\dots\circ\phi_{\tau_{+}(\underline{\phi})}}\,\mathrm{d}\lambda^{\otimes\mathbb{N}}(\underline{\phi})$. Then by the strong Markov property, (see [3, Lemme A.2]), the measure $\sigma$ is $\lambda^{*\tau_{+}}$-stationary. Moreover, $\lambda^{*\tau_{+}}$ has finite exponential moment (because $\tau_{+}$ does). Its support $\operatorname{supp}\lambda^{*\tau_{+}}$ does not have common fixed point on $\mathbb{R}$, for otherwise the group generated by $\operatorname{supp}\mu$ would have an orbit of cardinality $2$ and hence fixes the barycenter of this orbit. ∎
Now that $\lambda$ is supported on $\operatorname{Aff}(\mathbb{R})^{+}$, we denote by $\mu$ the corresponding measure on $P$. We relate the $\eta_{t}$-process with the $\mu$-random walk thanks to the following lemma.
Lemma 5.4 ($\eta_{t}$-process vs $\mu$-walk).
Given $t>0$, $n\geq 0$, we have
$\eta_{t}=\int_{P}\eta_{t\mathtt{r}_{g}}*\delta_{g}\,\mathrm{d}\mu^{*n}(g).$
Proof.
Observe that for any $s\in\mathbb{R}$ and $g\in P$,
$a(t\mathtt{r}_{g})u(s)g=a(t\mathtt{r}_{g})u(s)a(\mathtt{r}_{g})^{-1}u(\mathtt{b}_{g})=a(t)u(\mathtt{r}_{g}s+\mathtt{b}_{g}).$
The claim then follows from the $\lambda^{*n}$-stationarity and the fact that $\mu^{*n}$ and $\lambda^{*n}$ are related by the anti-isomorphism between $P$ and $\operatorname{Aff}(\mathbb{R})^{+}$. ∎
We now discretize the set of values of $\mathtt{r}_{g}$ that appears in the part $\eta_{t\mathtt{r}_{g}}$. Given $r_{0},r_{1}>0$ observe that
$\eta_{tr_{0}}=\delta_{a(r_{0}r_{1}^{-1})}*\eta_{tr_{1}}$
Hence, for any finite Borel measure $\nu$ on $X$, we get
(38) $|\eta_{tr_{0}}*\nu(f)-\eta_{tr_{1}}*\nu(f)|\ll|\log(r_{0}r_{1}^{-1})|\,\nu(X){\mathcal{S}}_{\infty,1}(f).$
Let $\rho>0$, consider a parameter $\alpha\in(0,1)$ to be specified later depending on $\mu$, and set :=${\mathscr{R}}:=\{(1+\rho^{\alpha})^{k}\,:\,k\in\mathbb{Z}\}$. Combining (38) with Lemma 5.4, we get for any $x\in X$, $f\in B^{\infty}_{\infty,1}(X)$, that
(39) $\lvert\eta_{t}*\delta_{x}(f)\rvert\leq\sum_{r\in{\mathscr{R}}}\lvert\eta_{tr}*\mu^{n}_{r}*\delta_{x}(f)\rvert+O(\rho^{\alpha}{\mathcal{S}}_{\infty,1}(f))$
where $\mu^{n}_{r}$ denotes the restriction of $\mu^{*n}$ to the set $\{\,g\in P:\mathtt{r}_{g}\in[r,r(1+\rho^{\alpha})[\,\}$.
Let $\kappa=\kappa(\mu)>0$ as in Proposition 5.1. Assume $\operatorname{inj}(x)\geq\rho$. By Proposition 4.1, there are constants $C=C(\mu)>1$ and $\varepsilon_{1}=\varepsilon_{1}(\mu)>0$ such that, provided $\rho\lll 1$, the measure $\mu^{*n}*\delta_{x}$ on $X$ is $(1-\kappa,{\mathcal{B}}_{\rho},\rho^{\varepsilon_{1}})$-robust for any $n\geq C\lvert\log\rho\rvert$. For the rest of this proof, we specify $n,t$ in terms of $\rho$ as
(40) $n=\lceil C\lvert\log\rho\rvert\rceil,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,t=\rho^{-C\ell-3/8}.$
Consider
=${\mathscr{R}}^{\prime}=\{\,r\in{\mathscr{R}}:\rho^{-1/4}\leq tr\leq\rho^{-1/2}\,\}={\mathscr{R}}\cap[\rho^{-C\ell+1/8},\rho^{-C\ell-1/8}].$
On the one hand, for each $r\in{\mathscr{R}}^{\prime}$, note that $\mu^{n}_{r}*\delta_{x}\leq\mu^{*n}*\delta_{x}$ is still $(1-\kappa,{\mathcal{B}}_{\rho},\rho^{\varepsilon_{1}})$-robust. Therefore the choice of ${\mathscr{R}}^{\prime}$ allows us to use Proposition 5.1 to obtain
(41) $\lvert\eta_{tr}*\mu^{n}_{r}*\delta_{x}(f)\rvert\leq(\rho^{\kappa}+\rho^{\varepsilon_{1}}){\mathcal{S}}_{\infty,1}(f).$
On the other hand, by the large deviation estimates for sums of i.i.d real random variables, there is a constant $\varepsilon_{2}=\varepsilon_{2}(\mu,C)>0$ such that
$\mu^{*n}\{\,g:\lvert n\ell+\log\mathtt{r}_{g}\rvert>\lvert\log\rho\rvert/10\,\}<\rho^{\varepsilon_{2}}.$
whenever $\rho\lll_{C}1$. This implies an upper bound on the total mass $\sum_{r\in{\mathscr{R}}\smallsetminus{\mathscr{R}}^{\prime}}\mu^{n}_{r}(P)\leq\rho^{\varepsilon_{2}}$ and hence
(42) $\sum_{r\in{\mathscr{R}}\smallsetminus{\mathscr{R}}^{\prime}}\lvert\eta_{tr}*\mu^{n}_{r}*\delta_{x}(f)\rvert\leq\rho^{\varepsilon_{2}}{\mathcal{S}}_{\infty,1}(f).$
Putting (39), (41), (42) together, we have
$\lvert\eta_{t}*\delta_{x}(f)\rvert$ (#$\leq\bigl{(}\#{\mathscr{R}}^{\prime}(\rho^{\kappa}+\rho^{\varepsilon_{1}})+\rho^{\varepsilon_{2}}\bigr{)}{\mathcal{S}}_{\infty,1}(f)$
$\leq\bigl{(}\rho^{\kappa/2}+\rho^{\varepsilon_{1}/2}+\rho^{\varepsilon_{2}}\bigr{)}{\mathcal{S}}_{\infty,1}(f)$
where the second bound uses $\sharp{\mathscr{R}}^{\prime}\ll\frac{2C\ell|\log\rho|}{\log(1+\rho^{\alpha})}\sim|\log\rho|\rho^{-\alpha}$ and assumes $\alpha\leq\kappa\varepsilon_{1}/4$, $\rho\lll 1$.
Viewing $\rho$ as varying with $t$ according to (40), we can summarize the above as the following. For $t>1$ sufficiently large, for any $x\in X$ with $\operatorname{inj}(x)\geq t^{-(C\ell+3/8)^{-1}}$,
$\lvert\eta_{t}*\delta_{x}(f)\rvert\leq t^{-\varepsilon_{3}}{\mathcal{S}}_{\infty,1}(f)$
with $\varepsilon_{3}=\frac{\min\{\kappa,\varepsilon_{1},\varepsilon_{2}\}}{8C\ell+3}$. Finally, if $x$ is a point with $\operatorname{inj}(x)<t^{-(C\ell+3/8)^{-1}}$ then $\operatorname{inj}(x)^{-1}t^{-\varepsilon_{3}}\geq 1$. ∎
Effective equidistribution for the $\mu$-walk on $X$ can also be handled similarly (and more simply).
Proof of C’.
Proposition 5.2 and Proposition 5.1 are still valid with $(t,\eta_{t})$ replaced by $(e^{n},\mu^{n})$ where $n$ is an integer parameter (essentially same proof, using Lemma 2.3 instead of Lemma 2.1). Combining Proposition 5.1 with Proposition 4.1, we get the theorem. More precisely, given $\rho\lll 1$, Proposition 4.1 tells us that for $m\ggg|\log\rho|+|\log\operatorname{inj}(x)|$, the measure $\nu:=\mu^{*m}*\delta_{x}$ satisfies the conditions required to apply Proposition 5.1. Then choosing $n>m$ such that $n-m\in[\frac{1}{4}|\log\rho|,\frac{1}{2}|\log\rho|]$, we obtain that $\mu^{*n}*\delta_{x}$ is $\rho^{\varepsilon}$-equidistributed for some small constant $\varepsilon=\varepsilon(\mu)>0$. This finishes the proof. ∎
6. From dynamics to Diophantine approximation
In this final section, we justify why B’, case $\Lambda=\operatorname{SL}_{2}(\mathbb{Z})$, implies the Khintchine dichotomy in A’ (and in particular A). Throughout the section, we take $X=\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$.
Theorem 6.1.
The convergence part in A’ holds.
Proof.
Consequence of [19, Theorem 9.1] and B’, case $\Lambda=\operatorname{SL}_{2}(\mathbb{Z})$. ∎
In [19, Theorem 12.1], the divergence part is also obtained from effective equidistribution of expanding translates of $u(s)x_{0}\,\mathrm{d}\sigma(s)$ in $X$. However, [19] requires for this the extra assumption that $\sigma$ arises from a contractive IFS which satisfies the open set condition, the latter meaning the existence of a non-empty open set $O\subseteq\mathbb{R}$ whose images under the affine maps in the IFS are pairwise disjoint and included in $O$. We briefly explain how to adapt the proof of [19, Theorem 12.1] to our setting, in order to remove the assumptions of contractiveness and open set condition.
Theorem 6.2.
The divergence part in A’ holds.
As in [19], we first reduce to a slightly weaker estimate.
Proposition 6.3 (Divergence part bis).
Assume the approximation function $\psi$ satisfies $\sum_{q\in\mathbb{N}}\psi(q)=+\infty$. There exists $\delta_{\psi}>0$ depending only on $\psi$ such that
$\sigma(W(\psi))\geq\delta_{\psi}.$
To check that the above proposition does imply the divergence part in A’, we need Lemma 6.4 below which extends [19, Lemma 2.7] to the case of measures with potential overlaps.
Lemma 6.4.
Let $E\subseteq\mathbb{R}$ be a measurable set and let $\delta>0$. Assume that for every $n\in\mathbb{N}$, $\lambda^{*n}$-almost every $\phi\in\operatorname{Aff}(\mathbb{R})$, we have $\sigma(\phi^{-1}E)\geq\delta$. Then $\sigma(E)=1$.
Proof.
Set $\Omega=\operatorname{Aff}(\mathbb{R})^{\mathbb{N}}$ the space of instructions. By [2, Theorem 2.4] there exists a measurable map $\Omega\rightarrow\mathbb{R},\,{\underline{\phi}}\mapsto\mathtt{b}_{{\underline{\phi}}}$ such that for $\lambda^{\otimes\mathbb{N}}$-almost every ${\underline{\phi}}=(\phi_{i})_{i\geq 1}\in\Omega$, we have
$\mathtt{b}_{{\underline{\phi}}}=\lim_{k\to+\infty}\phi_{1}\circ\dots\circ\phi_{k}(0),$
and $\sigma$ coincides with the law of $\mathtt{b}_{{\underline{\phi}}}$ as ${\underline{\phi}}\overset{law}{\sim}\lambda^{\otimes\mathbb{N}}$.
Consider the function $f\colon\Omega\to\mathbb{R}$ defined by $f({\underline{\phi}})=\operatorname{{\bf 1}}_{E}(\mathtt{b}_{{\underline{\phi}}})$. Work on the probability space $(\Omega,\lambda^{\otimes\mathbb{N}})$ and let $({\mathcal{F}}_{n})_{n\in\mathbb{N}}$ denote the standard filtration of $\Omega=\operatorname{Aff}(\mathbb{R})^{\mathbb{N}}$, i.e. ${\mathcal{F}}_{n}$ is the sub $\sigma$-algebra of $\Omega$ spanned by the first $n$ coordinates. Using the equivariance relation
$\mathtt{b}_{\underline{\phi}}=\phi_{1}\circ\dots\circ\phi_{n}(\mathtt{b}_{(\phi_{n+1},\phi_{n+2},\dotsc)})$
we see that the assumption implies that the conditional expectation $\mathbb{E}[f|{\mathcal{F}}_{n}]$ satisfies
$\forall n\in\mathbb{N},\;\quad\mathbb{E}[f|{\mathcal{F}}_{n}]\geq\delta,\quad\lambda^{\otimes\mathbb{N}}-\text{almost everywhere}.$
By the martingale convergence theorem, $\mathbb{E}[f|{\mathcal{F}}_{n}]$ almost-surely converges to $f$. Thus
$f\geq\delta,\quad\lambda^{\otimes\mathbb{N}}-\text{almost everywhere}.$
But $f$ takes value in $\{0,1\}$, hence $f=1$ almost everywhere. This concludes the proof of $\sigma(E)=\lambda^{\otimes\mathbb{N}}(f)=1$. ∎
Lemma 6.5.
Proof.
For every $n\geq 0$ and $\lambda^{*n}$-almost every $\phi\in\operatorname{Aff}(\mathbb{R})$, the pushforward measure $\phi_{\star}\sigma$ is self-similar. If Proposition 6.3 holds, we obtain in particular $\sigma(\phi^{-1}W(\psi))\geq\delta_{\psi}$. Then $\sigma(W(\psi))=1$ thanks to Lemma 6.4. ∎
It remains to prove Proposition 6.3.
Proof of Proposition 6.3. Under the additional assumptions that the IFS $\Phi:=\operatorname{supp}\lambda$ is finite, contractive, and satisfies the open-set condition, the result is proven in [19, Theorem 12.3]. The only two steps in the proof of [19, Theorem 12.3] that rely on these extra assumptions are [19, Theorem 10.1] (long range decorrelation) and [19, Theorem 12.11]. We explain how to obtain those two results without extra assumption, i.e. in the setting of A’.
As for the proof of B’, we may assume without loss of generality that $\lambda$ is supported on $\operatorname{Aff}(\mathbb{R})^{+}$ (not only $\operatorname{Aff}(\mathbb{R})$), see Lemma 5.3. We may also argue under the conditions $1\geq 2^{n}\psi(2^{n})=o(1)$ and $\psi(q)\geq\frac{1}{q(\log q)^{1.1}}$ for all $n\geq 0$, $q\geq 3$, see [19, Lemmas 12.5, 12.6, 12.8].
Let us first adapt [19, Theorem 10.1] (long range decorrelation) to the overlapping case. Below, $x_{0}=\operatorname{SL}_{2}(\mathbb{Z})/\operatorname{SL}_{2}(\mathbb{Z})$ denotes the basepoint of $X$.
Proposition 6.6.
There exist constants $\delta>0,C_{*}\geq 1$ such that for every functions $f_{1},f_{2}\in B^{\infty}_{\infty,1}(X)$, for all $t\geq s>0$ satisfying $t\geq C_{*}s,$ we have
$\int f_{1}(a(e^{t})u(r)x_{0})f_{2}(a(e^{s})u(r)x_{0})\,\mathrm{d}\sigma(r)\\ =\int f_{1}\,\mathrm{d}m_{X}\int f_{2}(a(e^{s})u(r)x_{0})\,\mathrm{d}\sigma(r)+O({\mathcal{S}}_{\infty,1}(f_{1}){\mathcal{S}}_{\infty,1}(f_{2})e^{-\delta|t-s|}).$
Proof.
We set $\xi_{t}=a(e^{t})$ and abbreviate ${\mathcal{S}}_{\infty,1}$ by ${\mathcal{S}}$. We recall the notations $\mathtt{r}_{\phi}$, $\mathtt{b}_{\phi}$ have been defined in (4), and $\ell=-\int\log\mathtt{r}_{\phi}\,\mathrm{d}\lambda(\phi)>0$.
By Lemma 2.1, for any $R>1$, we have
(43) $\sigma\{r\,:\,|r|>R\}\ll R^{-\gamma}.$
By the large deviation principle for sums of i.i.d real random variables, there exists a constant $\varepsilon_{1}=\varepsilon_{1}(\lambda)>0$ such that
(44) $\lambda^{*n}\underbrace{\{\phi:|n\ell+\log\mathtt{r}_{\phi}|\leq\frac{n\ell}{10}\}}_{{\mathscr{C}}}\geq 1-e^{-n\varepsilon_{1}},$
provided $n\ggg 1$. In the following, $n\ggg 1$ is an integer to be specified later, and
${\mathscr{C}}:=\{\phi\in\operatorname{Aff}(\mathbb{R})^{+}:|n\ell+\log\mathtt{r}_{\phi}|\leq\frac{n\ell}{10}\}.$
We note that for any $\phi\in{\mathscr{C}}$ and $r,s\in\mathbb{R}$,
(45) $|f_{2}(\xi_{s}u(\phi(r))x_{0})-f_{2}(\xi_{s}u(\mathtt{b}_{\phi})x_{0})|\ll e^{s-9n\ell/10}|r|{\mathcal{S}}(f_{2}).$
Hence, writing =$\mathbb{R}=\{|r|>e^{n\ell/2}\}\cup\{|r|\leq e^{n\ell/2}\}$ and using (43),(45),
(46) $\int_{\mathbb{R}}|f_{2}(\xi_{s}u(\phi(r))x_{0})-f_{2}(\xi_{s}u(\mathtt{b}_{\phi})x_{0})|\,\mathrm{d}\sigma(r)\ll(e^{-n\ell\gamma/2}+e^{s-2n\ell/5}){\mathcal{S}}(f_{2}).$
On the other hand, direct computation shows that for any $t>0$, $r\in\mathbb{R}$,
(47) $\xi_{t}u(\phi(r))=\xi_{t+\log\mathtt{r}_{\phi}}u(r)h_{\phi},$
where $h_{\phi}=\xi_{-\log\mathtt{r}_{\phi}}u(\mathtt{b}_{\phi})$. We deduce from (47) and B’ that for any $\phi\in{\mathscr{C}}$, $t>0$,
$\int_{\mathbb{R}}f_{1}(\xi_{t}u(r)x_{0})\,\mathrm{d}({\phi}_{\star}\sigma)(r)$ $=\int_{\mathbb{R}}f_{1}(\xi_{t+\log\mathtt{r}_{\phi}}u(r)h_{\phi}x_{0})\,\mathrm{d}\sigma(r)$
$=\int_{X}f_{1}\,\mathrm{d}m_{X}+O(\operatorname{inj}(h_{\phi}x_{0})^{-1}{\mathcal{S}}(f_{1})e^{-c(t+\log\mathtt{r}_{\phi})})$
(48) $=\int_{X}f_{1}\,\mathrm{d}m_{X}+O({\mathcal{S}}(f_{1})e^{-ct+c_{1}n\ell})$
where $c=c(\lambda)>0$ comes from B, $c_{1}=\frac{11}{10}(1+c)$, and the last equality follows from observing that the orbit $u(\mathbb{R})x_{0}$ is bounded whence $\operatorname{inj}(h_{\phi}x_{0})^{-1}\ll r_{\phi}^{-1}\leq e^{\frac{11}{10}n\ell}$.
We then have the following estimates:
×$\,\,\,\,\,\int_{\mathbb{R}\times{\mathscr{C}}}f_{1}(\xi_{t}u(r)x_{0})f_{2}(\xi_{s}u(r)x_{0})\,\mathrm{d}\phi_{\star}\sigma(r)\,\mathrm{d}\lambda^{*n}(\phi)$
=(46)×${=}\int_{\mathbb{R}\times{\mathscr{C}}}f_{1}(\xi_{t}u(\phi(r))x_{0})f_{2}(\xi_{s}u(\mathtt{b}_{\phi})x_{0})\,\mathrm{d}\sigma(r)\,\mathrm{d}\lambda^{*n}(\phi)+O((e^{-n\ell\gamma/2}+e^{s-2n\ell/5}){\mathcal{S}}(f_{1}){\mathcal{S}}(f_{2}))$ )
[…]
(49) $\,\,\,\,\,\,\,\,\,\,\,+O\left((e^{-n\ell\gamma/2}+e^{s-2n\ell/5}+e^{-ct+c_{1}n\ell}){\mathcal{S}}(f_{1}){\mathcal{S}}(f_{2})\right).$
In view of (44), this estimate still holds with the domain of integration ${\mathscr{C}}$ replaced by $\operatorname{Aff}(\mathbb{R})$, up to adding $O(e^{-n\varepsilon_{1}}{\mathcal{S}}(f_{1}){\mathcal{S}}(f_{2}))$ to the error term. Invoking on top of that the $\lambda^{*n}$-stationarity of $\sigma$, we conclude that
(50) $\int_{\mathbb{R}}f_{1}(\xi_{t}u(r)x_{0})f_{2}(\xi_{s}u(r)x_{0})\,\mathrm{d}\sigma(r)=\int_{X}f_{1}\,\mathrm{d}m_{X}\int_{\mathbb{R}}f_{2}(\xi_{s}u(r)x_{0})\,\mathrm{d}\sigma(r)\\ +O\left(e^{-n\varepsilon_{1}}+e^{-n\ell\gamma/2}+e^{s-2n\ell/5}+e^{-ct+c_{1}n\ell}\right){\mathcal{S}}(f_{1}){\mathcal{S}}(f_{2}).$
To finish the proof, we need to choose parameters $C_{*},\delta>0$ depending on $\lambda$, and an integer parameter $n=n(\lambda,s,t)\ggg 1$ so that the above error term is $O(e^{-\delta t})$ provided $t\geq C_{*}s$ and $s\ggg 1$. To do this, we take $n=\lceil s+\alpha t\rceil$ where $\alpha=\alpha(\mu)>0$ is small enough to satisfy $c-c_{1}\alpha\geq c/2$. Then taking $C_{*}$ large enough in terms of $\mu$ (and $\alpha$), the claim becomes transparent for each of the exponentials in the error term. ∎
We now establish the analogue of [19, Theorem 12.11] in our context. Denote by ${\mathscr{P}}(\mathbb{Z}^{2})$ the collection of primitive vectors in $\mathbb{Z}^{2}$ and consider
$A^{*}_{n}:=\{r\in\mathbb{R}\,:\,\exists(p,q)\in{\mathscr{P}}(\mathbb{Z}^{2})\text{ such that }\,2^{n-1}\leq q<2^{n}\text{ and }|qr-p|<\psi(2^{n})\}.$
We need to show
Proposition 6.7.
There exist constants $C,\gamma,\varepsilon>0$ depending on $\lambda$ such that for all large enough integers $m\leq n\leq(1+\varepsilon)m$, we have
$\sigma(A^{*}_{m}\cap A^{*}_{n})\leq C\sigma(A^{*}_{m})\max\left\{\sigma(A^{*}_{n})^{\gamma},2^{-\gamma(n-m)}\right\}.$
Remark.
[19, Proposition 12.11] only assumes $m\leq n$, while here we look at short range correlations i.e. $m\leq n\leq(1+\varepsilon)m$. Ultimately, the proof of [19, Theorem 12.3] in Khalil-Luethi only uses [19, Proposition 12.11] in the short correlation case. Indeed, [19, Proposition 12.11] is used to verify condition (3) of [19, Proposition 11.1], while condition (3) is used to handle short range independence of $A^{*}_{m}$ and $A^{*}_{n}$ when $n\in[m,m+K\lceil-a\log\sigma(A^{*}_{m})\rceil]$, where $K,a>0$ are constants allowed to depend on $\psi$ and $\sigma$ (see Step 3 in the proof of [19, Proposition 11.1]). Using the reduction $\psi(q)\geq\frac{1}{q(\log q)^{1.1}}$ for all $q\geq 3$ and [19, Equations 12.16, 12.17], we have
$\lim_{m\to\infty}\frac{K\lceil-a\log\sigma(A^{*}_{m})\rceil]}{m}=0.$
Therefore, the above statement is sufficient to apply their argument.
To prove Proposition 6.7, we introduce another family of sets $B^{*}_{n}$, which slightly enlarges the sets $A^{*}_{n}$,
$B^{*}_{n}:=\{r\in\mathbb{R}\,:\,\exists(p,q)\in{\mathscr{P}}(\mathbb{Z}^{2})\text{ such that }\,2^{n-2}\leq q<2^{n+1}\text{ and }\lvert qr-p\rvert<2\psi(2^{n})\}.$
We will also need a dynamical characterization of the sets $A^{*}_{n}$, $B^{*}_{n}$. Consider the function $(1,+\infty)\rightarrow\mathbb{R},t\mapsto r(t)$ from [19, Lemma 8.1]. Then, as in [19], given sufficiently large $n\geq 1$, we let $t_{n}>1$ be the unique solution of
(51) $e^{t_{n}-r(t_{n})}=2^{n}.$
Then, by the construction of [19, Lemma 8.1], we have
(52) $\psi(2^{n})=e^{-t_{n}-r(t_{n})}.$
We also use the shorthand $g_{t}=a(e^{2t})$. We introduce the following domains in $\mathbb{R}^{2}$,
${\mathcal{S}}:=(-1,1)^{2},$
${\mathcal{A}}:=\left\{z\in\mathbb{R}^{2}\,:\,|z_{2}|<\frac{1}{2}\right\},\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\mathcal{B}}:=\left\{z\in\mathbb{R}^{2}\,:\,|z_{2}|<\frac{1}{4}\right\}.$
${\mathcal{S}}$ is the standard open square of size $2$, and ${\mathcal{A}}$ are ${\mathcal{B}}$ horizontal bands around the absciss axis of respective width $1$ and $1/2$.
We then define $U^{*}_{n}\subseteq X$ as the subset of points representing lattices that meet the renormalized square $e^{-r(t_{n})}{\mathcal{S}}$ but not its intersection with the band $e^{-r(t_{n})}{\mathcal{A}}$. Similarly, we define $V^{*}_{n}\subseteq X$, asking for lattices meeting $e^{-r(t_{n})}2{\mathcal{S}}$ but not its intersection with the band $e^{-r(t_{n})}{\mathcal{B}}$. Here we make requirements for lattices, but we could just as well consider the systole of such lattice (meaning its vector of smallest non-zero norm), as this would not affect definitions provided $n\ggg_{\psi}1$.
Using (51) and (52), it is straightforward to check from definitions that $A^{*}_{n}$ (resp. $B^{*}_{n}$) coincides with the set of $s\in\mathbb{R}$ such that $g_{t_{n}}u(s)x_{0}\in U^{*}_{n}$ (resp. $V^{*}_{n}$). Moreover, we note that for some small absolute constant $\eta>0$, the $\eta$-neighborhood of $U^{*}_{n}$ is included in $V^{*}_{n}$ (provided $n\ggg_{\psi}1$).
Proof of Proposition 6.7.
For $n\ggg_{\psi}1$, we write $\varphi^{+}_{n}:X\rightarrow[0,1]$ the $\eta/2$-mollification of ${\mathbf{1}}_{\{\operatorname{dist}(U^{*}_{n},.)<\eta/2\}}$ the characteristic function of the $\eta/2$-neighborhood of $U^{*}_{n}$. In particular, one has
${\mathbf{1}}_{U^{*}_{n}}\leq\varphi^{+}_{n}\leq{\mathbf{1}}_{V^{*}_{n}}\,\,\,\,\text{ and }\,\,\,\,{\mathcal{S}}_{\infty,1}(\varphi^{+}_{n})\ll 1.$
The bound ${\mathbf{1}}_{U^{*}_{n}}\leq\varphi^{+}_{n}$ yields for any $m,n\ggg_{\psi}1$
$\sigma(A^{*}_{m}\cap A^{*}_{n})$ $\leq\int_{\mathbb{R}}\varphi^{+}_{m}(g_{t_{m}}u(s)x_{0})\varphi^{+}_{n}(g_{t_{n}}u(s)x_{0})\,\mathrm{d}\sigma(s)$
$=\int_{\mathbb{R}}F(g_{t_{m}}u(s)x_{0})\,\mathrm{d}\sigma(s)$
where $F:X\rightarrow\mathbb{R}$ is defined by the formula
$F(y):=\varphi^{+}_{m}(y)\varphi^{+}_{n}(g_{t_{n}-t_{m}}y).$
We now estimate the above integral. We let $\varepsilon>0$ be a parameter to specify below, and assume $m\leq n\leq(1+\varepsilon)m$. Noting that ${\mathcal{S}}_{\infty,1}(F)\ll e^{l(t_{n}-t_{m})}$ where $l>0$ is an absolute constant, we get by effective equidistribution (B’) that for some $c=c(\lambda)>0$,
$\int_{\mathbb{R}}F(g_{t_{m}}u(s)x_{0})\,\mathrm{d}\sigma(s)$ $=\int_{X}F\,\mathrm{d}m_{X}+O({\mathcal{S}}_{\infty,1}(F)e^{-ct_{m}})$
(53) $=\int_{X}F\,\mathrm{d}m_{X}+O(e^{l|t_{n}-t_{m}|-ct_{m}})$
However, by [19, Equation 12.32], one has
(54) $t_{n}-t_{m}=\frac{\log 2}{2}(n-m)+\frac{1}{2}\log\frac{\psi(2^{m})}{\psi(2^{n})}.$
Recalling the conditions added on $\psi$, and assuming $m\ggg_{\varepsilon}1$, one gets
(55) $0\leq t_{n}-t_{m}\ll(n-m)+\log m\ll\varepsilon m\ll\varepsilon t_{m},$
where the last inequality relies on the lower bound $t_{m}\geq\frac{1}{8}m$, which follows from Equation 54. Combining (6), (55) and assuming $\varepsilon\lll c/l$, we get
(56) $\int_{\mathbb{R}}F(g_{t_{m}}u(s)x_{0})\,\mathrm{d}\sigma(s)$ $=\int_{X}F\,\mathrm{d}m_{X}+O(e^{-\frac{c}{16}m}).$
In particular,
(57) $\sigma(A^{*}_{m}\cap A^{*}_{n})$ $\leq\int_{X}F\,\mathrm{d}m_{X}+O(e^{-\frac{c}{16}m}).$
Note that $\sigma$ is absent from the integral term in the upper bound.
We now work to reverse the sense of the inequality, i.e. get an upper bound for $\int_{X}F\,\mathrm{d}m_{X}$ up to enlarging the set $A^{*}_{n}$. Recall the sets $(B^{*}_{n},V^{*}_{n})$ defined earlier and the inequality ${\mathbf{1}}_{V^{*}_{n}}\geq\varphi^{+}_{n}$. We have that
$\sigma(B^{*}_{m}\cap B^{*}_{n})$ $\geq\int_{\mathbb{R}}\varphi^{+}_{m}(g_{t_{m}}u(r)x_{0})\varphi^{+}_{n}(g_{t_{n}}u(r)x_{0})\,\mathrm{d}\sigma(r)$
(58) $=\int_{X}F\,\mathrm{d}m_{X}+O(e^{-\frac{c}{16}m})$
where the second line is given by (56). Note that we could have performed the proof of (6) for any other self-similar measure $\sigma_{0}$ in the place of $\sigma$, say $\sigma_{0}:=\text{leb}_{[0,1]}$. Combining this remark with (57) and (6), we get
$\sigma(A^{*}_{m}\cap A^{*}_{n})$ $\leq\sigma_{0}(B^{*}_{m}\cap B^{*}_{n})+O(e^{-\frac{c}{16}m}).$
up to adjusting the value of $c$ from the start (to account for $\sigma_{0}$).
We may now conclude the proof as follows. Observe that $A^{*}_{m}\cap A^{*}_{n}\subseteq B^{*}_{m}\cap B^{*}_{n}$ and each $B^{*}_{i}$ is included in the union of the three sets corresponding to $A^{*}_{i-1},A^{*}_{i},A^{*}_{i+1}$ but for approximation functions $2\psi(2q),2\psi(q),2\psi(q/2)$ (whose respective sequences $r(t_{n})$ have bounded mutual differences). This allows to apply [19, Proposition 12.11] to $\sigma_{0}$ to bound
$\sigma_{0}(B^{*}_{m}\cap B^{*}_{n})\ll\sigma_{0}(B^{*}_{m})\max\left\{\sigma_{0}(B^{*}_{n})^{\gamma},2^{-\gamma(n-m)}\right\},$
where $\gamma>0$ is an absolute constant. Then applying [19, Equation 12.16], we see that the ratio $\sigma_{0}(B^{*}_{i})/\sigma(A^{*}_{i})$ is bounded below and above by two absolute positive constants (provided $i\ggg 1$ depending on $\sigma$, $\psi$), hence we may replace each $\sigma_{0}(B^{*}_{i})$ by $\sigma(A^{*}_{i})$ in the right-hand side. Finally observe that the additional error term $O(e^{-\frac{c}{16}m})$ is $O(\sigma(A^{*}_{m})\sigma(A^{*}_{n}))$. This last claim follows from the fact that $\sigma(A^{*}_{m})$ decays at most polynomially in $m$ (by [19, Equations 12.16, 12.17] again and the conditions $\psi(q)\geq q^{-1}(\log g)^{-1.1}$) and the requirement $n\in[m,(1+\varepsilon)m]$. ∎