### v3

# Khintchine dichotomy for self-similar measures

Abstract.
We establish the analogue of Khintchine’s theorem for all self-similar probability measures on the real line. When specified to the case of the Hausdorff measure on the middle-thirds Cantor set, the result is already new and provides an answer to an old question of Mahler. The proof consists in showing effective equidistribution in law of expanding upper-triangular random walks on $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$, a result of independent interest.
2010 Mathematics Subject Classification:
Primary 37A99, 11J83; Secondary 22F30.
W.H. is supported by the National Key R&D Program of China (No. 2022YFA1007500) and the National Natural Science Foundation of China (No. 12288201).
H.Z. is supported by the startup grant of Soochow University.
1. Introduction
A Borel probability measure $\sigma$ on the real line $\mathbb{R}$ is called self-similar if it satisfies
(1.1) $\sigma=\sum_{i=1}^{\mathtt{m}}\lambda_{i}\,\phi_{i\star}\sigma$
for some integer $m\geq 1$, some probability vector $(\lambda_{1},\cdots,\lambda_{\mathtt{m}})\in\mathbb{R}_{>0}^{\mathtt{m}}$, and some invertible affine maps $\phi_{1},\dotsc,\phi_{\mathtt{m}}:\mathbb{R}\to\mathbb{R}$ without common fixed point. This includes Hausdorff measures on missing digit Cantor sets. For example, the one on the middle-thirds Cantor set satisfies (1.1) with $\lambda_{1}=\lambda_{2}=1/2$ and $\phi_{1}:t\mapsto t/3$ and $\phi_{2}:t\mapsto t/3+2/3$. Another standard definition of self-similar measures requires that all the maps $\phi_{i}$ are contracting. We do not impose such a condition, see §2.1 for further discussion.
It is particularly intriguing to explore the Diophantine properties of points within the support of a self-similar measure. This research topic was proposed by Mahler in [38, Section 2], asking how close irrational numbers in the middle-thirds Cantor set can be approximated by rational numbers. One approach to framing Mahler’s question is by investigating whether Khintchine’s theorem extends to the middle-thirds Cantor measure (as asked by Kleinbock-Lindenstrauss-Weiss in [29, Section 10.1]).
Let us recall the classical Khintchine theorem. Here and hereafter, $\psi:\mathbb{N}\to\mathbb{R}_{>0}$ is a function that will be referred to as an approximation function. A point $s\in\mathbb{R}$ is called $\psi$-approximable if there exist infinitely many $(p,q)\in\mathbb{Z}\times\mathbb{N}$ such that
(1.2) $|qs-p|<\psi(q).$
Denote by $W(\psi)$ the set of $\psi$-approximable points in $\mathbb{R}$. The classical Khintchine theorem for the Lebesgue measure [26, 27] states that given a non-increasing approximation function $\psi$, the set $W(\psi)$ has null Lebesgue measure if the series $\sum_{q\in\mathbb{N}}\psi(q)$ is convergent, and full Lebesgue measure otherwise.
In this paper, we establish the analogue of Khintchine’s theorem for self-similar measures on $\mathbb{R}$.
Theorem A (Khintchine’s theorem for self-similar measures).
Let $\sigma$ be a self-similar probability measure on $\mathbb{R}$, let $\psi:\mathbb{N}\to\mathbb{R}_{>0}$ be a non-increasing function. Then
(1.3) $\sigma(W(\psi))=\left\{\begin{array}[]{ll}0&\text{if }\sum_{q\in\mathbb{N}}\psi(q)<\infty,\\ &\\ 1&\text{if }\sum_{q\in\mathbb{N}}\psi(q)=\infty.\end{array}\right.$
In the divergent case, given a $\sigma$-typical $s\in\mathbb{R}$, we also obtain estimates on the number of solutions $(p,q)$ of the inequality (1.2) with bounded $q$, see (7.4) and (7.5).
Let us briefly present the state of the art surrounding Khintchine’s theorem on fractals.
For the convergence part, the case $\psi(q)=1/q^{1+\varepsilon}$ was treated by Weiss [52] for measures satisfying certain decay conditions, comprising the case of the middle-thirds Cantor measure. Weiss’ result was later generalized to friendly measures on $\mathbb{R}^{d}$ for arbitrary positive integer $d$ by Kleinbock-Lindenstrauss-Weiss [29]. See also the related work of Pollington-Velani [40] on absolutely friendly measures, and that of Das-Fishman-Simmons-Urbański on quasi-decaying measures [16, 17].
For the divergence part, the case $\psi(q)=\varepsilon/q$ was treated by Einsiedler-Fishman-Shapira [20] for missing digit Cantor measures. Simmons-Weiss [46] then significantly generalized their result, promoting it to arbitrary self-similar measures on $\mathbb{R}^{d}$ (along with several refinements).
All the above works focus on specific approximation functions $\psi$. Under the sole condition that $\psi$ is non-increasing, Khalil and Luethi [24] were able to extend Khintchine’s theorem to self-similar measures $\sigma$ on $\mathbb{R}^{d}$, provided $\sigma$ has large dimension and the underlying IFS $(\phi_{i})_{1\leq i\leq m}$ is contractive, rational, and satisfies the open set condition. In particular, they derived Khintchine’s theorem for one-missing digit Cantor sets in base $5$. With a different approach, based on Fourier analysis, Yu [54] also achieved the convergence part of Khintchine’s theorem for general approximation functions, provided $\sigma$ is a measure with sufficiently fast average Fourier decay. The divergence part was very recently settled by Datta-Jana [18] under similar restrictions, reaching some cases that are not covered by [24] such as a $3$-missing digit Cantor set in base $450$.
All the aforementioned works impose various constraints for the Khintchine dichotomy (1.3) to hold for a fractal measure. Specifically, none of them establishes (1.3) in the case of the middle-thirds Cantor measure advertised by Mahler. A is already innovative in this regard, and goes way beyond this example.
Other related research topics. As Mahler pointed out in [38], it is also interesting to investigate intrinsic Diophantine approximation on a Cantor set. This means asking how well points on a fractal set can be approximated by rational points sitting inside the fractal set itself. We refer to recent works [50, 13] and references therein for related research in this direction.
In addition to fractals, Khintchine’s theorem has been extensively studied on submanifolds of $\mathbb{R}^{d}$. Major works in this area include [33, 51, 9, 10].
A is derived from an effective equidistribution result in homogeneous dynamics which we now present. Consider the real algebraic group $G=\operatorname{SL}_{2}(\mathbb{R})$, a lattice $\Lambda\subseteq G$, and the quotient space $X=G/\Lambda$, endowed with the standard hyperbolic metric (§2.1) and the Haar probability measure $m_{X}$. Write $\operatorname{inj}(x)$ the injectivity radius at $x\in X$2.1). Denote by $B^{\infty}_{\infty,1}(X)$ the set of smooth functions on $X$ which are bounded and have bounded order-$1$ derivatives, write ${\mathcal{S}}_{\infty,1}(\cdot)$ the corresponding Sobolev norm (§2.1). For $t>0$, $s\in\mathbb{R}$, write $a(t),u(s)\in G$ the elements given by
$a(t)=\begin{pmatrix}t^{1/2}&0\\ 0&t^{-1/2}\end{pmatrix}\qquad u(s)=\begin{pmatrix}1&s\\ 0&1\end{pmatrix}.$
Theorem B (Effective equidistribution of expanding fractals).
Let $\sigma$ be a self-similar probability measure on $\mathbb{R}$. There exists a constant $c=c(\Lambda,\sigma)>0$ such that for all $t>1$, $x\in X$, $f\in B^{\infty}_{\infty,1}(X)$, we have
(1.4) $\int_{\mathbb{R}}f(a(t)u(s)x)\,\mathrm{d}\sigma(s)=\int_{X}f\,\mathrm{d}m_{X}\,+\,O\bigl{(}\operatorname{inj}(x)^{-1}{\mathcal{S}}_{\infty,1}(f)t^{-c}\bigr{)}$
where the implicit constant in $O(\cdot)$ only depends on $\Lambda$ and $\sigma$.
B states the exponential equidistribution of the measure $\sigma$ seen on a piece of horocycle based at $x$ and expanded by the action of the geodesic flow. The exponent $c$ in the rate of equidistribution is uniform in $x$, however, equidistribution may take more time to start when $x$ is high in the cusp. This is reflected by the term $\operatorname{inj}(x)^{-1}$ in the rate. A refinement of B tackling double equidistribution will also be established, see Equation 6.3.
The link between homogeneous dynamics and Diophantine approximation is known as Dani’s correspondence [15]. In [34], Kleinbock-Margulis explicitely demonstrated how to use dynamics to obtain a new proof of the classical Khintchine theorem for the Lebesgue measure, see also the variant [49] by Sullivan and the seminal work of Patterson [39]. This dynamical perspective laid the foundation for many subsequent works generalizing Khintchine’s theorem in various aspects, see e.g. [31, 14, 24]. In particular, the implication from (1.4) to the convergent case of A is given in the work of Khalil-Luethi [24, Theorem 9.1]. Under the extra assumptions that $\sigma$ arises from a contractive IFS satisfying the open set condition, they also show that (1.4) is sufficient to establish the divergent case of A, see [24, Theorem 12.1]. Their proof relies on a subtle inverse Borel-Cantelli Lemma. Here, we adopt an approach that is closer to Schmidt’s original proof of the quantitative Khintchine theorem [43]. Taking advantage of B, this enables us to get rid of extra assumptions and has the double advantage of being shorter and quantitative, see Section 7.
Besides its applications to Diophantine approximation, B is interesting in its own right. It can be seen as a fractal and effective version of Ratner’s equidistribution theorem for unipotent flows on $X$. Recall that Ratner’s theorem states that any unipotent orbit on a finite-volume homogeneous space equidistributes within the smallest finite-volume homogeneous subspace that contains it. Unfortunately, the proof gives no information on the rate of equidistribution. Over the past years, substantial efforts were made to obtain an effective version of Ratner’s theorem, in other terms quantify the equidistribution of large but bounded pieces of unipotent orbits. In the case where the unipotent orbit arises from the action of a horospherical subgroup, Kleinbock and Margulis established in [32, 35] the effective equidistribution of expanding translates under the corresponding diagonal flow. More recently, significant progress on effective Ratner was made by Einsiedler-Margulis-Venkatesh [21], Strömbergsson [48], Kim [28], Lindenstrauss-Mohammadi [36], Lindenstrauss-Mohammadi-Wang [37], and Yang [53]. We note that these works focus on the expanding translates of the Haar measure on a piece of unipotent orbit. In [24], Khalil and Luethi obtain the first effective equidistribution of expanding fractal measures on a unipotent orbit in $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$. They argue under the assumption that the underlying IFS is contractive, rational, satisfies the open set condition, and the measure $\sigma$ is thick enough. They also require that the starting point $x$ belongs to a specific countable set related to the IFS. In Datta-Jana [18], effective equidistribution for expanding measures are also obtained in $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$ assuming sufficiently fast average Fourier decay and restrictions on the starting point $x$. B generalizes Khalil-Luethi’s and Datta-Jana’s equidistribution results in $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$ in so far as it allows for an arbitrary lattice $\Lambda$, any starting point $x$, and most importantly any self-similar measure $\sigma$. The dependence of our error term on the starting point is also more precise.
Remark. The weak-* convergence $\lim_{t\to+\infty}a(t)u(s)x\,\mathrm{d}\sigma(s)=m_{X}$ resulting from B is also new. Convergence without rate is also addressed in the independent concurrent work of Khalil-Luethi-Weiss [25] for rational carpet IFS’s in all dimensions. Note however that effectivity, and more precisely an exponential convergence rate, is crucial to derive the Khintchine dichotomy (1.3) through Dani’s correspondence.
We prove B from the point of view of random walks. The connection between the asymptotic behaviour of an expanding fractal and that of a random walk is rooted in the work of Simmons-Weiss [46] and then further exploited in [41, 42, 24, 19, 1]. In our paper, this connection takes the form of Lemma 5.4.
We establish the following effective equidistribution in law for random walks driven by expanding upper triangular matrices on $X$. In the statement below, $\mathbb{R}^{2}$ is endowed with the usual Euclidean structure and we write $e_{1}:=(1,0)\in\mathbb{R}^{2}$.
Theorem C (Effective equidistribution for random walks).
Let $\mu$ be a finitely supported probability measure on the group
$\{\,a(t)u(s):t>0,\,s\in\mathbb{R}\,\}\subseteq G.$
Assume that the support of $\mu$ is not simultaneously diagonalisable, and $\mu$ satisfies $\int_{G}\log\|ge_{1}\|\,\mathrm{d}\mu(g)>0$. Then there exists a constant $c=c(\Lambda,\mu)>0$ such that for all $x\in X$, $n\geq 1$ and $f\in B^{\infty}_{\infty,1}(X)$, we have
$\mu^{*n}*\delta_{x}(f)=m_{X}(f)+O\bigl{(}\operatorname{inj}(x)^{-1}{\mathcal{S}}_{\infty,1}(f)e^{-cn}\bigr{)}$
where the implicit constant in $O(\cdot)$ only depends on $\Lambda$ and $\mu$.
The proof of C is inspired by [5], where the first-named and second-named authors establish effective equidistribution for random walks on $X$ which are driven by a Zariski-dense probability measure on $G$. In our context however, the acting group is solvable. The proof consists of three phases. Each step concerns the dimension of the distribution of the random walk at some scale.
First, we show that the random walks gains some initial positive dimension: there exist constants $\kappa>0$, $A>0$ determined by $\mu$ such that for every small $\rho>0$, $x,y\in X$, every $n\geq|\log\rho|+A|\log\operatorname{inj}(x)|$,
${\mu^{*n}}*\delta_{x}(B_{\rho}y)\leq\rho^{\kappa}$
where the notation $B_{\rho}y$ refers to the open ball of radius $\rho$ centered at $y$ in $X$.
Second, we bootstrap the value of the exponent $\kappa$ arbitrarily close to $3$, say up to $3-\varepsilon$ provided $\rho\leq\rho_{0}(\varepsilon,\mu)$ and $n\geq C_{0}(\varepsilon,\mu)|\log\rho|+A|\log\operatorname{inj}(x)|$. The method is based on the multislicing argument from [5], which, in turn, relies on discretized projection theorems à la Bourgain.
Finally, once the dimension is close to full, we conclude using the spectral gap of the convolution operator $f\mapsto\mu*f$ acting on $L^{2}(X)$.
B follows from C, using Lemma 5.4 and a probabilistic argument. The convergence part of A is then a direct consequence of B, case $\Lambda=\operatorname{SL}_{2}(\mathbb{Z})$, and [24, Theorem 9.1]. The divergence part is obtained from a refinement of B about double equidistribution, inspired by [30], and builds upon Schmidt’s original proof of the quantitative classical Khintchine theorem [43].
Allowing $\lambda$ to have infinite support. Our method allows for slightly more general statements, extending the aforementioned Khintchine dichotomy and equidistribution results to measures arising from randomized IFS with potentially infinite support, provided a finite exponential moment.
We let $\operatorname{Aff}(\mathbb{R})$ denote the affine group of $\mathbb{R}$. For any $\phi\in\operatorname{Aff}({\mathbb{R}})$, we let $\mathtt{r}_{\phi}\in\mathbb{R}^{*}$, $\mathtt{b}_{\phi}\in\mathbb{R}$ denote the unique numbers such that
(1.5) $\phi(t)=\mathtt{r}_{\phi}t+\mathtt{b}_{\phi},\quad\forall t\in\mathbb{R}.$
We say a probability measure $\lambda$ on $\operatorname{Aff}(\mathbb{R})$ has a finite exponential moment if there exists $\varepsilon>0$ such that
(1.6) $\int_{\operatorname{Aff}(\mathbb{R})}|\mathtt{r}_{\phi}|^{\varepsilon}+|\mathtt{r}_{\phi}^{-1}|^{\varepsilon}+|\mathtt{b}_{\phi}|^{\varepsilon}\,\mathrm{d}\lambda(\phi)<\infty.$
Theorem A’.
Let $\lambda$ be a probability measure on $\operatorname{Aff}(\mathbb{R})$ with a finite exponential moment and such that $\operatorname{supp}\lambda$ does not have a global fixed point. Let $\sigma$ be a probability measure on $\mathbb{R}$ satisfying $\lambda*\sigma=\sigma$. Then $\sigma$ satisfies the Khintchine dichotomy (1.3).
Theorem B’.
Under the same assumptions, $\sigma$ satisfies the effective equidistribution for expanding translates from Equation 1.4.
Recall that a probability measure $\mu$ on $G$ has a finite exponential moment if for some $\varepsilon>0$, we have
(1.7) $\int_{G}\|g\|^{\varepsilon}\,\mathrm{d}\mu(g)<\infty.$
Theorem C’.
C is valid when the finite support assumption on $\mu$ is relaxed into a finite exponential moment condition.
Structure of the paper. In Section 2, we fix notations for the rest of the paper, we highlight basic facts concerning self-similar measures, and we recall some recurrence properties of the $\mu$-walk on $X$. In Section 3, we deduce positive dimension of ${\mu^{*n}}*\delta_{x}$ at exponentially small scales. In Section 4, we bootstrap the dimension until it reaches a number arbitrarily close to $3=\dim X$. In Section 5, we deduce the equidistribution statements, namely B’ and C’. In Section 6, we upgrade B’ into a double equidistribution statement. In Section 7, we prove the Khintchine dichotomy for every probability measure on $\mathbb{R}$ satisfying certain equidistribution properties, yielding in particular A’.
Acknowledgements. The authors thank Nicolas de Saxcé for sharing his insight on random walks and Diophantine approximation, as well as Tushar Das, Shreyasi Datta, Dmitry Kleinbock, David Simmons, and Sanju Velani for helpful comments on a previous version of this paper. W.H. and H.Z. thank Barak Weiss for enlightening discussions. H.Z. thanks Ronggang Shi for his encouragement.
2. Preliminaries
In this section, we set up notations and collect basic facts that will be useful for the rest of the paper.
2.1. Notation and Conventions
Throughout this paper, $G=\operatorname{SL}_{2}(\mathbb{R})$, $\Lambda\subseteq G$ is a lattice, and $X=G/\Lambda$.
Metric. We fix a basis $(e_{-},e_{0},e_{+})$ of the Lie algebra $\mathfrak{g}=\operatorname{Lie}(G)$ given by
$e_{-}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}\,\,\,\,\,\,e_{0}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}\,\,\,\,\,\,e_{+}=\begin{pmatrix}0&1\\ 0&0\end{pmatrix}$
We assume throughout that $G$ is endowed with the unique right-invariant Riemannian metric for which $(e_{-},e_{0},e_{+})$ is orthonormal. This induces a distance on $G$ and the quotient $X$ that we denote by $\operatorname{dist}$ in both cases. Given $\rho>0$, we write $B_{\rho}$ to denote the open ball of radius $\rho>0$ centered at the neutral element $\operatorname{Id}$ in $G$. Then the open ball of radius $\rho$ centered at a point $x\in X$ can be written as $B_{\rho}x$.
The injectivity radius of $X$ at a point $x$ is
$\operatorname{inj}(x)=\sup\{\,\rho>0\,:\,\text{the map $B_{\rho}\rightarrow X,g\mapsto gx$ is injective}\,\}.$
Sobolev norms. Set $\Xi_{l}$ the words on the alphabet $\{e_{-},e_{0},e_{+}\}$ of length at most $l$. Each $D\in\Xi_{l}$ acts as a differential operator on the space of smooth functions $C^{\infty}(X)$. Given $f\in C^{\infty}(X)$, $k,l\in\mathbb{N}\cup\{\infty\}$, we set
${\mathcal{S}}_{k,l}(f)=\sum_{D\in\Xi_{l}}\|Df\|_{L^{k}},$
where $\|\cdot\|_{L^{k}}$ refers to the $L^{k}$-norm for the Haar probability measure on $X$. We let $B^{\infty}_{k,l}(X)$ denote the space of smooth functions $f$ on $X$ such that ${\mathcal{S}}_{k,l}(f)<\infty$.
Haar measure. Let $m_{G}$ denote the Haar measure on $G$ normalised so that the $G$-invariant Borel measure $m_{X}$ it induces on $X$ is a probability measure.
Driving measures $\lambda$ and $\mu$. Let $\operatorname{Aff}(\mathbb{R})^{+}$ denote the group of orientation preserving affine transformations of the real line. Denote by
$P=\{\,a(t)u(s):t>0,\,s\in\mathbb{R}\,\}\subseteq G$
the subgroup of upper triangular matrices with positive diagonal entries. For every $g\in P$, we let $\mathtt{r}_{g}\in\mathbb{R}_{>0}$ and $\mathtt{b}_{g}\in\mathbb{R}$ be the unique numbers such that
$g=a(\mathtt{r}_{g})^{-1}u(\mathtt{b}_{g})=\begin{pmatrix}\mathtt{r}_{g}^{-1/2}&\mathtt{r}_{g}^{-1/2}\mathtt{b}_{g}\\ 0&\mathtt{r}_{g}^{1/2}\end{pmatrix}.$
We identify $P$ with $\operatorname{Aff}(\mathbb{R})^{+}$ by mapping $g\in P$ with the similarity $s\mapsto\mathtt{r}_{g}s+\mathtt{b}_{g}$. This is an anti-isomorphism between the two groups.
Fix a probability measure $\lambda$ on $\operatorname{Aff}(\mathbb{R})^{+}$ with support $\operatorname{supp}\lambda$, denote by $\mu$ the corresponding probability measure on $P$ via the above anti-isomorphism. Throughout this paper, $\lambda$ and $\mu$ determine each other in this way. For $n\in\mathbb{N}$, we write $\lambda^{*n}=\lambda*\dotsm*\lambda$ to denote the $n$-fold convolution of $\lambda$ with itself, we define $\mu^{*n}$ similarly.
We assume that $\lambda$, and equivalently $\mu$, has a finite exponential moment. With our notations, this means there exists $\varepsilon>0$ such that
$\int_{P}|\mathtt{r}_{g}|^{\varepsilon}+|\mathtt{r}_{g}^{-1}|^{\varepsilon}+|\mathtt{b}_{g}|^{\varepsilon}\,\mathrm{d}\mu(g)<\infty.$
We assume that $\operatorname{supp}\lambda$ does not have a global fixed point in $\mathbb{R}$. This amounts to saying that $\operatorname{supp}\mu$ does not have two common fixed points on the projective line, or alternatively, that the matrices in $\operatorname{supp}\mu$ are not simultaneously diagonalizable.
Self-similar measure $\sigma$. Throughout this paper, we let $\sigma$ denote a probability measure on $\mathbb{R}$ that is $\lambda$-stationary, which means
$\sigma=\int_{\operatorname{Aff}(\mathbb{R})}\phi_{\star}\sigma\,\mathrm{d}\lambda(\phi).$
By a theorem of Bougerol-Picard [11, Theorem 2.5], the existence of such $\sigma$ is equivalent to the condition:
(2.1) $\int_{P}\log\mathtt{r}_{g}\,\mathrm{d}\mu(g)<0,$
i.e. the random walk on $\mathbb{R}$ driven by $\lambda$ is contractive in average. Moreover, provided existence, the measure $\sigma$ is uniquely determined by $\lambda$, see [11, Corollary 2.7].
Lyapunov exponent. Let $\operatorname{Ad}:G\rightarrow\operatorname{Aut}(\mathfrak{g})$ be the adjoint representation. We denote by $\ell$ the top Lyapunov exponent associated to $\operatorname{Ad}_{\star}\mu$. It is determined only by the diagonal terms and is equal to
(2.2) $\ell=-\int_{P}\log\mathtt{r}_{g}\,\mathrm{d}\mu(g)>0.$
[…]
2.2. Hölder-regularity of self-similar measures
We first recall that the measure $\sigma$ has finite moment of positive order and is Hölder-regular. We often refer to the second property as having positive dimension (at all scales).
Lemma 2.1 (Moment and Hölder-regularity of $\sigma$).
There exists $\gamma>0$ such that
$(i)\,\,\int_{\mathbb{R}}|s|^{\gamma}\,\mathrm{d}\sigma(s)<\infty,\qquad\quad(ii)\,\,\forall r>0,\,\sup_{s\in\mathbb{R}}\sigma(s+[-r,r])\ll r^{\gamma}.$
Proof.
Item (i) follows from [23, Proposition 5.1] due to Guivarc’h and Le Page and item (ii) is a consequence of [2, Theorem 2.12] due to Aoun and Guivarc’h. If one is only interested in self-similar measures arising from contracting IFS’s, then (i) is trivial and is not needed for our proof below and item (ii) dates back at least to a work of Feng–Lau [22, Proposition 2.2]. ∎
Given an integer $n\in\mathbb{N}$, denote by $\sigma^{(n)}$ the image measure of $\mu^{*n}$ under the map $g\in P\mapsto\mathtt{b}_{g}\in\mathbb{R}$. Equivalently, $\sigma^{(n)}={\lambda^{*n}}*\delta_{0}$, where $\delta_{0}$ denotes the Dirac measure at $0\in\mathbb{R}$. We show that the measures $\sigma^{(n)}$ have a uniformly finite positive moment, and uniform positive dimension above an exponentially small scale. For this, we first observe that $\sigma^{(n)}$ converges toward $\sigma$ at exponential rate. We denote by $\operatorname{Lip}(\mathbb{R})$ the space of bounded Lipschitz functions on $\mathbb{R}$ with the norm $\|f\|_{\operatorname{Lip}}=\|f\|_{\infty}+\sup_{s\neq t}\frac{|f(s)-f(t)|}{|s-t|}$.
Lemma 2.2.
There exists $\varepsilon>0$ such that for all $n\geq 0$, all $f\in\operatorname{Lip}(\mathbb{R})$, we have
$|\sigma^{(n)}(f)-\sigma(f)|\ll e^{-\varepsilon n}\|f\|_{\operatorname{Lip}}.$
Proof.
We may assume $\|f\|_{\operatorname{Lip}}=1$. Then we have,
$|\sigma^{(n)}(f)-\sigma(f)|$ $=|\lambda^{*n}*\delta_{0}(f)-\lambda^{*n}*\sigma(f)|$
$\leq\int_{\operatorname{Aff}(\mathbb{R})^{+}\times\mathbb{R}}|f(\phi(0))-f(\phi(s))|\,\mathrm{d}(\lambda^{*n}\otimes\sigma)(\phi,s)$
(2.3) $\leq\int_{\operatorname{Aff}(\mathbb{R})^{+}\times\mathbb{R}}\min(2,\,\mathtt{r}_{\phi}|s|)\,\mathrm{d}(\lambda^{*n}\otimes\sigma)(\phi,s),$
where $\mathtt{r}_{\phi}>0$ denotes the dilation factor in the affine map $\phi$, see (1.5). Using the principle of large deviations and that $\lambda$ is contracting in average (2.1), we have for $\varepsilon\lll 1$,
(2.4) $\lambda^{*n}\{\,\phi\,:\,\mathtt{r}_{\phi}>e^{-\ell n/2}\,\}\ll e^{-\varepsilon n}.$
On the other hand, up to taking smaller $\varepsilon$, Lemma 2.1(i) guarantees that
(2.5) $\sigma\{\,s\,:\,|s|>e^{\ell n/4}\,\}\ll e^{-\varepsilon n}.$
The claim follows from the combination of (2.2), (2.4), (2.5). ∎
We now deduce our claim on the measures $\sigma^{(n)}$.
Lemma 2.3 (Moment and Hölder-regularity of $\sigma^{(n)}$).
There exists $\gamma>0$ such that
$(i)\quad\sup_{n\geq 1}\int_{\mathbb{R}}|s|^{\gamma}\,\mathrm{d}\sigma^{(n)}(s)<\infty$
and
$(ii)\quad\forall n\geq 1,\,\forall r>e^{-n},\quad\sup_{s\in\mathbb{R}}\sigma^{(n)}(s+[-r,r])\ll r^{\gamma}.$
Proof.
Fix $\gamma,\varepsilon\in(0,1)$ as in Lemma 2.2.
For (i), we need to check that the map $t\mapsto\sup_{n}\sigma^{(n)}\{\,s:|s|\geq t\,\}$ has polynomial decay as $t\to+\infty$. Given $t>2$, Lemma 2.2 and Lemma 2.1(i) imply that for every $n\geq 0$, one has
$\sigma^{(n)}\{\,s\,:\,|s|\geq t\,\}\leq\sigma\{\,s\,:\,|s|\geq t-1\,\}+O(e^{-\varepsilon n})\ll t^{-\gamma}+e^{-\frac{\varepsilon}{2}n}.$
Let $R>0$ be a parameter. The above justifies that uniformly in $n$, we have polynomial decays of tail probabilities of $\sigma^{(n)}$ for $t\leq e^{Rn}$. Taking $R\ggg 1$, the exponential moment assumption on $\lambda$ takes care of the case $t>e^{Rn}$, using the observation
$\sigma^{(n)}\{\,s\,:\,|s|\geq t\,\}\leq\lambda^{\otimes n}\Bigl{\{}(\phi_{1},\dots,\phi_{n})\,:\,n\prod_{k=1}^{n}\max(1,\mathtt{r}_{\phi_{k}},|\mathtt{b}_{\phi_{k}}|)\geq t\Bigr{\}}$
and the Markov inequality. This justifies (i), with a potentially smaller value of $\gamma$.
Let us check (ii). For $n\geq 0$, $s\in\mathbb{R}$ and $r\geq e^{-\frac{\varepsilon}{2}n}$, Lemma 2.2 guarantees
$\sigma^{(n)}([s-r,s+r])\leq\sigma([s-2r,s+2r])+O(e^{-\frac{\varepsilon}{2}n})\ll r^{\gamma},$
whence the claim (with $\frac{\varepsilon}{2}\gamma$ in place of $\gamma$ to treat all scales above $e^{-n}$). ∎
2.3. Recurrence of the random walk
We recall the following result of non-escape of mass for the $\mu$-walk on $X$.
Proposition 2.4 (Effective recurrence on $X$).
There exist constants $c,c^{\prime}>0$ depending on $\mu$ only such that for every $x\in X$, $n\in\mathbb{N}$, and $\rho>0$, we have
$\mu^{*n}*\delta_{x}\{\operatorname{inj}<\rho\}\ll({\operatorname{inj}^{-c}(x)}e^{-c^{\prime}n}+1)\rho^{c}.$
This result is proven in [42] in a more general setting. In our context, we present a brief proof for completeness.
Proof of Proposition 2.4.
[…]
(2.6) $\mu^{*n}*\delta_{x}(\operatorname{sys}^{-c})\ll e^{-c^{\prime}n}\operatorname{sys}(x)^{-c}+1.$
Let $k,R\geq 2$, $c\in(0,1)$ be parameters to be specified later depending on $\mu$. Recall also the notation $\ell=-\int_{G}\log\mathtt{r}_{g}\,\mathrm{d}\mu(g)>0$. Given $x\in X$, we write $w_{x}=v_{x}/\|v_{x}\|$ and set
$E:=\{g\in G\,:\,\|g\|>R^{k}\}\quad\text{and}\quad F_{x}:=\{g\in G\,:\,\|gw_{x}\|<e^{k\ell/4}\}.$
Assume $\operatorname{sys}(x)<R^{-k}$. Note that for any $g\in G$, we have $\operatorname{sys}(gx)\geq\|g\|^{-1}\operatorname{sys}(x)$, and also $\operatorname{sys}(gx)=\|gv_{x}\|\geq e^{k\ell/4}\operatorname{sys}(x)$ if $g\notin(E\cup F_{x})$. It follows that
(2.7) $\mu^{*k}*\delta_{x}(\operatorname{sys}^{-c})$ $\leq\int_{E\cup F_{x}}\|g\|^{c}\,\mathrm{d}\mu^{*k}(g)\operatorname{sys}(x)^{-c}+e^{-k\ell c/4}\operatorname{sys}(x)^{-c}.$
Moreover, using the Cauchy-Schwarz inequality and the submultiplicativity of the norm, we have
(2.8) $\int_{E\cup F_{x}}\|g\|^{c}\,\mathrm{d}\mu^{*k}(g)\leq\mu^{*k}(E\cup F_{x})^{1/2}M_{2c}(\mu)^{k/2}$
where, given $s>0$, we write $M_{s}(\mu):=\int_{G}\|g\|^{s}\,\mathrm{d}\mu(g)$ the $s$-exponential moment of $\mu$.
We claim that for some $\alpha=\alpha(\mu)>0$, and up to taking parameters $k,R\ggg 1$, we may further bound
(2.9) $\mu^{*k}(E\cup F_{x})\leq e^{-k\alpha}.$
Indeed, on the one hand, by the Markov inequality, for any $\varepsilon>0$,
$\mu^{*k}(E)\leq\left(M_{\varepsilon}(\mu)R^{-\varepsilon}\right)^{k}$
whence the claim on $\mu^{*k}(E)$ provided $R\ggg 1$. On the other hand, writing $w_{x}=(w_{x,1},w_{x,2})\in\mathbb{R}^{2}$, we have, by examining the first coordinate,
$\lVert gw_{x}\rVert\geq\mathtt{r}^{-1/2}_{g}\lvert w_{x,1}+\mathtt{b}_{g}w_{x,2}\rvert,$
whence
$\mu^{*k}(F_{x})$ $\leq\mu^{*k}\{g\in G\,:\,|w_{x,1}+\mathtt{b}_{g}w_{x,2}|<\mathtt{r}^{1/2}_{g}e^{k\ell/4}\}$
$\leq\mu^{*k}\{g\in G\,:\,|w_{x,1}+\mathtt{b}_{g}w_{x,2}|<e^{-k\ell/5}\}+Ce^{-k\varepsilon},$
where $C,\varepsilon>0$ are constants depending on $\mu$, arising from the large deviation principle for $\log\mathtt{r}_{g}$. Now the desired bound on $\mu^{*k}(F_{x})$ follows provided $k\ggg 1$: when $|w_{x,2}|\leq e^{-k\ell/10}$ it is a consequence of the moment estimate on $\mathtt{b}_{g}$, see Lemma 2.3(i), and otherwise it stems from its non-concentration estimate, see Lemma 2.3(ii). This finishes the proof of (2.9).
Combining (2.8) and (2.9), we get
(2.10) $\int_{E\cup F_{x}}\|g\|^{c}\,\mathrm{d}\mu^{*k}(g)\leq e^{-k\alpha/2}M_{2c}(\mu)^{k/2}\leq e^{-k\alpha/3},$
where the second inequality assumes additionally $c\lll 1$. Plugging (2.10) into (2.7), we have shown in the end that for $k,R\ggg 1$, and $c\lll 1$, for any $x\in X$ with $\operatorname{sys}(x)<R^{-k}$, we have
(2.11) $\mu^{*k}*\delta_{x}(\operatorname{sys}^{-c})\leq r\operatorname{sys}(x)^{-c}$
where $r=r(\mu,k,c)\in(0,1)$. On the other hand, if $\operatorname{sys}(x)\geq R^{-k}$ we have by finite exponential moment of $\mu$ and the inequality $\operatorname{sys}(gx)\geq\|g\|^{-1}\operatorname{sys}(x)$ that
(2.12) $\mu^{*k}*\delta_{x}(\operatorname{sys}^{-c})\leq C$
where $C=C(\mu,R,k,c)>0$.
Together (2.11) and (2.12) yield for every $x\in X$,
(2.13) $\mu^{*k}*\delta_{x}(\operatorname{sys}^{-c})\leq r\operatorname{sys}(x)^{-c}+C.$
The claim (2.6) follows from (2.13) inductively, setting $c^{\prime}=|\log r|$ and using that the sum $\sum_{k\geq 0}r^{k}C$ is finite. This finishes the proof of the proposition in the case where $\Lambda=\operatorname{SL}_{2}(\mathbb{Z})$.
For general $\Lambda$, note we may assume that $\Lambda$ is not cocompact, note also that each cusp of $X$ has a neighborhood that is isometric to a neighborhood of the cusp in $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$. Hence the behavior of the walk outside of a compact set is captured by the behavior in $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$, more precisely the proof of (2.6) (with $\operatorname{inj}$ in place of $\operatorname{sys}$ here) still goes through. ∎
3. Positive dimension
We show that the $n$-step distribution of the $\mu$-walk starting from a point $x$ acquires positive dimension at an exponential rate, tempered by the possibility that $x$ may be high in the cusp.
Proposition 3.1 (Positive dimension).
There exists $A,\kappa>0$ such that for every $x\in X$, $\rho>0$, $n\geq|\log\rho|+A|\log\operatorname{inj}(x)|$, we have
(3.1) $\forall y\in X,\quad\mu^{*n}*\delta_{x}(B_{\rho}y)\ll\rho^{\kappa}.$
Proof.
Let $\kappa>0$ be a parameter to specify below. Let $\rho\in(0,1/10)$, $n\geq|\log\rho|$, $x,y\in X$, and assume
(3.2) $\mu^{*n}*\delta_{x}(B_{\rho}y)\geq\rho^{\kappa}.$
Fix $\alpha=\alpha(\mu)>0$ so small that, setting $m=\lfloor\alpha|\log\rho|\rfloor$, we have
$\max_{h\in\operatorname{supp}\mu}\lVert\operatorname{Ad}(h)\rVert^{m}\leq\rho^{-1/2}.$
Writing $\mu^{*n}*\delta_{x}=\mu^{*m}*\mu^{*(n-m)}*\delta_{x}$, Equation (3.2) implies that
$\mu^{*(n-m)}*\delta_{x}(Z)\geq\rho^{2\kappa}\,\,\text{ where }\,\,Z:=\{z:\mu^{*m}*\delta_{z}(B_{\rho}y)\geq\rho^{2\kappa}\},$
up to assuming $\rho$ small enough in terms of $\kappa$. Indeed,
$\rho^{\kappa}\leq\mu^{*m}*\mu^{*(n-m)}*\delta_{x}(B_{\rho}y)$ $=\int_{Z\cup(X\smallsetminus Z)}\mu^{*m}*\delta_{z}(B_{\rho}y)\,\mathrm{d}\mu^{*(n-m)}*\delta_{x}(z)$
$\leq\rho^{2\kappa}+\mu^{*(n-m)}*\delta_{x}(Z),$
so we obtain $\mu^{*(n-m)}*\delta_{x}(Z)\geq\rho^{\kappa}-\rho^{2\kappa}\geq\rho^{2\kappa}$, provided $\rho\leq 2^{-1/\kappa}$.
We now show that $Z$ must be included in a small neighborhood of the cusp. Fix $z\in Z$. By the choice of $m$ and the lower bound $\mu^{*m}*\delta_{z}(B_{\rho}y)\geq\rho^{2\kappa}$, we have
(3.3) ${\mu^{*m}}\{\,g\,:\,g^{-1}y\in B_{\rho^{1/2}}z\,\}\geq\rho^{2\kappa}.$
On the other hand, by Lemma 2.3(i), we have for some $\gamma=\gamma(\mu)>0$, and $\rho\lll 1$,
(3.4) ${\mu^{*m}}\{\,g\,:\,|\mathtt{b}_{g}|\leq\rho^{-4\gamma^{-1}\kappa}\,\}\geq 1-\rho^{3\kappa}.$
By finite exponential moment of $\mu$, there exists also $\varepsilon>0$ depending only on $\mu$ such that
(3.5) $\mu^{*m}\{g\,:\,\log\mathtt{r}_{g}\in[(\ell-1)m,(\ell+1)m]\}\geq 1-\rho^{\alpha\varepsilon}.$
Let $C>1$ be a parameter to be specified below depending on $\mu$ only. Cutting the intervals $[-\rho^{-4\gamma^{-1}\kappa},\rho^{-4\gamma^{-1}\kappa}]$ and $[(\ell-1)m,(\ell+1)m]$ into subintervals of length $\rho^{C\kappa}$, then using the pigeonhole principle, we deduce from (3.3) (3.4), (3.5) that there exists $(r_{0},b_{0})\in\mathbb{R}^{2}$ with $r_{0}>0$, $|b_{0}|\leq\rho^{-4\gamma^{-1}\kappa}$ such that the set
$E:=\{g\,:\,g^{-1}y\in B_{\rho^{1/2}}z\text{ and }|1-\mathtt{r}_{g}r^{-1}_{0}|\leq\rho^{C\kappa}\text{ and }|\mathtt{b}_{g}-b_{0}|\leq\rho^{C\kappa}\}$
has $\mu^{*m}$-measure
(3.6) $\mu^{*m}(E)\geq\frac{\rho^{2\kappa}-\rho^{3\kappa}-\rho^{\alpha\varepsilon}}{\lceil 2\rho^{-4\gamma^{-1}\kappa}\rho^{-C\kappa}\rceil\lceil 2m\rho^{-C\kappa}\rceil}\geq\rho^{4C\kappa}$
where the last lower bound assumes $C\geq 4\gamma^{-1}$, $3\kappa\leq\alpha\varepsilon$, and $\rho\lll_{\kappa}1$.
Consider $g_{1},g_{2}\in E$, in particular $\operatorname{dist}(g_{1}^{-1}y,g_{2}^{-1}y)\ll\rho^{1/2}$. Recalling $g^{-1}_{i}=u(-\mathtt{b}_{g_{i}})a(\mathtt{r}_{g_{i}})$ and setting $z_{2}=a(\mathtt{r}_{g_{2}})y$, the inequality can be rewritten as
$\operatorname{dist}(u(-\mathtt{b}_{g_{1}})a(\mathtt{r}_{g_{1}}\mathtt{r}^{-1}_{g_{2}})z_{2},u(-\mathtt{b}_{g_{2}})z_{2})\ll\rho^{1/2}.$
In view of the condition $|\mathtt{b}_{g_{2}}|\leq\rho^{-4\gamma^{-1}\kappa}$, this implies
(3.7) $\operatorname{dist}(u(\mathtt{b}_{g_{2}}-\mathtt{b}_{g_{1}})a(\mathtt{r}_{g_{1}}\mathtt{r}^{-1}_{g_{2}})z_{2},z_{2})\ll\rho^{1/2-O(\kappa)}\ll\rho^{1/4}$
provided $\kappa\lll 1$. Moreover, the combination of (3.6) and the non-concentration estimate from Lemma 2.3(ii) allows us to choose the elements $g_{1},g_{2}\in E$ such that
(3.8) $|\mathtt{b}_{g_{2}}-\mathtt{b}_{g_{1}}|\geq\rho^{\gamma^{-1}5C\kappa},$
provided $\kappa\lll_{C}1$ to guarantee $\rho^{\gamma^{-1}5C\kappa}\geq e^{-m}$ as required by Lemma 2.3(ii). Observing that
$\operatorname{dist}(u(\mathtt{b}_{g_{2}}-\mathtt{b}_{g_{1}})a(\mathtt{r}_{g_{1}}\mathtt{r}^{-1}_{g_{2}}),\operatorname{Id})$ $\simeq|\mathtt{b}_{g_{2}}-\mathtt{b}_{g_{1}}|+|1-\mathtt{r}_{g_{1}}\mathtt{r}^{-1}_{g_{2}}|$ $\in[\rho^{\gamma^{-1}5C\kappa},4\rho^{C\kappa}]$
we deduce from (3.7) that for $\kappa\lll_{C}1$ so that $\rho^{\gamma^{-1}5C\kappa}\geq\rho^{1/5}$, we have
$\operatorname{inj}(z_{2})\ll\rho^{C\kappa}.$
Recalling $\operatorname{dist}(z,z_{2})\leq\operatorname{dist}(z,g_{2}^{-1}y)+\operatorname{dist}(g_{2}^{-1}y,z_{2})\ll\rho^{-4\gamma^{-1}\kappa}$, we get
$\operatorname{inj}(z)\ll\rho^{\frac{C}{2}\kappa}$
provided $C\ggg 1$.
In conclusion, we have shown that for $C\ggg 1$, for $\kappa\lll_{C}1$, for $\rho\lll_{\kappa}1$, and $n\geq m\simeq\alpha|\log\rho|$, we have
${\mu^{*(n-m)}}*\delta_{x}\{\operatorname{inj}\leq\rho^{C\kappa}\}\geq\rho^{2\kappa}$
By the effective recurrence statement from Proposition 2.4, this is absurd if $n-m\ggg{|\log\operatorname{inj}(x)|}$. This concludes the proof of the proposition. ∎
4. Dimensional bootstrap
In this section, we explain how the positive dimension estimate for $\mu^{*n}*\delta_{x}$ established in the previous section can be upgraded to a high-dimension estimate, up to applying more convolutions by $\mu$ and throwing away some small part of the measure. The notion of robust measures from [44] is well adapted to our purpose.
Definition 4.1 (Robustness).
Let $\alpha,\tau>0$, $I\subseteq(0,1]$. A Borel measure $\nu$ on $X$ is $(\alpha,{\mathcal{B}}_{I},\tau)$-robust if $\nu$ can be decomposed as the sum of two Borel measures $\nu=\nu^{\prime}+\nu^{\prime\prime}$ such that $\nu^{\prime\prime}(X)\leq\tau$, and $\nu^{\prime}$ satisfies
(4.1) $\nu^{\prime}\{\operatorname{inj}<\sup I\}=0,$
as well as for all $\rho\in I$, $y\in X$,
(4.2) $\nu^{\prime}(B_{\rho}y)\leq\rho^{3\alpha}$
Condition (4.2) means that $\nu^{\prime}$ has normalised dimension at least $\alpha$ with respect to balls of radius $\rho$. Note that in this definition, $\nu$ may not be a probability measure, this flexibility will be convenient for us.
The goal of the section is to establish the following high dimension estimate.
Proposition 4.2 (High dimension).
Let $\kappa\in(0,1/10)$. For $\eta,\rho\lll_{\kappa}1$ and for all $n\ggg_{\kappa}|\log\rho|+|\log\operatorname{inj}(x)|$, the measure $\mu^{*n}*\delta_{x}$ is $(1-\kappa,B_{\rho},\rho^{\eta})$-robust.
4.1. Multislicing
The proof of Proposition 4.2 relies on a multislicing estimate established in [5]. We recall the case of interest in our context.
We consider $\Theta$ a measurable space. We let $(\varphi_{\theta})_{\theta\in\Theta}$ denote a measurable family of $C^{2}$-embeddings $\varphi_{\theta}:B^{\mathbb{R}^{3}}_{1}\rightarrow\mathbb{R}^{3}$, and $(L_{\theta})_{\theta\in\Theta}$ a measurable family of constants $L_{\theta}\geq 1$ such that each map $\varphi_{\theta}$ is $L_{\theta}$-bi-Lipschitz:
$\forall x,y\in B^{\mathbb{R}^{3}}_{1},\quad\frac{1}{L_{\theta}}\|x-y\|\leq\|\varphi_{\theta}(x)-\varphi_{\theta}(y)\|\leq L_{\theta}\|x-y\|$
and has second order derivatives bounded by $L_{\theta}$:
$\forall x,h\in B^{\mathbb{R}^{3}}_{1},\quad\lVert\varphi_{\theta}(x+h)-\varphi_{\theta}(x)-(D_{x}\varphi_{\theta})(h)\rVert\leq L_{\theta}\lVert h\rVert^{2}.$
Given $\rho>0$, we denote by ${\mathcal{D}}_{\rho}$ (resp. ${\mathcal{R}}_{\rho}$) the collection of subsets of $\mathbb{R}^{3}$ that are translates of the $\rho$-cube $[0,\rho]^{3}$ (resp. the rectangle $R_{\rho}:=[0,1]e_{1}+[0,\rho^{1/2}]e_{2}+[0,\rho]e_{3}$).
We will also need to measure the angle between subspaces in $\mathbb{R}^{3}$. For each $k=1,2,3$, endow $\wedge^{k}\mathbb{R}^{3}$ with the unique Euclidean structure with respect to which the standard basis is orthonormal. Given subspaces $V,W\subseteq\mathbb{R}^{3}$, we set
$\operatorname{d_{\measuredangle}}(V,W)=\|v\wedge w\|$
where $v,w$ are unit vectors in $\wedge^{*}\mathbb{R}^{3}$ spanning respectively the lines $\wedge^{\dim V}V$, $\wedge^{\dim W}W$.
The multislicing estimate presented in 4.3 below is a special case of [5, Corollary 2.2]. It takes as input a Borel measure $\nu$ on the unit ball $B^{\mathbb{R}^{3}}_{1}$ that has normalized dimension at least $\alpha$ with respect to balls of radius above $\rho$. The output is a dimensional gain when the balls are replaced by (non-linear) rectangles of the form $(\varphi^{-1}_{\theta}(x+R_{\rho}))_{x\in\mathbb{R}^{3}}$ provided $\theta$ is chosen almost typically via a probability measure for which $\varphi_{\theta}$ satisfies suitable bounds on the derivatives as well as non-concentration estimates. The proof relies on Shmerkin’s nonlinear version [44] of Bourgain’s discretized projection theorem [12], local conditioning arguments, and a submodularity inequality for covering numbers.
Theorem 4.3 (Multislicing [5]).
Given $\kappa\in(0,1/2)$, there exist $\varepsilon=\varepsilon(\kappa)>0$ and $\rho_{0}=\rho_{0}(\kappa)>0$ such that the following holds for all $\rho\in(0,\rho_{0}]$.
Let $\nu$ be a Borel measure on $B^{\mathbb{R}^{3}}_{1}$ satisfying: $\exists\alpha\in(\kappa,1-\kappa)$, $\forall r\in[\rho,\rho^{\varepsilon}]$,
$\sup_{Q\in{\mathcal{D}}_{r}}\nu(Q)\leq r^{3\alpha}.$
Let $\Xi$ be a probability measure on $\Theta$ satisfying:
  • (i)
    $\Xi\{\,\theta\in\Theta:L_{\theta}\leq\rho^{-\varepsilon}\,\}=1.$
  • (ii)
    $\forall k\in\{1,2\}$, $\forall x\in B^{\mathbb{R}^{3}}_{1}$, $\forall r\in[\rho,\rho^{\varepsilon}]$, $\forall W\in\operatorname{Gr}(\mathbb{R}^{3},3-k)$,
    $\Xi\{\,\theta\in\Theta:\operatorname{d_{\measuredangle}}((D_{x}\varphi_{\theta})^{-1}V_{k},W)\leq r\,\}\leq r^{\kappa},$
    where $V_{k}=\operatorname{span}_{\mathbb{R}}(e_{1},\dotsc,e_{k})$.
Then there exists ${\mathcal{F}}\subseteq\Theta$ such that $\sigma({\mathcal{F}})\geq 1-\rho^{\varepsilon}$ and for every $\theta\in{\mathcal{F}}$, there exists $A_{\theta}\subseteq B^{\mathbb{R}^{3}}_{1}$ with $\nu(A_{\theta})\geq 1-\rho^{\varepsilon}$ and satisfying
$\sup_{Q\in{\mathcal{R}}_{\rho}}\nu_{|A_{\theta}}(\varphi_{\theta}^{-1}Q)\leq\rho^{\frac{3}{2}\alpha+\varepsilon}.$
4.2. Straightening charts
In order to apply the multislicing estimates from 4.3, we need special macroscopic charts in which the preimage by $g\in\operatorname{supp}\mu^{*n}$ of a ball looks like a rectangle. The goal of the present subsection is to define those charts.
Recall that $\mathfrak{g}$ admits the rootspace decomposition
$\mathfrak{g}=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{+},$
where
$\mathfrak{g}_{-}=\mathbb{R}e_{-},\quad\mathfrak{g}_{0}=\mathbb{R}e_{0}\quad\text{and}\quad\mathfrak{g}_{+}=\mathbb{R}e_{+}.$
We then define $\Psi:\mathfrak{g}\rightarrow G$ by the formula: $\forall(v_{-},v_{0},v_{+})\in\mathfrak{g}_{-}\times{\mathfrak{g}_{0}}\times\mathfrak{g}_{+}$,
$\Psi(v_{-}+v_{0}+v_{+})=\exp(v_{-})\exp(v_{0})\exp(v_{+}).$
Recall also the notation
$a(t)=\begin{pmatrix}t^{1/2}&0\\ 0&t^{-1/2}\end{pmatrix}.$
The next lemma tells us that, in the chart $\Psi$, the image of a ball $B_{\rho}$ in $G$ by some diagonal element $a(t)$ with small $t>0$ is included in a rectangle whose volume is comparable.
Lemma 4.4.
There is an absolute constant $r_{0}>0$ such that for any $t,\rho\in(0,1)$ with $|t^{-1}\rho|\leq r_{0}$ and any $h\in G$, there is $w\in\mathfrak{g}$ such that
(4.3) $\{\,v\in B^{\mathfrak{g}}_{r_{0}}:\Psi(v)\in a(t)B_{\rho}h\,\}\subseteq\operatorname{Ad}(a(t))B^{\mathfrak{g}}_{10\rho}+w.$
This result is a particular case of [5, Lemma 4.10]. We give a shorter proof in our context for completeness.
Proof.
Fix a vector $w$ in the left hand side of (4.3). If $v$ belongs to the left hand side of (4.3) as well, then by the triangle inequality, we have
$\Psi(v)\in a(t)B_{2\rho}a(t^{-1})\Psi(w).$
We can choose $r_{0}>0$ small so that the image of $\Psi$ contains $B_{2r_{0}}$. In particular, using that conjugation commutes with the exponential map, there is $u=u_{-}+u_{0}+u_{+}\in\mathfrak{g}$ such that
$u_{-}\in B^{\mathfrak{g}_{-}}_{t^{-1}\rho},\quad u_{0}\in B^{\mathfrak{g}_{0}}_{\rho},\quad u_{+}\in B^{\mathfrak{g}_{+}}_{t\rho}$
and
$\Psi(v)=\Psi(u)\Psi(w).$
Consider $x,y,s\in\mathbb{R}$ with $s\neq 0$ and $1+xy\neq 0$. Note that we have in $G$ the following equality
(4.4) $\begin{pmatrix}1&x\\ 0&1\end{pmatrix}\begin{pmatrix}1&0\\ y&1\end{pmatrix}\begin{pmatrix}s&0\\ 0&s^{-1}\end{pmatrix}=\begin{pmatrix}1&0\\ y^{\prime}&1\end{pmatrix}\begin{pmatrix}s^{\prime}&0\\ 0&s^{\prime-1}\end{pmatrix}\begin{pmatrix}1&x^{\prime}\\ 0&1\end{pmatrix}$
where
$x^{\prime}=\frac{x}{(1+xy)s^{2}},\quad s^{\prime}=(1+xy)s,\quad y^{\prime}=\frac{y}{1+xy}.$
Similarly,
(4.5) $\begin{pmatrix}s&0\\ 0&s^{-1}\end{pmatrix}\begin{pmatrix}1&0\\ y&1\end{pmatrix}=\begin{pmatrix}1&0\\ s^{-2}y&1\end{pmatrix}\begin{pmatrix}s&0\\ 0&s^{-1}\end{pmatrix}.$
Observe that
$\Psi(v)=\Psi(u)\Psi(w)=\exp(u_{-})\exp(u_{0})\exp(u_{+})\exp(w_{-})\exp(w_{0})\exp(w_{+}).$
Assuming $r_{0}\lll 1$, applying (4.4) to the factor $\exp(u_{+})\exp(w_{-})\exp(w_{0})$ then (4.5) to the factor $\exp(u_{0})\exp(w^{\prime}_{-})$, we obtain
$\Psi(v)\in\exp(B^{\mathfrak{g}_{-}}_{10t^{-1}\rho}+w_{-})\exp(B^{\mathfrak{g}_{0}}_{10\rho}+w_{0})\exp(B^{\mathfrak{g}_{+}}_{10t\rho}+w_{+}).$
Noting that $\Psi$ is injective (by direct computation again), this finishes the proof. ∎
In view of Lemma 4.4 and the formula
$g=a(\mathtt{r}^{-1}_{g})u(\mathtt{b}_{g}),$
we define a family of straightening charts $(\varphi_{\theta})$ as follows. Let
$\Theta=u(\mathbb{R}).$
Fix $r_{1}>0$ such that $\Psi$ is a smooth diffeomorphism between $B^{\mathfrak{g}}_{r_{1}}$ and a neighborhood $\mathcal{O}$ of $\operatorname{Id}\in G$. Given $\theta\in\Theta$, define $\varphi_{\theta}:\mathcal{O}\rightarrow\mathfrak{g}$ by
$\varphi_{\theta}:=\operatorname{Ad}(\theta^{-1})\circ(\Psi_{|B^{\mathfrak{g}}_{r_{1}}})^{-1}.$
Using that $\psi$ commutes with conjugation, we have the alternative formula $\varphi_{\theta}=(\Psi_{|\operatorname{Ad}(\theta^{-1})B^{\mathfrak{g}}_{r_{1}}})^{-1}\circ{\mathscr{C}}_{\theta^{-1}}$ where ${\mathscr{C}}_{\theta^{-1}}:h\mapsto\theta^{-1}h\theta$. Note that $\varphi_{\theta}$ is $L_{\theta}$-bi-Lipschitz and satisfies $\|\varphi_{\theta}\|_{C^{2}}\leq L_{\theta}$ for some quantity
(4.6) $L_{\theta}:=L\|\theta\|^{4}.$
where $L>1$ is a constant depending only on $r_{1}$.
Given an element $g\in P$, write
$g^{-1}=\theta_{g}a(\mathtt{r}_{g})$
with
(4.7) $\theta_{g}:=u(-\mathtt{b}_{g})\in\Theta.$
Lemma 4.4 tells us that for any $h\in G$, $\varphi_{\theta_{g}}(g^{-1}B_{\rho}h)$ is essentially an additive translate of the rectangle $\operatorname{Ad}(a(\mathtt{r}_{g}))B^{\mathfrak{g}}_{\rho}$ provided that $\mathtt{r}_{g}\in(0,1)$, and that both $g^{-1}B_{\rho}h$ and $a(\mathtt{r}_{g})B_{\rho}h\theta_{g}$ sit inside a prescribed (macroscopic) neighborhood of the identity.
4.3. Control of the charts
We now check that the charts ($\varphi_{\theta}$) from the previous section satisfy distortion control and non-concentration estimates. These will be required in order to apply 4.3 in the next section. The constants $r_{0}$ from Lemma 4.4 and $r_{1}>0$ in the definition of $\varphi_{\theta}$ are assumed fixed in a canonical way (so that dependence on them does not appear in subscript asymptotic notations).
Recall $L_{\theta}$ and $\theta_{g}$ are respectively defined in (4.6) and (4.7).
Lemma 4.5 (Distortion control).
Given $\varepsilon>0$, there exists $\gamma=\gamma(\mu,\varepsilon)>0$ such that for $n\ggg_{\varepsilon}1$, we have
$\mu^{*n}\{\,g\,:\,L_{\theta_{g}}>e^{\varepsilon n}\,\}\leq e^{-\gamma n}.$
Proof.
This is a direct consequence of $L_{\theta_{g}}\ll\|\theta_{g}\|^{4}\ll(1+|\mathtt{b}_{g}|)^{4}$ and Lemma 2.3(i), stating that the variable $\mathtt{b}_{g}$, where $g\overset{law}{\sim}\mu^{*n}$, has a moment of positive order that is bounded independently of $n$. ∎
Set $\mathfrak{g}_{-,0}=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}$.
Lemma 4.6 (Non-concentration).
There exists a constant $\kappa>0$ such that for $n\ggg 1$, $h\in B_{r_{0}}$ and $\rho\geq e^{-n}$, we have
$\forall W\in\operatorname{Gr}(T_{h}G,2),\quad\mu^{*n}\{\,g\,:\,\operatorname{d_{\measuredangle}}((D_{h}\varphi_{\theta_{g}})^{-1}\mathfrak{g}_{-},W)\leq\rho\,\}\ll\rho^{\kappa},$
and
$\forall W\in\operatorname{Gr}(T_{h}G,1),\quad\mu^{*n}\{\,g\,:\,\operatorname{d_{\measuredangle}}((D_{h}\varphi_{\theta_{g}})^{-1}\mathfrak{g}_{-,0},W)\leq\rho\,\}\ll\rho^{\kappa}.$
Proof.
Unwrapping definitions, we observe that the distribution of subspaces $h\mapsto(D_{h}\varphi_{\theta})^{-1}\mathfrak{g}_{-}$ is right-invariant and coincides with $\operatorname{Ad}(\theta)\mathfrak{g}_{-}$ at the identity. The same holds for $(D_{h}\varphi_{\theta})^{-1}\mathfrak{g}_{-,0}$. Recalling that $\theta_{g}=u(-\mathtt{b}_{g})$ and $\sigma^{(n)}$ is the law of $\mathtt{b}_{g}$ as $g\overset{law}{\sim}\mu^{*n}$, we are then led to proving the following non-concentration estimates:
$\sup_{W\in\operatorname{Gr}(\mathfrak{g},2)}\sigma^{(n)}\{s\,:\,\operatorname{d_{\measuredangle}}(\operatorname{Ad}(u(-s))\mathfrak{g}_{-},W)\leq\rho\}\ll\rho^{\kappa},\quad\text{and}$
$\sup_{W\in\operatorname{Gr}(\mathfrak{g},1)}\sigma^{(n)}\{s\,:\,\operatorname{d_{\measuredangle}}(\operatorname{Ad}(u(-s))\mathfrak{g}_{-,0},W)\leq\rho\}\ll\rho^{\kappa}.$
Let us check the first estimate, where $W\in\operatorname{Gr}(\mathfrak{g},2)$. Set $e_{-,0}=e_{-}\wedge e_{0}$, $e_{-,+}=e_{-}\wedge e_{+}$, $e_{0,+}=e_{0}\wedge e_{+}$. Write $\wedge^{2}W=\mathbb{R}(ae_{-,0}+be_{-,+}+ce_{0,+})$ where $a,b,c\in\mathbb{R}$ with $\max(|a|,|b|,|c|)=1$. Then direct computation yields for any $s\in\mathbb{R}$,
$\operatorname{d_{\measuredangle}}(\operatorname{Ad}(u(-s))\mathfrak{g}_{-},W)\simeq|as^{2}-bs-c|.$
In the case $\max(|a|,|b|)\leq\rho^{1/10}$, we must have $|c|=1$ while $|as^{2}-bs|\leq 1/2$ with $\sigma^{(n)}$-probability at least $1-O(\rho^{\gamma^{-1}/20})$ for some $\gamma=\gamma(\mu)$, see (Lemma 2.3(i)), whence the claim. In the case $\max(|a|,|b|)\geq\rho^{1/10}$, the desired non-concentration follows from combining Lemma 4.7 below and Lemma 2.3 (items (i) and (ii)), the latter claiming that $\sigma^{(n)}$ has finite moment and positive dimension at scales above $e^{-n}$, uniformly in $n$.
The second estimate is similar: writing $W=\mathbb{R}(ae_{+}+be_{0}+ce_{-})$ where $\max(\lvert a\rvert,\lvert b\rvert,\lvert c\rvert)=1$, we find $\operatorname{d_{\measuredangle}}(\operatorname{Ad}(u(-s))\mathfrak{g}_{-},W)\simeq\lvert a-2bs-cs^{2}\rvert$. ∎
The next elementary lemma was used in the proof Lemma 4.6 of above.
Lemma 4.7.
Let $(a,b,c)\in\mathbb{R}^{3}$ and $\varepsilon\in(0,1)$ with $\max(|a|,|b|)\geq\varepsilon$, let $\rho\in(0,1)$. The set $E:=\{s\in[-\rho^{-1/4},\rho^{-1/4}]\,:\,|as^{2}+bs+c|\leq\rho\}$ is included in at most two balls of radius $8\varepsilon^{-1}\rho^{1/4}$.
Proof.
We may assume $8\varepsilon^{-1}\rho^{1/4}\leq 2\rho^{-1/4}$, i.e. $\rho^{1/2}\leq\varepsilon/4$. If $s_{1},s_{2}\in E$, then $|as_{1}^{2}+bs_{1}-as_{2}^{2}+bs_{2}|\leq 2\rho$, i.e.
$|(s_{1}-s_{2})(b+a(s_{1}+s_{2}))|\leq 2\rho.$
Then either $|s_{1}-s_{2}|\leq 2\rho^{1/2}$ or $|b+a(s_{1}+s_{2})|\leq\rho^{1/2}$. In the second case, the conditions $\max(|a|,|b|)\geq\varepsilon$ and $\rho^{1/2}\leq\varepsilon/4$ force $a\geq\rho^{1/4}\varepsilon/4$, then $s_{1}$ belongs to the ball of radius $8\varepsilon^{-1}\rho^{1/4}$ and center $(-a^{-1}b-s_{2})$. ∎
4.4. Dimension increment
In this subsection, we apply the multislicing estimate from 4.3 to show that convolution by a well chosen power of $\mu$ increases dimensional properties of a measure at a given scale.
Proposition 4.8 (Dimension increment).
Let $\kappa,\varepsilon,\rho\in(0,1/10)$, $\alpha\in{[\kappa,1-\kappa]}$, $\tau\geq 0$ be some parameters. Consider on $X$ a Borel measure $\nu$ which is $(\alpha,{\mathcal{B}}_{[\rho,\rho^{\varepsilon}]},\tau)$-robust. Denote by $n_{\rho}\geq 0$ the integer part of $\frac{1}{2\ell}|\log\rho|$.
Assume $\varepsilon,\rho\lll_{\kappa}1$, then
[…]
Remark. Recall here that $\ell$ denotes the Lyapunov exponent of the $\operatorname{Ad}_{\star}\mu$-walk on $\mathfrak{g}$. Hence our choice for $n_{\rho}$ guarantees that the operator norm of $\operatorname{Ad}g^{-1}$ is roughly $\rho^{-1/2}$ when $g\overset{law}{\sim}{\mu^{*n_{\rho}}}$.
Proof.
[…]
Note that the compact set $X_{\rho^{\varepsilon}}:=\{\operatorname{inj}\geq\rho^{\varepsilon}\}$ can be covered by $\rho^{-O(\varepsilon)}$ balls of radius $\rho^{2\varepsilon}$, more precisely
$X_{\rho^{\varepsilon}}\subseteq\cup_{i\in I}B_{\rho^{2\varepsilon}}x_{i}$
where $\sharp I\leq\rho^{-O(\varepsilon)}$, $x_{i}\in X_{\rho^{\varepsilon}}$ for all $i$. As $\nu$ is supported on $X_{\rho^{\varepsilon}}$, we can then write
$\nu=\sum_{i\in I}\nu_{i}*\delta_{x_{i}}$
where $\nu_{i}$ is a Borel measure on $G$ with support in $B_{\rho^{2\varepsilon}}$. Note that the assumption that $\nu$ is $(\alpha,{\mathcal{B}}_{[\rho,\rho^{\varepsilon}]},0)$-robust implies that each $\nu_{i}$ is $(\alpha,{\mathcal{B}}_{[\rho,\rho^{\varepsilon}]},0)$-robust.
We now apply 4.3 to each $\nu_{i}$. We consider the family of charts $\varphi_{\theta}:\mathcal{U}\rightarrow\mathfrak{g}$ introduced in Section 4.3. In order to guarantee the distortion control requirement for $\varphi_{\theta}$, we introduce the renormalized truncation of $\mu^{*n}$ defined by
$\mu^{\prime}_{n}=\frac{\mu^{*n}_{|L_{\theta_{g}}\leq\rho^{-\varepsilon}}}{\mu^{*n}\{L_{\theta_{g}}\leq\rho^{-\varepsilon}\}}.$
By Lemma 4.5, this probability measure satisfies $\|\mu^{\prime}_{n}-\mu^{*n}\|\leq\rho^{\gamma}$ for some $\gamma=\gamma(\mu,\varepsilon)>0$. In particular, provided $\rho\lll_{\varepsilon}1$, the measure $\mu^{\prime}_{n}$ also satisfies the non-concentration estimates from Lemma 4.6. This allows us to apply 4.3 with $\Xi$ the law of $\theta_{g}$ when $g\overset{law}{\sim}{\mu^{\prime}_{n}}$ ($n=n_{\rho}$). We obtain some constant $\varepsilon_{1}>0$ depending only on $\kappa$, $\mu$ such that up to assuming $\varepsilon\lll_{\kappa}1$, $\rho\lll_{\kappa,\varepsilon}1$, there exists $\mathcal{G}_{i}\subseteq P$ with ${\mu^{\prime}_{n}}(\mathcal{G}_{i})\geq 1-\rho^{\varepsilon_{1}}$ satisfying for every $g\in\mathcal{G}_{i}$, that there exists a component measure $\nu_{i,g}\leq\nu_{i}$ with $\nu_{i,g}(G)\geq\nu_{i}(G)-\rho^{\varepsilon_{1}}$ and such that
(4.8) $\sup_{Q\in{\mathcal{R}}_{\rho}}\nu_{i,g}(\varphi_{\theta_{g}}^{-1}Q)\leq\rho^{\frac{3}{2}\alpha+\varepsilon_{1}}.$
On the other hand, the large deviation principle for the walk on $\mathbb{R}$ driven by $-\log\mathtt{r}_{g}\,\mathrm{d}\mu(g)$ guarantees that
the set $\mathcal{G}_{\mathtt{r}}=\{\,g\,:\,{\mathtt{r}_{g}^{-1}}\in[\rho^{-1/2+\varepsilon},\rho^{-1/2-\varepsilon}]\,\}\,\,$ satisfies $\,\,{\mu^{\prime}_{n}(\mathcal{G}_{\mathtt{r}})}\geq 1-\rho^{\varepsilon_{2}}$
for some $\varepsilon_{2}=\varepsilon_{2}(\mu,\varepsilon)>0$.
Setting $\mathcal{G}_{i,\mathtt{r}}=\mathcal{G}_{i}\cap\mathcal{G}_{\mathtt{r}}$ and using Lemma 4.4, observe that for $i\in I$, $g\in\mathcal{G}_{i,\mathtt{r}}$, for any ball $B_{\rho^{1/2}}y$ where $y\in X$, the intersection $(g^{-1}B_{\rho^{1/2}}y)\cap B_{\rho^{2\varepsilon}}x_{i}$ lifted to $B_{\rho^{2\varepsilon}}$ is included in at most $\rho^{-O(\varepsilon)}$ blocks of the form $\varphi_{\theta_{g}}^{-1}Q$ where $Q\in{\mathcal{R}}_{\rho}$. Hence, we get from (4.8),
(4.9) $\sup_{y\in X}\delta_{g}*\nu_{i,g}(B_{\rho^{1/2}}y)\leq\rho^{\frac{3}{2}\alpha+\varepsilon_{1}-O(\varepsilon)}.$
Setting $\mathcal{G}=\cap_{i\in I}\mathcal{G}_{i,\mathtt{r}}$ and recalling $\sharp I\leq\rho^{-O(\varepsilon)}$, we have $\mu^{\prime}_{n}(\mathcal{G})\geq 1-\rho^{\varepsilon_{1}-O(\varepsilon)}-\rho^{\varepsilon_{2}}$. We deduce
$\mu^{*n}(\{L_{\theta_{g}}\leq\rho^{-\varepsilon}\}\cap\mathcal{G})\geq 1-\rho^{\varepsilon_{1}-O(\varepsilon)}-\rho^{\varepsilon_{2}}-\rho^{\gamma}.$
Letting
$m_{n}^{\prime\prime}={\mu^{*n}}*\nu-\int_{G}\sum_{i}\delta_{g}*\nu_{i,g}{\mathbbm{1}}_{\{L_{\theta_{g}}\leq\rho^{-\varepsilon}\}\cap\mathcal{G}}\,\mathrm{d}\mu^{*n}(g),$
and taking $\varepsilon\lll_{\varepsilon_{1}}1$, we have $\|m_{n}^{\prime\prime}\|\leq\rho^{\varepsilon_{3}}$ where $\varepsilon_{3}=\varepsilon_{3}(\varepsilon,\varepsilon_{1},\varepsilon_{2},\gamma)>0$, while we see from (4.9) that $m_{n}^{\prime}:=({\mu^{*n}}*\nu)-m_{n}^{\prime\prime}$ satisfies
$\sup_{y\in X}m^{\prime}_{n}(B_{\rho^{1/2}}y)\leq\rho^{\frac{3}{2}\alpha+\varepsilon_{1}/2}.$
We now have checked the required dimensional increment. In order to conclude, we also need to check that $\mu^{*n}*\nu$ does not give too much mass to the cusp. This directly follows from Proposition 2.4 implying that for some constants $c,c^{\prime}>0$ depending on $\mu$, we have
$\mu^{*n}*\delta_{x}\{\operatorname{inj}<\rho^{1/2}\,\}\ll({\operatorname{inj}^{-c}(x)}e^{-c^{\prime}n}+1)\rho^{c/2},$
while by assumption, $\nu$ is supported on $\{\operatorname{inj}\geq\rho^{\varepsilon}\}$ with $\varepsilon<c^{\prime}/(2\ell c)$, and $n=n_{\rho}$. ∎
4.5. Proof of high dimension
We are finally able to show Proposition 4.2, namely that ${\mu^{*n}}*\delta_{x}$ reaches high dimension exponentially fast. The proof starts from positive dimension given by Proposition 3.1 and then proceeds by small increments using Proposition 4.8. Note however that Proposition 4.8 assumes non-concentration on a wide range of scales but the output dimensional increment only concerns a specific scale. Hence we need combine those single-scale increments to allow iterating the bootstrap. For this, we rely on the following lemma.
Lemma 4.9.
Let $\alpha,s,\rho\in(0,1]$, $\tau\in\mathbb{R}_{>0}$ be parameters. If $\nu$ is $(\alpha,\mathcal{B}_{r},\tau)$-robust for all $r\in[\rho,\rho^{s}]$, then for any $\varepsilon\in(0,\alpha)$, the measure $\nu$ is $(\alpha-\epsilon,\mathcal{B}_{[\rho,\rho^{s}]},\lceil\frac{\log s}{\log(1-\varepsilon)}\rceil\tau)$-robust.
Proof.
This is just a combination of two observations (1) if $\nu$ is $(\alpha,\mathcal{B}_{r},\tau)$-robust, then for every $t\in(0,1)$, it is $(t\alpha,\mathcal{B}_{[r^{1/t},r]},\tau)$-robust; (2) if $\nu$ is $(\alpha,\mathcal{B}_{I_{1}},\tau_{1})$-robust and $(\alpha,\mathcal{B}_{I_{2}},\tau_{2})$-robust, then $\nu$ is $(\alpha,\mathcal{B}_{I_{1}\cup I_{2}},\tau_{1}+\tau_{2})$-robust. See [5, Lemma 4.5] for details. ∎
Proof of Proposition 4.2.
Let $A>0$ be a large enough constant depending on the initial data $\mu$. Combining Proposition 3.1 and Proposition 2.4, we may assume $\kappa>0$ small enough from the start, so that for any $M>0$, for every $\rho\lll_{M}1$ and $n\geq M|\log\rho|+A|\log\operatorname{inj}(x)|$, the measure
[…]
By Proposition 4.8, there is some small constant $\varepsilon=\varepsilon(\mu,\kappa)>0$, such that up to imposing from the start $M\ggg_{\kappa}1$, we have for all $\rho\lll_{\kappa,M}1$, all $n\geq(\frac{1}{2\ell}+1)M|\log\rho|+A|\log\operatorname{inj}(x)|$ and $r\in[\rho^{2M\varepsilon},\rho^{1/(2M\varepsilon)}]$,
[…]
These estimates for single scales can be combined using Lemma 4.9 to get under the same conditions:
[…]
The argument in the last paragraph can be applied iteratively, adding at each step $k$ the value $+\varepsilon$ to the dimension provided the latter is not yet above $1-\kappa$ and provided $M\ggg_{\kappa,k}1$. As the value of $\varepsilon$ only depends on $\kappa$, we reach dimension $1-\kappa$ in a finite number of steps. This concludes the proof.
5. From high dimension to equidistribution
We consider the one-parameter family of probability measures $(\eta_{t})_{t>0}$ on $G$ given by
$\eta_{t}=a(t)u(s)\,\mathrm{d}\sigma(s).$
We show Proposition 5.1, stating that as $t\to+\infty$, a probability measure on $X$ with dimension close to $3$ equidistributes under the $\eta_{t}$-process toward the Haar measure on $X$, and does so with exponential rate. From this we deduce B’ (whence B) and C’ (whence C).
Proposition 5.1.
There exist $\kappa,\rho_{0}>0$ such that the following holds for all $\rho\in(0,\rho_{0}]$, $\tau>0$.
Let $\nu$ be a Borel measure on $X$ that is $(1-\kappa,{\mathcal{B}}_{\rho},\tau)$-robust and has mass at most $1$. Then for any $t\in[\rho^{-1/4},\rho^{-1/2}]$, for any $f\in B^{\infty}_{\infty,1}(X)$ with $m_{X}(f)=0$, we have
$|\eta_{t}*\nu(f)|\leq(\rho^{\kappa}+\tau){\mathcal{S}}_{\infty,1}(f).$
The argument relies on the quantitative decay of correlations for $X$. Consider the unitary representation of $G$ on $L^{2}(X)$ defined by the formula $g.f=f\circ g^{-1}$. From the combination of [3, Lemma 3] and [21, Equations (6.1), (6.9)], we know there exists $\delta_{0}=\delta_{0}(\Lambda)>0$ such that for any function $f\in B^{\infty}_{2,1}(X)$, any $g\in G$, we have
(5.1) $\lvert\langle g.f,f\rangle\rvert\ll\lVert g\rVert^{-\delta_{0}}{\mathcal{S}}_{2,1}(f)^{2}.$
From this we deduce a spectral gap for the family of Markov operators $P_{\eta_{t}}$. Recall that $P_{\eta_{t}}$ is the operator acting on non-negative measurable functions on $X$ given by the formula
$P_{\eta_{t}}f(x)=\int_{G}f(gx)\,\mathrm{d}\eta_{t}(g).$
$P_{\eta_{t}}$ extends continuously into an operator on $L^{2}(X)$ of norm $1$.
Proposition 5.2 (Spectral gap for $P_{\eta_{t}}$).
There exists $c>0$ such that for any function $f\in B^{\infty}_{2,1}(X)$ with $m_{X}(f)=0$, we have
$\forall t>1,\quad\|P_{\eta_{t}}f\|_{L^{2}}\ll t^{-c}{\mathcal{S}}_{2,1}(f).$
Proof.
Using (5.1), we have
$\|P_{\eta_{t}}f\|^{2}_{L^{2}}$ $=\iint_{G^{2}}\langle g^{-1}.f,h^{-1}.f\rangle\,\mathrm{d}\eta_{t}(g)\,\mathrm{d}\eta_{t}(h)$
$=\iint_{G^{2}}\langle hg^{-1}.f,f\rangle\,\mathrm{d}\eta_{t}(g)\,\mathrm{d}\eta_{t}(h)$
$\ll{\mathcal{S}}_{2,1}(f)^{2}\iint_{G^{2}}\lVert hg^{-1}\rVert^{-\delta_{0}}\,\mathrm{d}\eta_{t}(g)\,\mathrm{d}\eta_{t}(h)$
Plugging in the definition of $\eta_{t}$
$hg^{-1}\,\mathrm{d}\eta_{t}(g)\,\mathrm{d}\eta_{t}(h)=u(t(s_{1}-s_{2}))\,\mathrm{d}\sigma(s_{1})\,\mathrm{d}\sigma(s_{2}),$
we get
$\lVert P_{\eta_{t}}f\rVert^{2}_{L^{2}}\ll\iint_{\mathbb{R}^{2}}\max\{1,t|s_{1}-s_{2}|\}^{-\delta_{0}}\,\mathrm{d}\sigma(s_{1})\,\mathrm{d}\sigma(s_{2}){\mathcal{S}}_{2,1}(f)^{2}.$
Finally, the Hölder regularity of $\sigma$ from Lemma 2.1(ii) implies that
$\sigma^{\otimes 2}\underbrace{\{(s_{1},s_{2}):t|s_{1}-s_{2}|\leq t^{1/2}\}}_{E}\ll t^{-c},$
for some constant $c=c(\sigma)>0$. Hence,
$\iint_{\mathbb{R}^{2}}\max\{1,t\lvert s_{1}-s_{2}\rvert\}^{-\delta_{0}}\,\mathrm{d}\sigma(s_{1})\,\mathrm{d}\sigma(s_{2})$
$=$ $\iint_{E}1\,\mathrm{d}\sigma\otimes\sigma+\iint_{\mathbb{R}^{2}\smallsetminus E}(t\lvert s_{1}-s_{2}\rvert)^{-\delta_{0}}\,\mathrm{d}\sigma(s_{1})\,\mathrm{d}\sigma(s_{2})$
$\ll$ $t^{-c}+t^{-\delta_{0}/2},$
concluding the proof. ∎
To prove Proposition 5.1 we mollify the measure $\nu$ at some scale $\rho>0$: let $\nu_{\rho}$ be the Borel measure on $X$ defined by
$\nu_{\rho}(f)=\frac{1}{m_{G}(B_{\rho})}\int_{X}\int_{B_{\rho}}f(gx)\,\mathrm{d}m_{G}(g)\,\mathrm{d}\nu(x),$
where $f$ denotes here any non-negative measurable function on $X$.
Note that if $\nu$ is supported on the compact set $\{\operatorname{inj}\geq\rho\}$, then by a change of variable $g\in B_{\rho}\mapsto gx\in B_{\rho}x$ and the Fubini-Lebesgue theorem, we have
$\nu_{\rho}(f)$ $=\frac{1}{m_{G}(B_{\rho})}\iint_{X\times X}{\mathbbm{1}}_{y\in B_{\rho}x}f(y)\,\mathrm{d}m_{X}(y)\,\mathrm{d}\nu(x)$
$=\frac{1}{m_{G}(B_{\rho})}\int_{X}f(y)\int_{X}{\mathbbm{1}}_{x\in B_{\rho}y}\,\mathrm{d}\nu(x)\,\mathrm{d}m_{X}(y).$
This implies that $\nu$ is absolutely continuous with respect to $m_{X}$ and its Radon-Nikodym derivative is
(5.2) $\frac{\mathrm{d}\nu_{\rho}}{\mathrm{d}m_{X}}(x)=\frac{\nu(B_{\rho}x)}{m_{G}(B_{\rho})}.$
In particular, if $\nu$ is $(1-\kappa,{\mathcal{B}}_{\rho},0)$-robust, then
$\Bigl{\lVert}\frac{\mathrm{d}\nu_{\rho}}{\mathrm{d}m_{X}}\Bigr{\rVert}_{L^{\infty}}\ll\rho^{-3\kappa}.$
Proof of Proposition 5.1.
We let $\kappa,r,\rho_{0}>0$ be parameters to specify below, $\rho,\tau,\nu$ as in the proposition, and consider a test function $f\in B^{\infty}_{\infty,1}(X)$ with zero average. Clearly we may assume $\tau=0$, i.e. $\nu$ is $(1-\kappa,{\mathcal{B}}_{\rho},0)$-robust.
We can write for any $n\geq 0$,
$|\eta_{t}*\nu(f)|=\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu\right\rvert\leq\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu_{\rho}\right\rvert+\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu_{\rho}-\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu\right\rvert.$
The first term is bounded by
$\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu_{\rho}\right\rvert$ $\leq\lVert P_{\eta_{t}}f\rVert_{L^{1}}\Bigl{\lVert}\frac{\mathrm{d}\nu_{\rho}}{\mathrm{d}m_{X}}\Bigr{\rVert}_{L^{\infty}}$
$\leq\lVert P_{\eta_{t}}f\rVert_{L^{2}}\Bigl{\lVert}\frac{\mathrm{d}\nu_{\rho}}{\mathrm{d}m_{X}}\Bigr{\rVert}_{L^{\infty}}$
$\ll t^{-c}{\mathcal{S}}_{2,1}(f)\rho^{-3\kappa},$
where the last inequality uses the spectral gap estimate from Proposition 5.2 on the one hand, and the assumption that $\nu$ is $(1-\kappa,{\mathcal{B}}_{\rho},0)$-robust on the other hand.
From the definition of $\nu_{\rho}$, the second term is bounded by
$\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu_{\rho}-\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu\right\rvert\leq\rho{\mathcal{S}}_{\infty,1}(P_{\eta_{t}}f)\ll\rho t{\mathcal{S}}_{\infty,1}(f).$
Put together, we have obtained
$|\eta_{t}*\nu(f)|\ll\bigl{(}t^{-c}\rho^{-3\kappa}+\rho t){\mathcal{S}}_{\infty,1}(f)\ll\rho^{\kappa}{\mathcal{S}}_{\infty,1}(f)$
where the last upper bound holds for $t\in[\rho^{-1/4},\rho^{-1/2}]$, up to choosing $\kappa=c/16$ and $\rho_{0}$ is small enough in terms of $\kappa$. ∎
We now address the
Proof of B’.
Note first that until now, we considered a measure $\lambda$ supported on $\operatorname{Aff}(\mathbb{R})^{+}$ while B’ allows for a measure $\lambda$ on $\operatorname{Aff}(\mathbb{R})$. We reduce easily to the $\operatorname{Aff}(\mathbb{R})^{+}$-case via the following lemma.
Lemma 5.3.
We may assume the measure $\lambda$ is supported on $\operatorname{Aff}(\mathbb{R})^{+}$.
Proof.
Set $\Omega=\operatorname{Aff}(\mathbb{R})^{\mathbb{N}}$, consider the stopping time $\tau_{+}:\Omega\rightarrow\mathbb{N}$ defined for $\underline{\phi}=(\phi_{i})_{i\geq 1}\in\Omega$ by
$\tau_{+}(\underline{\phi})=\inf\{n\geq 1\,:\,\mathtt{r}_{\phi_{1}\circ\dots\circ\phi_{n}}>0\}.$
Write $\lambda^{*\tau_{+}}=\int_{\Omega}\delta_{\phi_{1}\circ\dots\circ\phi_{\tau_{+}(\underline{\phi})}}\,\mathrm{d}\lambda^{\otimes\mathbb{N}}(\underline{\phi})$. Then by the strong Markov property, (see [4, Lemme A.2]), the measure $\sigma$ is $\lambda^{*\tau_{+}}$-stationary. Moreover, $\lambda^{*\tau_{+}}$ has finite exponential moment (because $\tau_{+}$ does). Its support $\operatorname{supp}\lambda^{*\tau_{+}}$ does not have common fixed point on $\mathbb{R}$, for otherwise the group generated by $\operatorname{supp}\mu$ would have an orbit of cardinality $2$ and hence fixes the barycenter of this orbit. ∎
Now that $\lambda$ is supported on $\operatorname{Aff}(\mathbb{R})^{+}$, we denote by $\mu$ the corresponding measure on $P$. We relate the $\eta_{t}$-process with the $\mu$-random walk thanks to the following lemma.
Lemma 5.4 ($\eta_{t}$-process vs $\mu$-walk).
Given $t>0$, $n\geq 0$, we have
$\eta_{t}=\int_{P}\eta_{t\mathtt{r}_{g}}*\delta_{g}\,\mathrm{d}\mu^{*n}(g).$
Proof.
Observe that for any $s\in\mathbb{R}$ and $g\in P$,
$a(t\mathtt{r}_{g})u(s)g=a(t\mathtt{r}_{g})u(s)a(\mathtt{r}_{g})^{-1}u(\mathtt{b}_{g})=a(t)u(\mathtt{r}_{g}s+\mathtt{b}_{g}).$
The claim then follows from the $\lambda^{*n}$-stationarity and the fact that $\mu^{*n}$ and $\lambda^{*n}$ are related by the anti-isomorphism between $P$ and $\operatorname{Aff}(\mathbb{R})^{+}$. ∎
We now discretize the set of values of $\mathtt{r}_{g}$ that appears in the part $\eta_{t\mathtt{r}_{g}}$. Given $r_{0},r_{1}>0$ observe that
$\eta_{tr_{0}}=\delta_{a(r_{0}r_{1}^{-1})}*\eta_{tr_{1}}$
Hence, for any finite Borel measure $\nu$ on $X$, we get
(5.3) $|\eta_{tr_{0}}*\nu(f)-\eta_{tr_{1}}*\nu(f)|\ll|\log(r_{0}r_{1}^{-1})|\,\nu(X){\mathcal{S}}_{\infty,1}(f).$
Let $\rho>0$, consider a parameter $\alpha\in(0,1)$ to be specified later depending on $\mu$, and set :=${\mathscr{R}}:=\{(1+\rho^{\alpha})^{k}\,:\,k\in\mathbb{Z}\}$. Combining (5.3) with Lemma 5.4, we get for any $x\in X$, $f\in B^{\infty}_{\infty,1}(X)$, that
(5.4) $\lvert\eta_{t}*\delta_{x}(f)\rvert\leq\sum_{r\in{\mathscr{R}}}\lvert\eta_{tr}*\mu^{n}_{r}*\delta_{x}(f)\rvert+O(\rho^{\alpha}{\mathcal{S}}_{\infty,1}(f))$
where $\mu^{n}_{r}$ denotes the restriction of $\mu^{*n}$ to the set $\{\,g\in P:\mathtt{r}_{g}\in[r,r(1+\rho^{\alpha})[\,\}$.
Let $\kappa=\kappa(\mu)>0$ as in Proposition 5.1. Assume $\operatorname{inj}(x)\geq\rho$. By Proposition 4.2, there are constants $C=C(\mu)>1$ and $\varepsilon_{1}=\varepsilon_{1}(\mu)>0$ such that, provided $\rho\lll 1$, the measure $\mu^{*n}*\delta_{x}$ on $X$ is $(1-\kappa,{\mathcal{B}}_{\rho},\rho^{\varepsilon_{1}})$-robust for any $n\geq C\lvert\log\rho\rvert$. For the rest of this proof, we specify $n,t$ in terms of $\rho$ as
(5.5) $n=\lceil C\lvert\log\rho\rvert\rceil,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,t=\rho^{-C\ell-3/8}.$
Consider
=${\mathscr{R}}^{\prime}=\{\,r\in{\mathscr{R}}:\rho^{-1/4}\leq tr\leq\rho^{-1/2}\,\}={\mathscr{R}}\cap[\rho^{C\ell+1/8},\rho^{C\ell-1/8}].$
On the one hand, for each $r\in{\mathscr{R}}^{\prime}$, note that $\mu^{n}_{r}*\delta_{x}\leq\mu^{*n}*\delta_{x}$ is still $(1-\kappa,{\mathcal{B}}_{\rho},\rho^{\varepsilon_{1}})$-robust. Therefore the choice of ${\mathscr{R}}^{\prime}$ allows us to use Proposition 5.1 to obtain
(5.6) $\lvert\eta_{tr}*\mu^{n}_{r}*\delta_{x}(f)\rvert\leq(\rho^{\kappa}+\rho^{\varepsilon_{1}}){\mathcal{S}}_{\infty,1}(f).$
On the other hand, by the large deviation estimates for sums of i.i.d real random variables, there is a constant $\varepsilon_{2}=\varepsilon_{2}(\mu,C)>0$ such that
$\mu^{*n}\{\,g:\lvert n\ell+\log\mathtt{r}_{g}\rvert>\lvert\log\rho\rvert/10\,\}<\rho^{\varepsilon_{2}}.$
whenever $\rho\lll_{C}1$. This implies an upper bound on the total mass $\sum_{r\in{\mathscr{R}}\smallsetminus{\mathscr{R}}^{\prime}}\mu^{n}_{r}(P)\leq\rho^{\varepsilon_{2}}$ and hence
(5.7) $\sum_{r\in{\mathscr{R}}\smallsetminus{\mathscr{R}}^{\prime}}\lvert\eta_{tr}*\mu^{n}_{r}*\delta_{x}(f)\rvert\leq\rho^{\varepsilon_{2}}{\mathcal{S}}_{\infty,1}(f).$
Putting (5.4), (5.6), (5.7) together, we have
$\lvert\eta_{t}*\delta_{x}(f)\rvert$ (#$\leq\bigl{(}\#{\mathscr{R}}^{\prime}(\rho^{\kappa}+\rho^{\varepsilon_{1}})+\rho^{\varepsilon_{2}}\bigr{)}{\mathcal{S}}_{\infty,1}(f)$
$\leq\bigl{(}\rho^{\kappa/2}+\rho^{\varepsilon_{1}/2}+\rho^{\varepsilon_{2}}\bigr{)}{\mathcal{S}}_{\infty,1}(f)$
where the second bound uses $\sharp{\mathscr{R}}^{\prime}\ll\frac{2C\ell|\log\rho|}{\log(1+\rho^{\alpha})}\sim|\log\rho|\rho^{-\alpha}$ and assumes $\alpha\leq\kappa\varepsilon_{1}/4$, $\rho\lll 1$.
Viewing $\rho$ as varying with $t$ according to (5.5), we can summarize the above as the following. For $t>1$ sufficiently large, for any $x\in X$ with $\operatorname{inj}(x)\geq t^{-(C\ell+3/8)^{-1}}$,
$\lvert\eta_{t}*\delta_{x}(f)\rvert\leq t^{-\varepsilon_{3}}{\mathcal{S}}_{\infty,1}(f)$
with $\varepsilon_{3}=\frac{\min\{\kappa,\varepsilon_{1},\varepsilon_{2}\}}{8C\ell+3}$. Finally, if $x$ is a point with $\operatorname{inj}(x)<t^{-(C\ell+3/8)^{-1}}$ then $\operatorname{inj}(x)^{-1}t^{-\varepsilon_{3}}\geq 1$. ∎
Effective equidistribution for the $\mu$-walk on $X$ can also be handled similarly (and more simply).
Proof of C’.
Proposition 5.2 and Proposition 5.1 are still valid with $(t,\eta_{t})$ replaced by $(e^{n},\mu^{n})$ where $n$ is an integer parameter (essentially same proof, using Lemma 2.3 instead of Lemma 2.1). Combining Proposition 5.1 with Proposition 4.2, we get the theorem. More precisely, given $\rho\lll 1$, Proposition 4.2 tells us that for $m\ggg|\log\rho|+|\log\operatorname{inj}(x)|$, the measure $\nu:=\mu^{*m}*\delta_{x}$ satisfies the conditions required to apply Proposition 5.1. Then choosing $n>m$ such that $n-m\in[\frac{1}{4}|\log\rho|,\frac{1}{2}|\log\rho|]$, we obtain that $\mu^{*n}*\delta_{x}$ is $\rho^{\varepsilon}$-equidistributed for some small constant $\varepsilon=\varepsilon(\mu)>0$. This finishes the proof. ∎
6. Double equidistribution
In this section, we show effective double equidistribution properties for expanding fractals. This result refines B’ and will play a role in the proof of the divergent case of A’. We use the notations set in Section 2. In particular, $X=\operatorname{SL}_{2}(\mathbb{R})/\Lambda$ where $\Lambda$ is an arbitrary lattice, $x_{0}=\Lambda/\Lambda$ is the basepoint of $X$, and $\sigma$ is a probability measure on $\mathbb{R}$ that is stationary for a randomized orientation preserving IFS $\lambda$ with a finite exponential moment.
Given a probability measure $\xi$ on $\mathbb{R}$, bounded continuous functions $f_{1},f_{2}:X\rightarrow\mathbb{R}$, and times $t_{2}\geq t_{1}>0$, we introduce
(6.1) $\Delta^{\xi}_{f_{1},f_{2}}(t_{1},t_{2}):=\left\lvert\int_{\mathbb{R}}f_{1}\bigl{(}a(t_{1})u(s)x_{0}\bigr{)}f_{2}\bigl{(}a(t_{2})u(s)x_{0}\bigr{)}\,\mathrm{d}\xi(s)-m_{X}(f_{1})m_{X}(f_{2})\right\rvert.$
Hence the probability measure $u(s)x_{0}\,\mathrm{d}\xi(s)$ on $X$ enjoys double equidistribution toward $m_{X}$ under expansion by the diagonal flow if for any such $f_{1},f_{2}$, we have
$\Delta^{\xi}_{f_{1},f_{2}}(t_{1},t_{2})\to 0\,$ as $\,\inf(t_{2}t_{1}^{-1},t_{1})\to+\infty$.
In this section, we show that in the case where $\xi=\sigma$, double equidistribution holds with an effective rate.
Proposition 6.1 (Effective double equidistribution of expanding fractals).
For every $\eta>0$, there exist $C,c>0$ such that for all $t_{1},t_{2}>1$ with $t_{2}\geq t^{1+\eta}_{1}$ and $f_{1},f_{2}\in B^{\infty}_{\infty,1}(X)$, we have
(6.2) $\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{\infty,1}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,1}(f_{1}){\mathcal{S}}_{\infty,1}(f_{2})t_{2}^{-c}.$
Taking $f_{2}=1$ and letting $t_{2}\to+\infty$, we see that Proposition 6.1 implies B’. The proposition assumes that the times $t_{1}$, $t_{2}$ are slightly separated, via the condition $t_{2}\geq t^{1+\eta}_{1}>1$. In fact we will see later in Corollary 7.6 that (6.2) also implies an upper bound in the short-range regime $t^{1+\eta}_{1}\geq t_{2}\geq t_{1}$. Namely, for all $t_{2}\geq t_{1}>1$, we have
(6.3) $\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{2,1}(f_{1}){\mathcal{S}}_{2,1}(f_{2})t_{1}^{c}t_{2}^{-c}\\ +C{\mathcal{S}}_{\infty,1}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,1}(f_{1}){\mathcal{S}}_{\infty,1}(f_{2})t_{2}^{-c},$
for possibly different constants $C,c>0$, depending only on $\Lambda$, $\sigma$.
The proof of Proposition 6.1 is inspired by [30, Theorem 1.2], which deals with absolutely continuous measures, and [24, Proposition 10.1] which deals with fractal measures and either short-range or long-range regime (i.e. $t_{2}\in[t_{1},t_{1}^{1+\varepsilon}]$ or $t_{2}\geq t^{C}_{1}$ where $C^{-1},\varepsilon\lll 1$). Here is the main idea behind the proof. By self-similarity of $\sigma$, the distribution $(a(t_{1})u(s)x_{0},a(t_{2})u(s)x_{0})\,\mathrm{d}\sigma(s)$ is roughly that of $(hx_{0},ghx_{0})\,\mathrm{d}\mu^{*n_{1}}(h)\,\mathrm{d}\mu^{*n_{2}}(g)$ where $n_{1}\simeq\ell^{-1}\log t_{1}$ and $n_{2}\simeq\ell^{-1}\log(t_{2}/t_{1})$, with $\ell$ the Lyapunov exponent of $\operatorname{Ad}_{\star}\mu$, see (2.2). Then we apply B’ to the $\mu$-random walk starting at $hx_{0}$, to get that the variable in the second coordinate equidistributes conditionally to the first one. B’ tells us the first coordinate equidistributes as well, whence the result.
Proof.
To lighten notations, we write $\mathcal{S}=\mathcal{S}_{\infty,1}$ and ${\mathcal{S}}(f_{1},f_{2})={\mathcal{S}}(f_{1}){\mathcal{S}}(f_{2})$. Noting the relation
$\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2})\leq\Delta^{\sigma}_{f_{1},f_{2}-m_{X}(f_{2})}(t_{1},t_{2})+\left\lvert\int_{\mathbb{R}}f_{1}\bigl{(}a(t_{1})u(s)x_{0}\bigr{)}\,\mathrm{d}\sigma(s)-m_{X}(f_{1})\right\rvert\lvert m_{X}(f_{2})\rvert$
and that B’ provides us with a constant $c=c(\Lambda,\sigma)>0$ such that
$\left\lvert\int_{\mathbb{R}}f_{1}\bigl{(}a(t_{1})u(s)x_{0}\bigr{)}\,\mathrm{d}\sigma(s)-m_{X}(f_{1})\right\rvert\ll{\mathcal{S}}(f_{1})t_{1}^{-c},$
we can reduce to the case where $m_{X}(f_{2})=0$.
Thus, we are left to bound the integral
$I:=\int_{\mathbb{R}}f_{1}\bigl{(}a(t_{1})u(s)x_{0}\bigr{)}f_{2}\bigl{(}a(t_{2})u(s)x_{0}\bigr{)}\,\mathrm{d}\sigma(s)$
by a quantity of the form $O_{\eta}({\mathcal{S}}(f_{1},f_{2})t_{2}^{-\kappa})$ where $\kappa=\kappa(\Lambda,\sigma,\eta)>0$.
We first use the $\lambda$-stationarity of $\sigma$ to allow the argument in $f_{2}$ to vary randomly conditionally to that in $f_{1}$. Let $M,n\geq 1$ be (large) parameters to be specified later. By the large deviation principle for sums of i.i.d. real random variables, there exists $\varepsilon=\varepsilon(\lambda,M)>0$ such that, provided $n\ggg_{M}1$,
(6.4) $\lambda^{*n}({\mathscr{C}})\geq 1-e^{-n\varepsilon},\text{ where }{\mathscr{C}}:=\bigl{\{}\,\phi\in\operatorname{Aff}(\mathbb{R})^{+}:\lvert n\ell+\log\mathtt{r}_{\phi}\rvert\leq\frac{n\ell}{M}\,\bigr{\}}.$
Note that for any $\phi\in{\mathscr{C}}$, $t>1$, $s\in\mathbb{R}$, we have $|\phi(s)-\mathtt{b}_{\phi}|\leq\mathtt{r}_{\phi}|s|\leq e^{-(1-\frac{1}{M})n\ell}|s|$, whence
(6.5) $|f_{1}\bigl{(}a(t)u(\phi(s))x_{0}\bigr{)}-f_{1}\bigl{(}a(t)u(\mathtt{b}_{\phi})x_{0}\bigr{)}|\ll te^{-(1-\frac{1}{M})n\ell}|s|\mathcal{S}(f_{1}).$
Moreover, as $s$ varies with law $\sigma$, its size is controlled by the moment estimate of Lemma 2.1. Namely, there is some $\gamma=\gamma(\sigma)>0$ such that for all $R>1$,
(6.6) $\sigma\{\,s\in\mathbb{R}:|s|>R\,\}\ll R^{-\gamma}.$
Splitting the integral on $s$ according to whether $\lvert s\rvert\leq e^{\frac{n\ell}{M}}$ or not and using (6.5) and (6.6), we obtain
(6.7) $\int_{\mathbb{R}}\bigl{\lvert}f_{1}\bigl{(}a(t)u(\phi(s))x_{0}\bigr{)}-f_{1}\bigl{(}a(t)u(\mathtt{b}_{\phi})x_{0}\bigr{)}\bigr{\rvert}\,\mathrm{d}\sigma(s)\ll(e^{-\frac{n\ell\gamma}{M}}+te^{-(1-\frac{2}{M})n\ell})\mathcal{S}(f_{1}).$
Using the $\lambda$-stationarity of $\sigma$, then applying (6.4) followed by (6.7), we deduce
$I$ $=\int_{\operatorname{Aff}(\mathbb{R})^{+}}\int_{\mathbb{R}}f_{1}\bigl{(}a(t_{1})u(s)x_{0}\bigr{)}f_{2}\bigl{(}a(t_{2})u(s)x_{0}\bigr{)}\,\mathrm{d}(\phi_{\star}\sigma)(s)\,\mathrm{d}\lambda^{*n}(\phi)$
$=\int_{{\mathscr{C}}}f_{1}\bigl{(}a(t_{1})u(\mathtt{b}_{\phi})x_{0}\bigr{)}\int_{\mathbb{R}}f_{2}\bigl{(}a(t_{2})u(s)x_{0}\bigr{)}\,\mathrm{d}(\phi_{\star}\sigma)(s)\,\mathrm{d}\lambda^{*n}(\phi)+E_{1}$
where $E_{1}$ stands for the error term $E_{1}=O_{M}(e^{-\frac{n\ell\gamma}{M}}+t_{1}e^{-(1-\frac{2}{M})n\ell}+e^{-n\varepsilon}){\mathcal{S}}(f_{1},f_{2})$.
It follows that
(6.8) $\lvert I\rvert\leq{\mathcal{S}}(f_{1})\int_{{\mathscr{C}}}\left\lvert\int_{\mathbb{R}}f_{2}\bigl{(}a(t_{2})u(s)x_{0}\bigr{)}\,\mathrm{d}(\phi_{\star}\sigma)(s)\right\rvert\,\mathrm{d}\lambda^{*n}(\phi)+|E_{1}|$
The inner integral invovling $f_{2}$ can be bounded using B’. Indeed, recall that $m_{X}(f_{2})=0$ and that for any $t>0$ and any $s\in\mathbb{R}$,
$a(t)u(\phi(s))=a(t\mathtt{r}_{\phi})u(s)h_{\phi}$
where $h_{\phi}=a(\mathtt{r}_{\phi}^{-1})u(\mathtt{b}_{\phi})$. Hence, for any $\phi\in{\mathscr{C}}$,
$\int_{\mathbb{R}}f_{2}\bigl{(}a(t_{2})u(s)x_{0}\bigr{)}\,\mathrm{d}(\phi_{\star}\sigma)(s)$ $=\int f_{2}\bigl{(}a(t_{2}\mathtt{r}_{\phi})u(s)h_{\phi}x_{0}\bigr{)}\,\mathrm{d}\sigma(s)$
(6.9) $=O\bigl{(}\operatorname{inj}(h_{\phi}x_{0})^{-1}\mathcal{S}(f_{2})t_{2}^{-c}e^{c(1+1/M)n\ell}\bigr{)}.$
where $c=c(\Lambda,\sigma)>0$ is the exponent provided by B’. Note that $h_{\phi}$ has law $\mu^{*n}$ when $\phi$ varies randomly according to $\lambda^{*n}$. Hence, by the effective recurrence of the $\mu$-random walk on $X$ (Proposition 2.4), there exists $\delta=\delta(\Lambda,\lambda)>0$ such that
(6.10) $\lambda^{*n}\bigl{\{}\,\phi\in\operatorname{Aff}(\mathbb{R})^{+}:\operatorname{inj}(h_{\phi}x_{0})\leq e^{-\frac{c}{M}n\ell)}\,\bigr{\}}\ll e^{-\delta\frac{c}{M}n\ell}.$
Note that for $\phi\in{\mathscr{C}}$ not belonging to the set in (6.10), the error term in (6) is bounded by $O(\mathcal{S}(f_{2})t_{2}^{-c}e^{c(1+2/M)n\ell})$. Therefore, we see from (6.8), (6), (6.10) that
$\lvert I\rvert\ll(t_{2}^{-c}e^{c(1+2/M)n\ell}+e^{-\delta\frac{c}{M}n\ell}){\mathcal{S}}(f_{1},f_{2})+|E_{1}|.$
Recalling the value of $E_{1}$ and choosing $n$ such that $n\ell=\frac{1}{2}\log t_{1}+\frac{1}{2}\log t_{2}+O(1)$, we obtain
$\lvert I\rvert\ll_{M}{\mathcal{S}}(f_{1},f_{2})\Bigl{(}(t_{2}/t_{1})^{-c/2}(t_{1}t_{2})^{c/M}+(t_{1}t_{2})^{-c^{\prime}}+(t_{2}/t_{1})^{-1/2}(t_{1}t_{2})^{1/M}\Bigr{)}$
where $c^{\prime}>0$ only depends on $\Lambda$, $\lambda$, $\sigma$, $M$. The desired estimate $\lvert I\rvert\ll t_{2}^{-\kappa}$ follows, provided $M$ has been chosen large enough from the start depending on the separation parameter $\eta$. ∎
7. The dichotomy
We show that an arbitrary probability measure $\xi$ on $\mathbb{R}$ obeys the Khintchine dichotomy provided that the pushfoward $a(t)u(s)\operatorname{SL}_{2}(\mathbb{Z})\,\mathrm{d}\xi(s)$ exhibits certain effective equidistribution properties on $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$ for large $t$. We deduce A’ (whence A). We use the notations introduced in Section 2.
Definition 7.1.
Let $\xi$ be a probability measure on $\mathbb{R}$. We say that $\xi$ satisfies the effective single equidistribution property on $X$ if there are constants $C,c>0$ such that
(7.1) $\forall f\in B^{\infty}_{\infty,1}(X),\,\forall t>1,\\ \left\lvert\int_{\mathbb{R}}f\bigl{(}a(t)u(s)x_{0}\bigr{)}\,\mathrm{d}\xi(s)-m_{X}(f)\right\rvert\leq C{\mathcal{S}}_{\infty,1}(f)t^{-c}.$
We say that $\xi$ satisfies the effective double equidistribution property on $X$ if for any $\eta>0$, there are constants $C,c>0$ such that
(7.2) $\forall f_{1},f_{2}\in B^{\infty}_{\infty,1}(X),\,\forall t_{1}>1,\,\forall t_{2}>t_{1}^{1+\eta},\\ \Delta^{\xi}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{\infty,1}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,1}(f_{1}){\mathcal{S}}_{\infty,1}(f_{2})t_{2}^{-c}.$
where the notation $\Delta^{\xi}_{f_{1},f_{2}}(t_{1},t_{2})$ is defined in (6.1). See Corollary 7.6 for an alternative characterization.
In [24], Khalil-Luethi showed that effective single equidistribution implies the convergent case of the Khintchine dichotomy.
Theorem 7.2 (Convergent case [24, Theorem 9.1]).
Let $\xi$ be a probability measure on $\mathbb{R}$ satisfying the effective single equidistribution property (7.1) on $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$. Then for every non-increasing function $\psi:\mathbb{N}\to\mathbb{R}_{>0}$ such that $\sum_{q}\psi(q)<\infty$, we have
$\xi(W(\psi))=0.$
We show that effective double equidistribution implies the divergent case of the Khintchine dichotomy. Moreover our method yields quantitative estimates on the number of solutions of the Diophantine inequality when bounding the denominator. We set ${\mathcal{P}}(\mathbb{Z}^{2}):=\{\,(p,q)\in\mathbb{Z}^{2}:\gcd(p,q)=1\,\}$ the set of primitive elements in $\mathbb{Z}^{2}$. We let $\zeta(t)=\sum_{n\geq 1}n^{-t}$ denote the Riemann zeta function.
Theorem 7.3 (Divergent case).
Let $\xi$ be a probability measure on $\mathbb{R}$ satisfying the effective double equidistribution property (7.2) on $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$. Let $\psi:\mathbb{N}\to\mathbb{R}_{>0}$ be a non-increasing function satisfying $\sum_{q}\psi(q)=\infty$, as well as
(7.3) $\forall q\in\mathbb{N},\quad\psi(q)\leq q^{-1}.$
Then for $\xi$-almost every $s\in\mathbb{R}$, as $N\to+\infty$, we have
[…]
The same holds if we ask for $-\psi(q)<qs-p\leq 0$ instead.
Without the extra domination assumption (7.3) on the approximation function $\psi$, we still have a quantitative lower bound (which tends to infinity).
Corollary 7.4.
If $\xi$ satisfies (7.2) on $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$ and $\psi:\mathbb{N}\to\mathbb{R}_{>0}$ is non-increasing with $\sum_{q}\psi(q)=\infty$, then for $\xi$-almost every $s\in\mathbb{R}$, as $N\to+\infty$, we have
[…]
The same holds if we ask for $-\psi(q)<qs-p\leq 0$ instead.
Assuming 7.3, we establish Corollary 7.4 and A’.
Proof of Corollary 7.4.
It follows from 7.3 applied to the approximation function $q\mapsto\min(\psi(q),q^{-1})$. Indeed, this application is allowed because we have $\sum_{q}\min(\psi(q),q^{-1})=\infty$. To justify this, observe that given any non-increasing function $\Psi:\mathbb{N}\rightarrow\mathbb{R}^{+}$, we have $\sum_{q}\Psi(q)=\infty$ if and only if $\sum_{n}2^{n}\Psi(2^{n})=\infty$. Hence $\sum_{q}\min(\psi(q),q^{-1})=\infty$ amounts to $\sum_{n}\min(2^{n}\psi(2^{n}),1)=\infty$ which in turns follows from $\sum_{n}2^{n}\psi(2^{n})=\infty$. ∎
Proof of A’.
As in the proof of B’, we may assume $\lambda$ is supported on $\operatorname{Aff}(\mathbb{R})^{+}$, see Lemma 5.3. Hence we are reduced to the setting of Section 2. By Proposition 6.1, $\sigma$ satisfies the effective double equidistribution property (7.2) (and in particular (7.1)). Hence both 7.2 and Corollary 7.4 apply, yielding the announced dichotomy. ∎
We now pass to the proof of 7.3. In a first step we will show that effective double equidistribution in fact yields decorrelation estimates that are valid for all times $t_{2}\geq t_{1}\geq 1$. Then we will exploit these estimates through the mean of Dani’s correspondence to deduce the theorem.
7.1. Single vs double equidistribution
Note that effective double equidistribution (7.2) implies effective single equidistribution (7.1). Conversely, effective single equidistribution gives a double equidistribution estimate in the short-range regime. The proof exploits the decay of matrix coefficients as in [24, Theorem 10.1].
Lemma 7.5.
Let $\xi$ be a Borel probability measure on $\mathbb{R}$ satisfiying (7.1) with associated constants $C>1,c\in(0,1)$. Then for every $t_{1},t_{2}\geq 1$ such that $t^{1+c/2}_{1}\geq t_{2}\geq t_{1}$ and every $f_{1},f_{2}\in B^{\infty}_{\infty,1}(X)$,
(7.6) $\Delta^{\xi}_{f_{1},f_{2}}(t_{1},t_{2})\ll{\mathcal{S}}_{2,1}(f_{1}){\mathcal{S}}_{2,1}(f_{2})t_{1}^{\delta_{0}}t_{2}^{-\delta_{0}}+C{\mathcal{S}}_{\infty,1}(f_{1}){\mathcal{S}}_{\infty,1}(f_{2})t^{-c/3}_{2}.$
where $\delta_{0}=\delta_{0}(\Lambda)>0$ arises from (5.1).
Proof.
Let $t_{2}\geq t_{1}\geq 1$ and $f_{1},f_{2}\in B^{\infty}_{\infty,1}(X)$. Set $F:X\to\mathbb{R}$ to be
$F(x)=f_{1}(x)f_{2}(a(t_{2}/t_{1})x),\quad x\in X,$
so that $F\bigl{(}a(t_{1})u(s)x_{0}\bigr{)}=f_{1}\bigl{(}a(t_{1})u(s)x_{0}\bigr{)}f_{2}\bigl{(}a(t_{2})u(s)x_{0}\bigr{)}$ for all $s\in\mathbb{R}$. Then
${\mathcal{S}}_{\infty,1}(F)$ $\ll{\mathcal{S}}_{\infty,1}(f_{1}){\mathcal{S}}_{\infty,1}(a(t_{1}/t_{2}).f_{2})\ll{\mathcal{S}}_{\infty,1}(f_{1})\lVert\operatorname{Ad}(a(t_{1}/t_{2}))\rVert{\mathcal{S}}_{\infty,1}(f_{2})$
$\ll t_{2}/t_{1}{\mathcal{S}}_{\infty,1}(f_{1}){\mathcal{S}}_{\infty,1}(f_{2}).$
By (7.1) applied to $F$ and $t_{1}$,
$\left\lvert\int_{\mathbb{R}}F\bigl{(}a(t_{1})u(s)x_{0}\bigr{)}\,\mathrm{d}\xi(s)-\langle f_{1},a(t_{1}/t_{2}).f_{2}\rangle\right\rvert\leq C{\mathcal{S}}_{\infty,1}(F)t_{1}^{-c}$
By (5.1), we have
$\left\lvert\langle f_{1},a(t_{1}/t_{2}).f_{2}\rangle-m_{X}(f_{1})m_{X}(f_{2})\right\rvert\ll\lVert a(t_{1}/t_{2})\rVert^{-\delta_{0}}{\mathcal{S}}_{2,1}(f_{1}){\mathcal{S}}_{2,1}(f_{2}).$
Combining the above together, we obtain
$\Delta^{\xi}_{f_{1},f_{2}}(t_{1},t_{2})\ll{\mathcal{S}}_{2,1}(f_{1}){\mathcal{S}}_{2,1}(f_{2})t_{1}^{\delta_{0}}t_{2}^{-\delta_{0}}+C{\mathcal{S}}_{\infty,1}(f_{1}){\mathcal{S}}_{\infty,1}(f_{2})t_{2}t_{1}^{-1-c}$
whence the desired inequality in the regime $t^{1+c/2}_{1}\geq t_{2}\geq t_{1}$. ∎
We deduce that even though double equidistribution was formulated with the separation assumption $t_{2}>t^{1+\eta}_{1}$ on the parameters $t_{1},t_{2}\geq 1$, it still provides estimates in the short-range regime $t_{2}\leq t^{1+\eta}_{1}$. Put together, we obtain the following result.
Corollary 7.6.
A probability measure $\xi$ on $\mathbb{R}$ satisfies the effective double equidistribution property (7.2) if and only if there exist constants $c>0$ and $C>1$ such that for every $f_{1},f_{2}\in B^{\infty}_{\infty,1}(X)$ and all $t_{2}\geq t_{1}>1$.
(7.7) $\Delta^{\xi}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{2,1}(f_{1}){\mathcal{S}}_{2,1}(f_{2})t_{1}^{c}t_{2}^{-c}\\ +C{\mathcal{S}}_{\infty,1}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,1}(f_{1}){\mathcal{S}}_{\infty,1}(f_{2})t_{2}^{-c},$
Proof.
As the ${\mathcal{S}}_{2,1}$-norm is bounded by the ${\mathcal{S}}_{\infty,1}$-norm, the converse direction is clear. Assume $\xi$ satisfies (7.2). Recalling that (7.2) implies (7.1), Lemma 7.5 applies and yields the upper bound in the short-range regime (possibly with different values of $C,c$). It also holds in the non short-range regime by definition of (7.2). ∎
7.2. Lower bound estimate
In this subsection, we establish the lower bound in our quantitative Khintchine dichotomy 7.3. Notations refer to 7.3, in particular $X$ here is $\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$ and $\psi(q)\leq q^{-1}$. We extend $\psi$ to a function $\mathbb{R}^{+}\to\mathbb{R}_{>0}$ by setting $\psi(q)=\psi(\lceil q\rceil)$ for non-integer values of $q$, so that it is still non-increasing and smaller than $q^{-1}$.
For $N\geq 1$ and $s\in\mathbb{R}$, we write $\mathscr{T}_{N}(s)$ for the left-hand side of (7.4), on which we aim to obtain a lower bound. We fix a parameter $\tau\in(1,2]$ and define for $k\geq 0$,
[…]
Letting $n\geq 1$ be such that $\tau^{n}\leq N<\tau^{n+1}$ and using that $\psi$ is non-increasing, we have
(7.8) $\mathscr{T}_{N}(s)\geq\mathscr{T}_{\tau^{n}}(s)\geq\sum_{k=1}^{n}\mathscr{S}_{k}(s).$
We bound below the sum on the right hand side.
Proposition 7.7.
Under the assumptions of 7.3, for $\xi$-almost all $s\in\mathbb{R}$, for every $\varepsilon>0$, for all large enough $n$, we have
(7.9) $\sum_{k=1}^{n}\mathscr{S}_{k}(s)\geq(1-\varepsilon)\zeta(2)^{-1}\sum_{k=1}^{n}(\tau^{k}-\tau^{k-1})\psi(\tau^{k}).$
The lower bound in 7.3 follows at once.
Proof of lower bound in (7.4) using Proposition 7.7.
In view of (7.8) and Proposition 7.7, it suffices to show that for any $\varepsilon>0$, there is some $\tau>1$ such that
$\sum_{k=1}^{n}(\tau^{k}-\tau^{k-1})\psi(\tau^{k})\geq(1-3\varepsilon)\sum_{q=1}^{N}\psi(q)$
whenever $\tau^{n}\leq N<\tau^{n+1}$ and $N$ is large enough (in terms of $\psi$ and $\varepsilon$).
Indeed, we can pick $\tau=1+\varepsilon$. Because $\psi$ is non-increasing, we have
$\sum_{q=\lceil\tau^{k}\rceil}^{\lceil\tau^{k+1}\rceil-1}\psi(q)\leq(\lceil\tau^{k+1}\rceil-\lceil\tau^{k}\rceil)\psi(\tau^{k})\leq(\tau+\varepsilon)(\tau^{k}-\tau^{k-1})\psi(\tau^{k})$
for all $k\geq 1$ large enough. Summing up to $k=n$ yields the desired inequality. ∎
We now turn to the proof of Proposition 7.7.
First, we invoke Dani’s correspondence to give the quantity $\mathscr{S}_{k}(s)$ a dynamical interpretation. Consider $X=\operatorname{SL}_{2}(\mathbb{R})/\operatorname{SL}_{2}(\mathbb{Z})$ and $x_{0}=\operatorname{SL}_{2}(\mathbb{Z})/\operatorname{SL}_{2}(\mathbb{Z})\in X$ the identity coset. For a function $f:\mathbb{R}^{2}\to[0,+\infty]$, we denote by $\widetilde{f}:X\to[0,+\infty]$ its primitive Siegel transform. It is defined by: $\forall g\in G$,
$\widetilde{f}(gx_{0})=\sum_{v\in{\mathcal{P}}(\mathbb{Z}^{2})}f(gv).$
For each $k\geq 1$, consider the quantities $r_{k},t_{k}\in\mathbb{R}_{>0}$ such that
$\tau^{k}=r_{k}t_{k}^{1/2},\qquad\psi(\tau^{k})=r_{k}t_{k}^{-1/2},$
or equivalently
(7.10) $r_{k}^{2}=\tau^{k}\psi(\tau^{k}),\qquad t_{k}=\tau^{k}\psi(\tau^{k})^{-1}.$
Consider the rectangle $R_{k}=[0,r_{k})\times(\tau^{-1}r_{k},r_{k}]\subseteq\mathbb{R}^{2}$. Direct computation shows that for any $s\in\mathbb{R}$,
$\mathscr{S}_{k}(s)=\widetilde{\mathbbm{1}}_{R_{k}}\bigl{(}a(t_{k})u(s)x_{0}\bigr{)}.$
Next, we construct smooth lower approximations $(\varphi_{k})_{k\geq 1}$ of the functions $(\widetilde{\mathbbm{1}}_{R_{k}})_{k\geq 1}$. This substitution will allow us to use equidistribution estimates. Let $\varepsilon>0$ be a (small) parameter. Let $R_{k}^{-}:=\big{[}\varepsilon r_{k},(1-\varepsilon)r_{k}\bigr{)}\times\bigl{(}(\tau^{-1}+\varepsilon)r_{k},(1-\varepsilon)r_{k}\big{]}\subseteq R_{k}$ denote the rectangle shrunken by $\varepsilon r_{k}$ on each side from $R_{k}$. Note that for every $g\in B_{\varepsilon/10}\subseteq G$, we have $gR_{k}^{-}\subseteq R_{k}$ and hence $g_{*}\widetilde{\mathbbm{1}}_{R_{k}^{-}}\leq\widetilde{\mathbbm{1}}_{R_{k}}$. Let $\theta_{\varepsilon}:G\to\mathbb{R}_{\geq 0}$ be a smooth bump function supported on $B_{\varepsilon/10}$ such that $m_{G}(\theta_{\varepsilon})=1$ and ${\mathcal{S}}_{\infty,1}(\theta_{\varepsilon})\ll\varepsilon^{-4}$. We set for every $k\geq 1$,
$\varphi_{k}:=\theta_{\varepsilon}*\widetilde{\mathbbm{1}}_{R_{k}^{-}}.$
In particular, $\varphi_{k}\leq\widetilde{\mathbbm{1}}_{R_{k}}$, so for every $s\in\mathbb{R}$,
(7.11) $\varphi_{k}\bigl{(}a(t_{k})u(s)x_{0}\bigr{)}\leq\mathscr{S}_{k}(s).$
We now discuss the norm properties of the functions $\varphi_{k}$.
Lemma 7.8.
For every $k\geq 1$, we have
(7.12) $m_{X}(\varphi_{k})=\zeta(2)^{-1}r_{k}^{2}(1-2\varepsilon)(1-\tau^{-1}-2\varepsilon).$
(7.13) ${\mathcal{S}}_{\infty,1}(\varphi_{k})\ll\varepsilon^{-1}\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,{\mathcal{S}}_{2,1}(\varphi_{k})\ll\varepsilon^{-1}\sqrt{m_{X}(\varphi_{k})}.$
Proof.
Note that by our assumption on $\psi$, we have $r_{k}\leq 1$ hence $R_{k}\subseteq[0,1)\times[0,1]$ contains at most $2$ primitive vectors of any unimodular lattice in $\mathbb{R}^{2}$. It follows that $\lVert\widetilde{\mathbbm{1}}_{R_{k}}\rVert_{L^{\infty}}\leq 2$ and then $\lVert\varphi_{k}\rVert_{L^{\infty}}\leq 2$.
By Siegel’s summation formula [45, Equation 25],
$m_{X}(\varphi_{k})$ $=m_{G}(\theta_{\varepsilon})m_{X}(\widetilde{\mathbbm{1}}_{R_{k}^{-}})=\zeta(2)^{-1}\text{leb}_{\mathbb{R}^{2}}(R^{-}_{k})$
$=\zeta(2)^{-1}r_{k}^{2}(1-2\varepsilon)(1-\tau^{-1}-2\varepsilon).$
Then
${\mathcal{S}}_{\infty,1}(\varphi_{k})\leq m_{G}(\operatorname{supp}\theta_{\varepsilon}){\mathcal{S}}_{\infty,1}(\theta_{\varepsilon})\lVert\widetilde{\mathbbm{1}}_{R_{k}^{-}}\rVert_{L^{\infty}}\ll\varepsilon^{-1}.$
Finally, using that $\widetilde{\mathbbm{1}}_{R_{k}}$ only takes integer values,
${\mathcal{S}}_{2,1}(\varphi_{k})\leq{\mathcal{S}}_{\infty,1}(\varphi_{k})\sqrt{m_{X}(\operatorname{supp}\varphi_{k})}\\ \ll\varepsilon^{-1}\sqrt{m_{X}(\operatorname{supp}\widetilde{\mathbbm{1}}_{R_{k}})}\leq\varepsilon^{-1}\sqrt{m_{X}(\widetilde{\mathbbm{1}}_{R_{k}})}\ll\varepsilon^{-1}\sqrt{m_{X}(\varphi_{k})}.$
where the last bound relies on (7.12) and Siegel’s summation formula again. ∎
We consider $(\mathbb{R},\xi)$ as a probability space. Expectation $\mathbb{E}[\,\cdot\,]$ refer implicitely to this probability space. Introduce for every $k\geq 1$, the random variable
$Y_{k}:\mathbb{R}\to\mathbb{R},\quad s\mapsto\varphi_{k}\bigl{(}a(t_{k})u(s)x_{0}\bigr{)}.$
Write
$y_{k}=m_{X}(\varphi_{k})\in[0,1]$
and set $Z_{k}=Y_{k}-y_{k}$ as the (quasi-recentered) companion of $Y_{k}$.
From the quantitative double equidistribution hypothesis on $\xi$, we deduce an upper bound on the second moment of a sum of $Z_{k}$’s.
Proposition 7.9.
In the setting of 7.3, assume additionally that
(7.14) $\psi(q)\geq q^{-1}\log^{-2}(q),\quad\forall q\geq 3.$
Then there is a constant $C^{\prime}$ such that for every subset $J\subseteq\mathbb{N}^{*}$ we have
$\mathbb{E}\Bigl{[}\Bigl{(}\sum\nolimits_{j\in J}Z_{j}\Bigr{)}^{2}\Bigr{]}\leq C^{\prime}\sum_{j\in J}y_{j}.$
Proof.
Let $C>1$ and $c>0$ be as in the full-range double equidistribution estimate (7.7). In this proof, the implied constants in the $\ll$ notation are allowed to depend on $C$, $\tau$ and $\varepsilon>0$.
By definition, for each $k,l\geq 1$,
$\mathbb{E}[Z_{k}Z_{l}]=\mathbb{E}[Y_{k}Y_{l}]-y_{k}y_{l}-\mathbb{E}[Z_{k}]y_{l}-y_{k}\mathbb{E}[Z_{l}].$
Combining (7.7) with the bounds on Sobolev norms from (7.13), we obtain for $k\leq l$,
$\lvert\mathbb{E}[Y_{k}Y_{l}]-y_{k}y_{l}\rvert\ll\sqrt{y_{k}y_{l}}t_{k}^{c}t_{l}^{-c}+y_{l}t_{k}^{-c}+t_{l}^{-c}.$
while by (7.1),
$\lvert\mathbb{E}[Z_{k}]\rvert\ll t_{k}^{-c}.$
By expanding the square power, using the above bounds, and recalling from (7.10) that $t_{k}\geq\tau^{k}$ and $t_{l}/t_{k}\geq\tau^{l-k}$ for $k\leq l$, we deduce
$\mathbb{E}\Bigl{[}\Bigl{(}\sum\nolimits_{j\in J}Z_{j}\Bigr{)}^{2}\Bigr{]}\ll\sum\nolimits_{k,l\in J,k\leq l}(\sqrt{y_{k}y_{l}}\tau^{-c(l-k)}+y_{l}\tau^{-ck}+\tau^{-cl}).$
Using $\sqrt{y_{k}y_{l}}\leq y_{k}+y_{l}$ and the convergence $\sum_{n=0}^{\infty}\tau^{-cn}<+\infty$, the first sum satisfies $\sum\nolimits_{k,l\in J,k\leq l}\sqrt{y_{k}y_{l}}\tau^{-c(l-k)}\ll\sum y_{j}$. The convergence $\sum_{n=0}^{\infty}\tau^{-cn}<+\infty$ bounds similarly the second sum. To bound the third sum, note that combining (7.10) with our assumption (7.14), then using Equation 7.12, we have
$\tau^{-ck}\ll(k\log\tau)^{-2}\leq r_{k}^{2}\ll y_{k}.$
Hence $\tau^{-cl}\ll y_{k}\tau^{-c(l-k)}$, so $\sum\nolimits_{k,l\in J,k\leq l}\tau^{-cl}\ll\sum y_{j}$ as for the first sum. ∎
The following lemma is a general fact about sequences of random variables. It is abstracted from Schmidt’s proof of the quantitative Khintchine theorem for the Lebesgue measure [43]. See also [47, Chapter I, Lemma 10], or [34, Lemma 2.6].
Lemma 7.10.
Let $(Y_{k})_{k\geq 1}$ be a sequence of non-negative real random variables. Let $(y_{k})_{k\geq 1}\in[0,1]^{\mathbb{N}^{*}}$ be a sequence of real numbers, set $Z_{k}=Y_{k}-y_{k}$. Assume that $\sum_{k=1}^{\infty}y_{k}=+\infty$ and for some $C_{1}\geq 1$
(7.15) $\forall n\geq m\geq 1,\quad\mathbb{E}\Bigl{[}\bigl{(}\sum_{k=m}^{n}Z_{k}\bigr{)}^{2}\Bigr{]}\leq C_{1}\sum_{k=m}^{n}y_{k}.$
Then almost surely, there is a constant $C_{2}>1$ such that
$\forall n\geq 1,\quad\Bigl{\lvert}\sum_{k=1}^{n}Z_{k}\Bigr{\rvert}\leq C_{2}\Bigl{(}\sum_{k=1}^{n}y_{k}\Bigr{)}^{1/2}\log^{2}\Bigl{(}\sum_{k=1}^{n}y_{k}\Bigr{)}.$
We are now able to conclude the proof of Proposition 7.7, whence that of the lower bound in 7.3.
Proof of Proposition 7.7.
The series $\sum_{q}q^{-1}\log^{-2}(q)$ is convergent. Thus, by the convergent case of the Khintchine dicothomy for measures satisfying (7.1) (7.2), we know that if we replace $\psi$ by $q\mapsto\max\{\psi(q),q^{-1}\log^{-2}(q)\}$ (say for $q\geq 3$, and by $q\mapsto 1/2$ else), then for $\xi$-almost every $s\in\mathbb{R}$, the left-hand side of (7.9) is increased by only a bounded amount. For this reason, without loss of generality, we can assume (7.14).
Note that in view of the inequality (7.11), we have $\sum_{k=1}^{n}\mathscr{S}_{k}\geq\sum_{k=1}^{n}Y_{k}$. Equations (7.12) and (7.10) yield $1\geq y_{k}\geq\zeta(2)^{-1}(1-O(\varepsilon))(\tau^{k}-\tau^{k-1})\psi(\tau^{k})$, in particular $\sum_{k=1}^{\infty}y_{k}=\infty$. This estimate, combined with the previous paragraph and the variance bound Proposition 7.9, allows to apply Lemma 7.10 to get that $\xi$-almost everywhere, $\sum_{k=1}^{n}Y_{k}\sim\sum_{k=1}^{n}y_{k}\geq\zeta(2)^{-1}(1-O(\varepsilon))\sum_{k=1}^{n}(\tau^{k}-\tau^{k-1})\psi(\tau^{k})$. This concludes the proof. ∎
7.3. Upper bound estimate
The proof of the upper bound in 7.3 (Equation 7.4) is similar. We extend $\psi$ to $\mathbb{R}^{+}$ by setting $\psi(q)=\min(q^{-1},\psi(\lfloor q\rfloor)$ for non-integer values of $q$. We have for $\tau^{n}\leq N<\tau^{n+1}$ and for every $s\in\mathbb{R}$,
$\mathscr{T}_{N}(s)\leq\sum_{k=0}^{n}\mathscr{S}_{k}^{+}(s)$
where for every $k\geq 0$,
[…]
Then
$\mathscr{S}_{k}^{+}(s)=\widetilde{{\mathbbm{1}}}_{[0,r_{k})\times[r_{k},\tau r_{k})}\bigl{(}a(t_{k})u(s)x_{0}\bigr{)}\leq\varphi_{k}^{+}\bigl{(}a(t_{k})u(s)x_{0}\bigr{)}$
where $\varphi_{k}^{+}=\theta_{\varepsilon}*\widetilde{\mathbbm{1}}_{[-\varepsilon r_{k},(1+\varepsilon)r_{k})\times[(1-\varepsilon)r_{k},(\tau+\varepsilon)r_{k})}$. For the rest of the proof, we can use mutatis mutandis the argument for the lower bound.